### v1

# Parameter estimation of range-migrating targets using OTFS signals from LEO satellites

Abstract
This study investigates a communication-centric integrated sensing and communication (ISAC) system that utilizes orthogonal time frequency space (OTFS) modulated signals emitted by low Earth orbit (LEO) satellites to estimate the parameters of space targets experiencing range migration, henceforth referred to as high-speed targets. Leveraging the specific signal processing performed by OTFS transceivers, we derive a novel input-output model for the echo generated by a high-speed target in scenarios where ideal and rectangular shaping filters are employed. Our findings reveal that the target response exhibits a sparse structure in the delay-Doppler domain, dependent solely upon the initial range and range-rate; notably, range migration causes a spread in the target response, marking a significant departure from previous studies. Utilizing this signal structure, we propose an approximate implementation of the maximum likelihood estimator for the target’s initial range, range-rate, and amplitude. The estimation process involves obtaining coarse information on the target response using a block orthogonal matching pursuit algorithm, followed by a refinement step using a bank of matched filters focused on a smaller range and range-rate region. Finally, numerical examples are provided to evaluate the estimation performance.
Index Terms:
Communication-centric ISAC, opportunistic sensing, OTFS modulation, LEO satellites, space targets, high-speed targets, range migration, maximum likelihood estimation.
I Introduction
Future wireless networks and Internet-of-Things (IoT) will utilize terrestrial and non-terrestrial stations to ensure seamless and ubiquitous connectivity [1]. Among spaceborne transceivers, low Earth orbit (LEO) satellites are attracting much attention for their lower latency and higher data rates [2]; also, they may efficiently provide integrated sensing and communications (ISAC) functionalities over large regions with the same signal waveform and hardware architecture [3, 4, 5].
Orthogonal time frequency space (OTFS) modulation is a promising transmission scheme for future LEO constellations [6, 7]. OTFS-based transceivers exploit channel diversity across both time and frequency, thus providing better communication performance in doubly-dispersive fading channels as compared to schemes using orthogonal frequency division multiplexing [8, 9, 10]. They are also suitable for radar sensing, as the underlying signal processing resembles that of pulsed-Doppler radars [11]; in particular, the range and range-rate [^1] In this study, range-rate refers to the radial component of velocity, i.e., the rate at which the distance between the object and the observer is changing. of prospective scattering objects can be inferred from the estimated delay-Doppler channel response [12]. These unique features enable OTFS-based LEO satellites to inherently support ISAC in various forms [13, 14]. Unlike terrestrial stations, LEO-based stations have the capability to also cover aerial and space regions, and this expanded coverage opens up new application scenarios, such as real-time environmental monitoring, enhanced global navigation systems, and space exploration. Communication-centric ISAC solutions appear particularly promising in this context, as existing data signals can be opportunistically reused to add the sensing functions at a low cost, without altering the underlying communication protocol or affecting its performance [12, 15, 16, 17, 18]; this approach is indeed considered in this study.
Several solutions for signal estimation and detection have been deployed for OTFS-based applications, which exploit the channel sparsity in the delay-Doppler domain, and different complexity-performance tradeoffs have been achieved by modeling the range and range-rate of any path component as either integer or fractional multiples of the range and range-rate resolutions, respectively: [^2] For a given path component, the delay and Doppler parameters are equivalent to its range and range-rate, respectively, with a convenient scaling factor; these terms are often used interchangeably in the literature, and, for the remainder of this study, we will prefer the latter terminology. see [19] and references therein. For example, the work in [8] has derived a widely-used input-output signal model under the assumption of integer range and fractional range-rate for both ideal and rectangular shaping filters and has presented a message passing algorithm for data symbol detection. [^3] Hereafter, integer (fractional) range or range-rate means that the range or range-rate is an integer (fractional) multiple of the respective resolution. Leveraging the model in [8], a pilot-aided channel estimation has been developed in [20]; here, pilot, guard, and data symbols are arranged so as to avoid interference between pilot and data symbols at the receiver, thus facilitating channel estimation and data detection. Another channel estimation method based on a sparse Bayesian learning algorithm has been provided in [21] and, then, extended in [22] to handle both fractional range and range-rate. More recently, a channel estimation method relying on the insertion of random pilots has been presented in [23]. On a different side, leveraging the model in [8], the work in [12] has proposed to exploit existing OTFS communication signals for radar sensing to implement a simplified form of communication-centric ISAC. This idea has been further elaborated in [24], where the Cramér-Rao lower bounds in the estimation of the range and range-rate are derived, in [25], where different pulse shapes are investigated, in [26], where a modified orthogonal matching pursuit algorithm with fractional refinement is proposed, and in [27], where the estimation of the target parameter is formulated as an atomic norm minimization. The 3D-structured sparsity of MIMO-OTFS channel is further explored in [28], where the downlink channel estimation problem is formulated as a sparse signal recovery problem. In addition, OTFS signals have been used for joint localization and communications in [29] and for target tracking in [30].
Previous studies assume that the length (range) of a channel path between the transmitter and receiver remains constant within the OTFS frame, which is typically valid for terrestrial applications. However, this assumption may not hold true for satellite-based space sensing, as the significant relative motion between the transmitter, monitored target, and receiver can cause range migration during the OTFS frame. With a transmitter on a LEO satellite and a radar receiver on the same or a different satellite, or an aircraft at a lower altitude, non-cooperative space targets (e.g., space debris, asteroids, spacecraft, and unresponsive satellites) may have a range-rate of several km/s; for a system with $M=512$ subcarriers, $N=128$ symbol intervals, a subcarrier spacing of $\Delta=15$ kHz, a symbol interval of $T=1/\Delta$, and a range-rate of $\bar{v}=15$ km/s, the range variation in the OTFS frame is $\bar{v}NT=128$ m and greatly exceeds the range resolution $c/(2M\Delta)\approx 19.5$ m, where $c$ is the speed of light.
In this study, we investigate the use of OTFS communication signals emitted by LEO satellites for estimating the initial range, range-rate, and amplitude of space targets experiencing range migration in the OTFS frame, henceforth referred to as high-speed targets for brevity. We follow a communication-centric approach where the existing waveform and protocol are not modified to maintain the communication performance unaltered; instead, the focus is on the design of the opportunistic radar receiver and on the analysis of its achievable performance. Sensing high-speed targets is a challenging problem, and previous studies in the radar literature have primarily utilized unmodulated or modulated coherent pulse trains (typically, a train of linear frequency modulated pulses) and developed motion compensation strategies operating in the time or frequency domain to integrate consecutive pulses [31, 32, 33, 34]. However, the impact of range migration on OTFS signals is unexplored, and existing motion compensation methodologies cannot directly be applied in the delay-Doppler domain when off-the-shelf OTFS modulator and demodulator are employed. This study aims to provide initial insights on parameter estimation of high-speed targets using OTFS communication signals, without changing the structure of existing transceivers. The major contributions are summarized below.
  • We derive a novel model for the echo produced by a high-speed target considering both ideal and rectangular shaping filters; in particular, the target range is assumed constant over a block of a few symbols and changes from block-to-block consistently with the range-rate. The proposed model subsumes the one in [8] when the target range migration is negligible and the target range is an integer multiple of the range resolution.
  • We demonstrate that the response of high-speed targets maintains a sparse structure in the delay-Doppler domain, determined entirely by the initial range and range-rate. Additionally, we show that range migration results in a spread of this response across both dimensions.
  • We propose an approximate implementation of the maximum likelihood (ML) estimator for target parameters, which avoids exhaustive delay-Doppler search. First, a block orthogonal matching pursuit (BOMP) algorithm [35] is used to obtain coarse information on the initial range and range-rate. Next, a refined estimate of the target parameters is derived using a bank of matched filters focused on a narrower delay-Doppler region. The proposed procedure can be combined with the CLEAN algorithm in [36, 37, 38, 39, 15] to handle multiple targets.
  • Numerical simulations are conducted to validate the proposed echo model and assess the performance of the target parameter estimator under both ideal and rectangular shaping filters. Results are also compared against those obtained using the methods in [20] and [28] that assume no range migration occurs within the OTFS frame.
Organization
The remainder of the paper is organized as follows. Sec. II presents the system description and the design assumptions. Secs. III and IV provide the proposed signal model for ideal and rectangular shaping filters, respectively. Sec. V illustrates the proposed procedure for target parameter estimation. Sec. VI discusses some numerical examples. Sec. VII contains the conclusions. Finally, some mathematical derivations are deferred to the Appendix.
Notation
In the following, $\mathbb{R}$ and $\mathbb{C}$ are the set of real and complex numbers, respectively. Column vectors and matrices are denoted by lowercase and uppercase boldface letters, respectively. The symbols $(\,\cdot\,)^{*}$, $(\,\cdot\,)^{\mathsf{T}}$, and $(\,\cdot\,)^{\mathsf{H}}$ denote conjugate, transpose, and conjugate-transpose, respectively. $\bm{I}_{M}$ is the $M\times M$ identity matrix. $(\bm{a})_{i}$ denotes the $i$-th entry of the vector $\bm{a}$. $(\bm{A})_{i,j}$ denotes the entry in the $i$-th row and $j$-th column of the matrix $\bm{A}$. $(\bm{A}_{1}\ \cdots\ \bm{A}_{n})$ and $(\bm{A}_{1};\ \cdots\ ;\bm{A}_{n})$ denote the horizontal and vertical concatenations of the matrices $\bm{A}_{1},\ldots,\bm{A}_{n}$. $\mathbbm{1}_{\mathcal{A}}$ is the indicator function of the condition $\mathcal{A}$, i.e., $\mathbbm{1}_{\mathcal{A}}=1$, if $\mathcal{A}$ holds true, and $\mathbbm{1}_{\mathcal{A}}=0$ otherwise. $[\,\cdot\,]_{M}$ denotes the mod $M$ operation. $\delta(t)$ and $\delta[n]$ denote the Dirac delta function and the discrete unit sample function, respectively. ${}\Re\left\{\cdot\right\}$ denotes the real part. Finally, $\mathrm{i}$, $\odot$, and $\mathrm{E}[\,\cdot\,]$ denote the imaginary unit, the Schur product, and the statistical expectation, respectively.
II System description
Consider a sensing system consisting of a communication transmitter and a radar receiver. The communication transmitter is mounted on a LEO satellite, uses an OTFS modulation, and operates with a carrier frequency $c/\lambda$, where $\lambda$ is the carrier wavelength; the emitted waveform is specified by the underlying communication protocol and cannot be modified to adjust the radar needs. The radar receiver aims to estimate the initial range (delay), range-rate (Doppler shift), and amplitude of high-speed targets illuminated by the communication transmitter; it may be mounted on the same platform hosting the communication transmitter (monostatic configuration) or on a separate device (bistatic configuration), e.g., another LEO satellites, an aircraft, a high altitude platform, or a terrestrial station. The receiver is synchronized with the transmitter and knows the emitted symbols: for a monostatic configuration, this is readily verified [15], whereas, for a bistatic configuration, this requires the presence of some coordination [^4] For example, when the radar receiver is on a different LEO satellite, the coordination may be established via an inter-satellite link. or a separate reference channel able to demodulate the communication signal [40, 41]. For a more illustration, we consider a bistatic scenario, as shown in Fig. 1, but the following developments similarly apply to the monostatic case. We initially assume the presence of a single target to simplify exposition.
Figure 1: System geometry.
II-A Transmit signal
[…]
$x(t)=\sum_{n=0}^{N-1}\underbrace{\sum_{m=0}^{M-1}X_{\rm TF}[n,m]g_{\rm tx}\!\left(t-nT\right)\mathrm{e}^{\mathrm{i}2\pi m\Delta(t-nT)}}_{x_{n}(t)},$ (1)
where $x_{n}(t)$ is the signal emitted in the $n$-th symbol interval, $X_{\rm TF}[n,m]$ is the time-frequency symbol sent in the $n$-th symbol interval on subcarrier $m$, and $g_{\rm tx}(t)$ is the transmit shaping filter with unit energy, effective temporal duration $T$, and effective spectral width $\Delta$. The time-frequency sequence $X_{\rm TF}[n,m]$ is obtained as the inverse Symplectic Finite Fourier Transform of the delay-Doppler sequence $X_{\rm DD}[k,l]$, i.e.,
$X_{\rm TF}[n,m]=\frac{1}{\sqrt{NM}}\sum_{k=0}^{N-1}\sum_{l=0}^{M-1}X_{\rm DD}[k,l]\mathrm{e}^{\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)},$ (2)
for $n=0,\ldots,N-1$ and $m=0,\ldots,M-1$. By considering $g_{\rm tx}(t)$ time limited in $[0,T]$, then $x(t)$ has support $[0,NT]$. Also, by assuming the spectrum of $g_{\rm tx}(t)$ contained in $[-\Delta/2,\Delta/2]$, then the spectrum of $x(t)$ is contained in $[-\Delta/2,M\Delta-\Delta/2]$. [^6] Since a time limited signal cannot be strictly band limited, some spectral leakage outside $[-\Delta/2,M\Delta-\Delta/2]$ is present. In practice, the transmit shaping filter may not be exactly zero outside $[0,T]$ to reduce such spectral leakage; in this latter case, $x(t)$ is only approximately contained in $[0,NT]$.
II-B Design assumptions
During an OTFS frame, the target, the communication transmitter, and the radar receiver can be assumed to move with a constant velocity. The position of the target at time $t\in[0,NT]$ is $\bm{p}(t)=\bm{p}-\bm{v}t\in\mathbb{R}^{3}$, where $\bm{p}$ and $\bm{v}$ are the initial position and the velocity, respectively; similarly, the positions of communication transmitter and the radar receiver at time $t\in[0,NT]$ are $\bm{p}_{\rm tx}(t)=\bm{p}_{\rm tx}-\bm{v}_{\rm tx}t\in\mathbb{R}^{3}$ and $\bm{p}_{\rm rx}(t)=\bm{p}_{\rm rx}-\bm{v}_{\rm rx}t\in\mathbb{R}^{3}$, respectively. Accordingly, for $t\in[0,NT]$, the (bistatic) target range is
$\bar{r}(t)=\left\|\bm{p}(t)-\bm{p}_{\rm tx}(t)\right\|+\left\|\bm{p}(t)-\bm{p}_{\rm rx}(t)\right\|.$ (3)
Upon ignoring the quadratic and higher order terms in the Taylor expansion of the above expression, we have that [32]
$\bar{r}(t)=\bar{d}-\bar{v}t,$ (4)
where $\bar{d}$ and $\bar{v}$ are the initial range and the range-rate, respectively, defined as
$\bar{d}$ $=\left\|\bm{p}-\bm{p}_{\rm tx}\right\|+\left\|\bm{p}-\bm{p}_{\rm rx}\right\|,$ (5a)
$\bar{v}$ $=\frac{\left(\bm{p}-\bm{p}_{\rm tx}\right)^{\mathsf{T}}\!\left(\bm{v}_{\rm tx}-\bm{v}\right)}{\left\|\bm{p}-\bm{p}_{\rm tx}\right\|}+\frac{\left(\bm{p}-\bm{p}_{\rm rx}\right)^{\mathsf{T}}\!\left(\bm{v}_{\rm rx}-\bm{v}\right)}{\left\|\bm{p}-\bm{p}_{\rm rx}\right\|}.$ (5b)
Denote by $\bar{r}_{\min}$ and $\bar{r}_{\max}$ the minimum and maximum ranges, respectively, and by $\bar{v}_{\min}$ and $\bar{v}_{\max}$ the minimum and maximum range-rates, respectively. These values are tied to the prior uncertainty on the target location and mobility, as resulting from a previous detection phase (if the radar is operating in search mode) or as predicted by the tracking stage (if the radar is operating in tracking mode), and are assumed known. Also, let $r_{\max}=\bar{r}_{\max}-\bar{r}_{\min}$ and $v_{\max}=\bar{v}_{\max}-\bar{v}_{\min}$ be the length of the inspected range interval and range-rate interval, respectively. Then, we assume
$\tau_{\max}$ $=\frac{r_{\max}}{c}<T,$ (6a)
$\nu_{\max}$ $=\frac{v_{\max}}{\lambda}<\Delta,$ (6b)
so as to avoid ambiguity in the estimation of the initial range and the range-rate of the target, respectively.
Finally, let $N$ be a positive integer multiple of $B$. Then, we assume that $\bar{r}(t)$ is constant over at least $N/B$ consecutive symbol intervals: we refer to this as the $B$-fold stop-and-go approximation on the range variation over the OTFS frame. Under this design assumption, we can express:
$\bar{r}(t)=\bar{r}(bT_{B}),\quad\forall t\in[bT_{B},(b+1)T_{B}),$ (7)
with $T_{B}=NT/B$ and $b=0,\ldots,B-1$. For a fixed system bandwidth, increasing $B$ improves the model’s accuracy, particularly when the target exhibits higher range-rates or the OTFS frame duration is long (see Sec. VI for further discussion). Notably, prior studies assume $B=1$ at the design stage, effectively neglecting any range migration across the entire OTFS frame. In contrast, setting $B=N$ limits this approximation to only a single symbol interval, thereby capturing range migration with a much finer granularity.
II-C Received signal
Upon neglecting the noise contribution, the radio-frequency target echo at the radar receiver is
$\bar{y}_{\rm RF}(t)=\Re\left\{\bar{\alpha}x\left(t-\frac{\bar{r}(t)}{c}\right)\mathrm{e}^{\mathrm{i}2\pi\frac{c}{\lambda}\left(t-\frac{\bar{r}(t)}{c}\right)}\right\},$ (8)
where $\bar{\alpha}\in\mathbb{C}$ is the target amplitude (accounting for the radiated energy, the transmit and receive antenna gains, the target radar cross-section, and any signal attenuation). For convenience, we advance the time axis by $\bar{r}_{\min}/c$, thus obtaining
$\bar{y}_{\rm RF}\left(t+\frac{\bar{r}_{\min}}{c}\right)=\Re\left\{y(t)\mathrm{e}^{\mathrm{i}2\pi\frac{c+\bar{v}_{\min}}{\lambda}t}\right\},$ (9)
where
$y(t)=\bar{\alpha}x\left(t-\frac{r(t)}{c}\right)\mathrm{e}^{-\mathrm{i}2\pi\frac{r(t)}{\lambda}}\mathrm{e}^{-\mathrm{i}2\pi\frac{\bar{v}_{\min}}{\lambda}t}$ (10)
is the baseband representation of the target echo with respect to the reference frequency $\frac{c+\bar{v}_{\min}}{\lambda}$ and $r(t)=\bar{r}(t)-\bar{r}_{\min}$ is the excess target range (i.e., the target range in excess with respect to $\bar{r}_{\min}$). The support of $y(t)$ is contained in the time interval $[0,NT+r_{\max}/c]$, and its spectrum is negligible outside the frequency interval $[-\Delta/2,M\Delta+v_{\max}-\Delta/2]$.
Define the excess initial range $d=\bar{d}-\bar{r}_{\min}\in[0,r_{\max}]$, the excess range-rate $v=\bar{v}-\bar{v}_{\min}\in[0,v_{\max}]$, and $\alpha=\bar{\alpha}\mathrm{e}^{-\mathrm{i}2\pi\frac{d}{\lambda}}$. If the variation of the target range in a symbol interval can be ignored (which is the case under the $B$-fold stop-and-go approximation), then (10) can be expanded as
$y(t)=\sum_{n=0}^{N-1}y_{n}(t),$ (11)
where
$y_{n}(t)$ $=\alpha x_{n}\left(t-\frac{r\left(nT\right)}{c}\right)\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}t}$
[…]
and
$h_{n}(\tau,\nu)=\alpha\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(nT\right)}{c}}\delta\left(\tau-\frac{r\left(nT\right)}{c}\right)\delta\left(\nu-\frac{v}{\lambda}\right).$ (13)
Notice here that $y_{n}(t)$ is the output of a doubly-selective channel with delay-Doppler spread function $h_{n}(\tau,\nu)$ when the input is $x_{n}(t)$; the spread function depends upon $n$, as the target range can change during the OTFS frame.
II-D OTFS demodulation
The radar receiver is implemented on a device than relies on its OTFS demodulator to obtain discrete samples of the received signal $y(t)$ in the delay-Doppler domain; such samples are then processed to estimate the target parameters as shown in Sec. V. The OTFS demodulator first transform $y(t)$ into the time-frequency domain via the Wigner Transform; in particular, the following measurements are obtained
[…]
for $n=0,\ldots,N-1$ and $m=0,\ldots,M-1$, where $g_{\rm rx}(t)$ is the receive shaping filter with unit energy, effective temporal duration $T$ and effective spectral width $\Delta$. Denote by
[…]
the cross-ambiguity function of $g_{\rm tx}(t)$ and $g_{\rm rx}(t)$; then, the following proposition establishes a connection between $\{Y_{\rm TF}[n,m]\}$ and $\{X_{\rm TF}[n,m]\}$ that generalizes the result in [8, Theorem 1] to the case where the variation of the target range is ignored only within each symbol interval.
Proposition 1.
If the variation of the target range within each symbol interval can be ignored, then
$Y_{\rm TF}[n,m]=\sum_{n^{\prime}=0}^{N-1}\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n^{\prime},m^{\prime}]X_{\rm TF}[n^{\prime},m^{\prime}],$ (16)
for $n=0,\ldots,N-1$ and $m=0,\ldots,M-1$, where
$H_{n,m}[n^{\prime},m^{\prime}]=\alpha\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}nT}\mathrm{e}^{-\mathrm{i}2\pi m^{\prime}\Delta\frac{r\left(n^{\prime}T\right)}{c}}\\ \times\gamma\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c},(m-m^{\prime})\Delta-\frac{v}{\lambda}\right).$ (17)
Proof.
Upon plugging (11) into (14), we have
$Y_{\rm TF}[n,m]=\sum_{n^{\prime}=0}^{N-1}\sum_{m^{\prime}=0}^{M-1}\alpha X_{\rm TF}[n^{\prime},m^{\prime}]$
×$\quad\times\int_{-\infty}^{\infty}g_{\rm tx}\!\left(t-n^{\prime}T-\frac{r\left(n^{\prime}T\right)}{c}\right)g^{*}_{\rm rx}\!\left(t-nT\right)$
×$\quad\times\mathrm{e}^{\mathrm{i}2\pi m^{\prime}\Delta\left(t-n^{\prime}T-\frac{r\left(n^{\prime}T\right)}{c}\right)}\mathrm{e}^{-\mathrm{i}2\pi m\Delta\left(t-nT\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}t}dt$
$=\sum_{n^{\prime}=0}^{N-1}\sum_{m^{\prime}=0}^{M-1}X_{\rm TF}[n^{\prime},m^{\prime}]$
×$\quad\times\alpha\mathrm{e}^{\mathrm{i}2\pi m\Delta\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c}\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\left(n^{\prime}T+\frac{r\left(n^{\prime}T\right)}{c}\right)}$
×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c}\right)}$
×$\quad\times\int_{-\infty}^{\infty}g_{\rm tx}\!\left(\beta\right)g^{*}_{\rm rx}\!\left(\beta-(n-n^{\prime})T+\frac{r\left(n^{\prime}T\right)}{c}\right)$
×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left(\beta-(n-n^{\prime})T+\frac{r\left(n^{\prime}T\right)}{c}\right)}d\beta$
$=\sum_{n^{\prime}=0}^{N-1}\sum_{m^{\prime}=0}^{M-1}X_{\rm TF}[n^{\prime},m^{\prime}]\alpha\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}nT}\mathrm{e}^{-\mathrm{i}2\pi m^{\prime}\Delta\frac{r\left(n^{\prime}T\right)}{c}}$
×$\quad\times\gamma\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c},(m-m^{\prime})\Delta-\frac{v}{\lambda}\right),$
which completes the proof. ∎
Finally, the OTFS demodulator transforms $Y_{\rm TF}[n,m]$ into the delay-Doppler domain via the Symplectic Finite Fourier Transform; in particular, we have
$Y_{\rm DD}[k,l]=\frac{1}{\sqrt{NM}}\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}Y_{\rm TF}[n,m]\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)},$ (18)
for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$. In the Secs. III and IV we will establish an explicit relationship between $\{Y_{\rm DD}[n,m]\}$ and $\{X_{\rm DD}[n,m]\}$ for the cases where ideal and rectangular shaping filters are employed, respectively.
III Ideal shaping filters
Consider ideal shaping filters satisfying the bi-orthogonality condition [8, Sec. III]. Under this assumption, the ambiguity function satisfies $\gamma(\tau,\nu)=1$ for $(\tau,\nu)\in[-\tau_{\max},\tau_{\max}]\times[-\nu_{\max},+\nu_{\max}]$ and $\gamma(\tau,\nu)=0$ for $(\tau,\nu)\in[nT-\tau_{\max},nT+\tau_{\max}]\times[m\Delta-\nu_{\max},m\Delta+\nu_{\max}]$ for all pairs of integers $(m,n)\neq(0,0)$. Although practical waveforms can only approximate this condition to a certain extent, the use of ideal shaping filters offers valuable design insights, and the resulting system performance serves as a benchmark for comparison with that achievable using rectangular shaping filters.
In this case, (16) becomes
$Y_{\rm TF}[n,m]=H_{n,m}[n,m]X_{\rm TF}[n,m],$ (19)
for $n=1,\ldots,N-1$ and $m=0,\ldots,M-1$, where
$H_{n,m}[n,m]=\alpha\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}nT}\mathrm{e}^{-\mathrm{i}2\pi\frac{r\left(nT\right)}{c}m\Delta}.$ (20)
Also, denote by
$\mathcal{D}_{Q}(\nu)=\frac{1}{Q}\sum_{q=0}^{Q-1}\mathrm{e}^{-\mathrm{i}2\pi\nu q}=\frac{\mathrm{e}^{-\mathrm{i}2\pi\nu Q}-1}{Q\left(\mathrm{e}^{-\mathrm{i}2\pi\nu}-1\right)}$ (21)
the periodic sinc function (also called Dirichlet function) of degree $Q$. Then, the following proposition establishes a connection between $\{Y_{\rm DD}[k,l]\}$ and $\{X_{\rm DD}[k,l]\}$ that generalizes the result in [8, Proposition 2] to the case where the target range can change every $N/B$ symbol intervals.
Proposition 2.
Assume that ideal shaping filters are employed. Then, under the $B$-fold stop-and-go approximation on the range variation over the OFTS frame, we have
$Y_{\rm DD}[k,l]=\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\!X_{\rm DD}[k^{\prime},l^{\prime}]\Phi[k-k^{\prime},l-l^{\prime}],$ (22)
for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$, where
$\Phi[k,l]$ $=\frac{1}{B}\sum_{b=0}^{B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k}{N}-\frac{vT}{\lambda}\right)$
×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (23)
Proof.
After exploiting (2) and (19), Eq. (18) becomes
$Y_{\rm DD}[k,l]=\frac{1}{NM}\!\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\!X_{\rm DD}[k^{\prime},l^{\prime}]\\ \times\underbrace{\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}H_{n,m}[n,m]\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}\right)n}\mathrm{e}^{\mathrm{i}2\pi\left(\frac{l-l^{\prime}}{M}\right)m}}_{(\star)}.$ (24)
Finally, notice that we have
(($\star$)( ⋆ ) $=M\alpha\sum_{n=0}^{N-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}$
×$\quad\times\left[\frac{1}{M}\sum_{m=0}^{M-1}\mathrm{e}^{\mathrm{i}2\pi\left(\frac{l-l^{\prime}}{M}-\frac{r\left(nT\right)\Delta}{c}\right)m}\right]$
$=M\alpha\sum_{n=0}^{N-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}$
×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(nT\right)\Delta}{c}\right)\right)$
$=NM\alpha\sum_{b=0}^{B-1}\frac{1}{B}\left[\frac{B}{N}\sum_{n=bN/B}^{(b+1)N/B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}\right]$
×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right)$ (25)
$=NM\alpha\Phi[k-k^{\prime},l-l^{\prime}],$
where (25) follows from (7). ∎
III-A Some insights on Proposition 2
Recall that $\mathcal{D}_{Q}(\nu)$ is a periodic function of $\nu$ with period one and is zero at multiple of $1/Q$, except for $\nu=0,\pm 1,\pm 2,\ldots$ where it is one; also, $|\mathcal{D}_{Q}(\nu)|$ has a main-lobe centered at 00 of width $2/Q$ (between zero crossings) and $Q-2$ sidelobes of width $1/Q$ (between zero crossings) in $[-0.5,0.5]$. The first sidelobe is approximately $13$ dB down from the main-lobe, and the subsequent sidelobes fall off at about $6$ dB per octave [42]; following [8], for large $Q$, we can assume $|\mathcal{D}_{Q}(\nu)|\approx 0$ for $\nu\notin[-D/Q,D/Q]$, with $0<2D+1\ll Q$, where the non-negative integer $D$ is a design parameter that depends upon the desired approximation error. Accordingly, the following remarks are in order.
Remark 1.
The sequence $\Phi[k,l]$ in (23) is a periodic function of $k$ and $l$ with period $N$ and $M$, respectively. This implies that (22) is a two-dimensional circular convolution, whereby
$Y_{\rm DD}[k,l]$ $=\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\!X_{\rm DD}[k^{\prime},l^{\prime}]\Phi\Big{[}[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\Big{]}$ (26a)
$=\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\!X_{\rm DD}\Big{[}[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\Big{]}\Phi[k^{\prime},l^{\prime}],$ (26b)
for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$. ∎
Remark 2.
For large $N/B$ and $M$, the sums in (26) only contain few dominant terms. To illustrate this point, write first the range at time $bT_{B}$ and the range-rate as
$r\left(bT_{B}\right)$ $=(l_{b,{\rm int}}+l_{b,{\rm fra}})R_{\rm r},$ (27a)
$v$ $=(k_{\rm int}+k_{\rm fra})R_{\rm rr},$ (27b)
for $b=0,\ldots,B-1$, where $l_{b,{\rm int}}\in\{0,1,\ldots,M-1\}$, $k_{\rm int}\in\{0,1,\ldots,N-1\}$, and $l_{b,{\rm fra}},\,k_{\rm fra}\in(-0.5,0.5]$. Then, from (23) and (27) we have
$\Phi[k,l]$ $=\frac{1}{B}\sum_{b=0}^{B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k_{\rm int}-k_{\rm fra}}{B}\right)b}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k_{\rm int}-k_{\rm fra}}{N}\right)$
×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l_{b,{\rm int}}-l_{b,{\rm fra}}}{M}\right)\right).$ (28)
Finally, from the properties of the Dirichlet function discussed in Sec. III-A, we have $|\Phi[k,l]|\approx 0$ if $[k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $[l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$, for $b=0,\ldots,B-1$: more on this in Remark 4. ∎
Remark 3.
Assume that $l_{b,{\rm int}}=l_{{\rm int}}$ and $l_{b,{\rm fra}}=l_{{\rm fra}}$ for $b=0,\ldots,B-1$, which means that there is no range migration and that the target range is an integer multiple of the range resolution. Then, upon plugging (28) in (26b) and after straightforward manipulations, we obtain
$Y_{\rm DD}[k,l]=\alpha\sum_{\bar{k}=-\lfloor N/2\rfloor}^{\lceil N/2\rceil-1}\sum_{\bar{l}=-\lfloor M/2\rfloor}^{\lceil M/2\rceil-1}\mathcal{D}_{N}\!\left(-\frac{\bar{k}+k_{\rm fra}}{N}\right)\\ \times\mathcal{D}_{M}\!\left(\frac{\bar{l}+l_{\rm fra}}{M}\right)X_{\rm DD}\Big{[}[k-k_{\rm int}+\bar{k}]_{N},[l-l_{\rm int}+\bar{l}]_{M}\Big{]},$ (29)
consistently with [8, Proposition 2]. In addition, if $k_{\rm fra}=l_{\rm fra}=0$, then (29) becomes
$Y_{\rm DD}[k,l]$ $=\alpha X_{\rm DD}\Big{[}[k-k_{\rm int}]_{N},[l-l_{{\rm int}}]_{M}\Big{]}.$ (30)
Remark 4.
Refer to caption      Refer to caption      Refer to caption
Figure 2: $|\Phi[k,l]|$ vs $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ for Cases 1, 2, and 3, when $B=4$, $N=M=64$, and ideal shaping filters are employed.
The two-dimensional sequence $\alpha\Phi[k,l]$ represent the target response in the delay-Doppler domain. For a given $B$, $\Phi[k,l]$ is specified by the initial range and the range-rate of the target and satisfy $|\Phi[k,l]|\leq 1$; to get more insights on its structure, we discuss a toy example with $B=4$ and $N=M=64$. Fig. 2 shows $|\Phi[k,l]|$ versus $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ in different cases:
  1. 1.
    $k_{\rm fra}=l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=0$, $k_{\rm int}=32$, and $l_{0,{\rm int}}=\cdots=l_{3,{\rm int}}=42$;
  2. 2.
    $k_{\rm fra}=l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=0.5$, $k_{\rm int}=32$, and $l_{0,{\rm int}}=\cdots=l_{3,{\rm int}}=42$;
  3. 3.
    $k_{\rm fra}=l_{0,{\rm fra}}=\cdots=l_{B-1,{\rm fra}}=0.5$, $k_{\rm int}=50$, $l_{0,{\rm int}}=42$, $l_{1,{\rm int}}=43$, $l_{2,{\rm int}}=44$, and $l_{3,{\rm int}}=45$.
In Case 1, the range of the target is not changing in the OTFS frame; also, the range and range-rate of the target are multiple of $R_{\rm r}$ and $R_{\rm rr}$, respectively. In this case, $|\Phi[k,l]|$ only contains one non-zero entry, whose location depends upon the range and range-rate of the target, in keeping with the analytical model in (30). By comparing Cases 1 and 2, it is seen that a fractional range and a fractional range-rate cause a spread of $|\Phi[k,l]|$ along both domains, in keeping with (29). By comparing Cases 2 and 3, it seen that a variation of the target range every $N/B$ symbol intervals causes an additional spread of $|\Phi[k,l]|$ in both domains: this is the novel result established by Proposition 2. Finally, notice that most elements of $|\Phi[k,l]|$ are much smaller than one, in keeping with Remark 2; hence, the response of a high-speed target still has a sparse structure in the delay-Doppler domain, determined entirely by the initial range and range-rate. ∎
III-B Vectorized model
The samples $\{Y_{\rm DD}[k,l]\}$ can be organized into the vector $\bm{y}\in\mathbb{C}^{NM}$, where $(\bm{y})_{lN+k+1}=Y_{\rm DD}[k,l]$, for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$. In the following, we refer to $\bm{y}$ as the target echo. From (26a) we have
$\bm{y}=\alpha\bm{\Phi}\bm{x},$ (31)
where $\bm{\Phi}$ is an $NM\times NM$ matrix with $(\bm{\Phi})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Phi\big{[}[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\big{]}$ and $\bm{x}$ is an $NM$-dimensional vector with $(\bm{x})_{lN+k+1}=X_{\rm DD}[k,l]$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$. Alternatively, from (26b) we have
$\bm{y}=\alpha\bm{X}\bm{\phi},$ (32)
where $\bm{X}$ is an $NM\times NM$ matrix with $(\bm{X})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=X_{\rm DD}\big{[}[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\big{]}$ and $\bm{\phi}$ is an $NM$-dimensional vector with $(\bm{\phi})_{lN+k+1}=\Phi[k,l]$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$.
In the following, we refer to
$\bm{e}(d,v,B)=\bm{\Phi}(d,v,B)\bm{x}=\bm{X}\bm{\phi}(d,v,B)$ (33)
as the normalized target echo, where we have made explicit the dependence upon the initial range $d$ of the target, the range-rate $v$ of the target, and the parameter $B$.
IV Rectangular shaping filters
Assume here $g_{\rm tx}(t)=g_{\rm rx}(t)=\sqrt{1/T}\mathrm{rect}(t/T)$. Then, $\gamma(\tau,\nu)=0$ for $|\tau|\geq T$; also, for large $M$, recall that [8]
$\gamma(\tau,\nu)=\frac{T}{M}\sum_{p=0}^{M-1}g_{\rm tx}\left(\frac{pT}{M}\right)g^{*}_{\rm rx}\left(\frac{pT}{M}-\tau\right)\mathrm{e}^{-\mathrm{i}2\pi\nu\left(\frac{pT}{M}-\tau\right)}\\ =\begin{cases}\displaystyle\frac{1}{M}\sum_{p=0}^{\lceil M+\tau M/T\rceil-1}\!\!\mathrm{e}^{-\mathrm{i}2\pi\nu T\left(\frac{p}{M}-\frac{\tau}{T}\right)},&\tau\in(-T,0),\\ \displaystyle\frac{1}{M}\sum_{p=\lceil\tau M/T\rceil}^{M-1}\!\!\mathrm{e}^{-\mathrm{i}2\pi\nu T\left(\frac{p}{M}-\frac{\tau}{T}\right)},&\tau\in[0,T).\end{cases}$ (34)
To proceed, notice that (16) becomes here
$Y_{\rm TF}[n,m]$ $=H_{n,m}[n,m]X_{\rm TF}[n,m]$
$\quad+\sum_{m^{\prime}=0,\,m^{\prime}\neq m}^{M-1}H_{n,m}[n,m^{\prime}]X_{\rm TF}[n,m^{\prime}]$
$\quad+\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n-1,m^{\prime}]X_{\rm TF}[n-1,m^{\prime}],$ (35)
for $n=1,\ldots,N-1$ and $m=0,\ldots,M-1$, where the second and third terms on the right hand side are the inter carrier interference (ICI) and inter symbol interference (ISI) [8], respectively. For $n=0$ no ISI is present, as a guard interval is inserted between consecutive OTFS frames, and (16) becomes
$Y_{\rm TF}[0,m]$ $=H_{0,m}[0,m]X_{\rm TF}[0,m]$
$\quad+\sum_{m^{\prime}=0,\,m^{\prime}\neq m}^{M-1}H_{0,m}[0,m^{\prime}]X_{\rm TF}[0,m^{\prime}],$ (36)
for $m=0,\ldots,M-1$. Then, the following proposition establishes a connection between $\{Y_{\rm DD}[k,l]\}$ and $\{X_{\rm DD}[k,l]\}$ that generalizes the result in [8, Theorem 2] to the case where the target range is not a multiple of $R_{\rm r}$ and can change every $N/B$ symbol intervals; the proof is in the Appendix.
Proposition 3.
Assume rectangular shaping filters. Then, under the $B$-fold stop-and-go approximation on the range variation over the OFTS frame, we have
$Y_{\rm DD}[k,l]=\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\!X_{\rm DD}[k^{\prime},l^{\prime}]\Psi_{k^{\prime},l^{\prime}}[k-k^{\prime},l-l^{\prime}],$ (37)
for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$, where
$\Psi_{k^{\prime},l^{\prime}}[k,l]$ $=\frac{1}{B}\sum_{b=0}^{B-1}\Psi_{b,k^{\prime},l^{\prime}}[k,l],$ (38)
while
$\Psi_{b,k^{\prime},l^{\prime}}[k,l]=\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}}{M}}$
×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k}{N}-\frac{vT}{\lambda}\right)$
×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right),$ (39)
if $10\leq l^{\prime}\leq\big{\lceil}M-\frac{r\left(bT_{B}\right)M}{cT}\big{\rceil}-$ 1, and
$\Psi_{b,k^{\prime},l^{\prime}}[k,l]=\mathrm{e}^{-\mathrm{i}2\pi\frac{k^{\prime}}{N}}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}-M}{M}}$
×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\left[\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k}{N}-\frac{vT}{\lambda}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]$
×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l+M}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right),$ (40)
if $\big{\lceil}M-\frac{r\left(bT_{B}\right)M}{cT}\big{\rceil}\leq l^{\prime}\leq M-1$.
The following remarks on Proposition 3 are now in order.
Remark 5.
The summations in (37) cannot be regarded as a two-dimensial convolution, since $\Psi_{k^{\prime},l^{\prime}}[k,l]$ depends upon $k^{\prime}$ and $l^{\prime}$: this is in sharp contrast with the result obtained for ideal shaping filters in Remark 1. Notice in passing that, for fixed $k^{\prime}$ and $l^{\prime}$, $\Psi_{k^{\prime},l^{\prime}}[k,l]$ is a periodic function of $k$ and $l$ with period $N$ and $M$, respectively, while, for fixed $k$ and $l$, it is a periodic function of $k^{\prime}$ with period $N$. ∎
Remark 6.
For large $N/B$ and $M$, the summations in (37) only contain few dominant dominant terms; indeed, from the properties of the Dirichlet function discussed in Sec. III-A, we have $|\Psi_{k^{\prime},l^{\prime}}[k,l]|\approx 0$ if $[k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $[l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$, for $b=0,\ldots,B-1$. This result is similar to that obtained for ideal shaping filters in Remark 2. ∎
Remark 7.
Proposition 3 subsumes [8, Theorem 2] as a special case. To verify this, use (27) to recast (37) as
$Y_{{\rm DD}}[k,l]=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{k^{\prime}=0}^{N-1}$
×$\quad\times\Bigg{[}\sum_{l^{\prime}=0}^{M-1-l_{b,{\rm int}}-\lfloor l_{b,{\rm fra}}\rfloor}X_{\rm DD}[k^{\prime},l^{\prime}]\Psi_{b,k^{\prime},l^{\prime}}[k-k^{\prime},l-l^{\prime}]$
$\quad+\sum_{l^{\prime}=M-l_{b,{\rm int}}-\lfloor l_{b,{\rm fra}}\rfloor}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\Psi_{b,k^{\prime},l^{\prime}}[k-k^{\prime},l-l^{\prime}]\Bigg{]}$
$=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{\bar{k}=k_{\rm int}-k}^{N+k_{\rm int}-k-1}$
×$\quad\times\Bigg{[}\sum_{\bar{l}_{b}=l_{b,{\rm int}}-l}^{M-l-1-\lfloor l_{b,{\rm fra}}\rfloor}\Psi_{b,k^{\prime},l^{\prime}}[k_{\rm int}-\bar{k},l_{b,{\rm int}}-\bar{l}_{b}]$
×$\quad\times X_{\rm DD}\Big{[}[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big{]}$
$\quad+\sum_{\bar{l}_{b}=-l-\lfloor l_{b,{\rm fra}}\rfloor}^{l_{b,{\rm int}}-l-1}\Psi_{b,k^{\prime},l^{\prime}}[k_{\rm int}-\bar{k},l_{b,{\rm int}}-\bar{l}_{b}-M]$
×$\quad\times X_{\rm DD}\Big{[}[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big{]}\Bigg{]}$
$=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{\bar{k}=-\lfloor N/2\rfloor}^{\lceil N/2\rceil-1}\sum_{\bar{l}_{b}=-l-\lfloor l_{b,{\rm fra}}\rfloor}^{M-l-1-\lfloor l_{b,{\rm fra}}\rfloor}$
×$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l_{b,{\rm int}}+l_{b,{\rm fra}}}{M}}\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l-l_{b,{\rm int}}+\bar{l}_{b}}{M}}$
×$\quad\times\mathrm{e}^{\mathrm{i}2\pi\left(\frac{\bar{k}+k_{\rm fra}}{B}\right)b}A_{b,\bar{k},\bar{l}_{b}}[k,l]\mathcal{D}_{M}\!\left(\frac{\bar{l}_{b}+l_{b,{\rm fra}}}{M}\right)$
×$\quad\times X_{\rm DD}\Big{[}[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big{]},$ (41)
where $A_{b,\bar{k},\bar{l}_{b}}[k,l]$ is equal to
$\mathcal{D}_{\frac{N}{B}}\!\left(-\frac{\bar{k}+k_{\rm fra}}{N}\right),$ (42)
if $l_{b,{\rm int}}-l\leq\bar{l}_{b}\leq M-l-1-\lfloor l_{b,{\rm fra}}\rfloor$, and to
$\left[\mathcal{D}_{\frac{N}{B}}\!\left(-\frac{\bar{k}+k_{\rm fra}}{N}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]\mathrm{e}^{-\mathrm{i}2\pi\frac{[k-k_{\rm int}+\bar{k}]_{N}}{N}},$ (43)
if $-l-\lfloor l_{b,{\rm fra}}\rfloor\leq\bar{l}_{b}\leq l_{b,{\rm int}}-l-1$. If $l_{b,{\rm int}}=l_{{\rm int}}$ and $l_{b,{\rm fra}}=0$ for $b=0,\ldots,B-1$, then (41) becomes
$Y_{{\rm DD}}[k,l]=\alpha\sum_{\bar{k}=-\lfloor N/2\rfloor}^{\lceil N/2\rceil-1}\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l_{{\rm int}}}{M}}\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l-l_{{\rm int}}}{M}}\\ \times A_{\bar{k}}[k,l]X_{\rm DD}\Big{[}[k-k_{\rm int}+\bar{k}]_{N},[l-l_{{\rm int}}]_{M}\Big{]},$ (44)
where $A_{\bar{k}}[k,l]$ is equal to
$\mathcal{D}_{N}\!\left(-\frac{\bar{k}+k_{\rm fra}}{N}\right),$ (45)
if $l_{{\rm int}}\leq l\leq M-1$, and to
$\left[\mathcal{D}_{N}\!\left(-\frac{\bar{k}+k_{\rm fra}}{N}\right)-\frac{1}{N}\right]\mathrm{e}^{-\mathrm{i}2\pi\frac{[k-k_{\rm int}+\bar{k}]_{N}}{N}},$ (46)
if $10\leq l\leq l_{{\rm int}}-$ 1, that is consistent with [8, Theorem 2]. ∎
Remark 8.
Following the discussion in Remark 4, the two-dimensional sequence $\alpha\Psi_{k^{\prime},l^{\prime}}[k,l]$ can be regarded as the target response in the delay-Doppler domain for an input symbol located on the delay-Doppler grid point indexed by $k^{\prime}$ and $l^{\prime}$. We underline that the target response depends now upon the location of the input symbol in the delay-Doppler plane, which is in sharp contrast to the result previously obtained for ideal shaping filters. For a given $B$, the sequence $\Psi_{k^{\prime},l^{\prime}}[k,l]$ in (38) is entirely specified by the initial range and the range-rate of the target and satisfy $|\Psi_{k^{\prime},l^{\prime}}[k,l]|\leq 1$: this latter result is instead similar to that for ideal shaping filters. ∎
IV-A Vectorized model
As in Sec III-B, the samples $\{Y_{\rm DD}[k,l]\}$ can be organized into the vector $\bm{y}\in\mathbb{C}^{NM}$ representing the target echo, where $(\bm{y})_{lN+k+1}=Y_{\rm DD}[k,l]$, for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$. From (37), we have
$\bm{y}=\alpha\bm{\Psi}\bm{x},$ (47)
where $\bm{\Psi}$ is an $(NM\times NM)$-dimensional matrix with $(\bm{\Psi})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Psi_{k^{\prime},l^{\prime}}\big{[}k-k^{\prime},l-l^{\prime}\big{]}$. Interestingly, the models in (31) and (47) present the same structure; however, the involved matrices $\bm{\Phi}$ and $\bm{\Psi}$ are different. For future reference, we also define the normalized target echo as
$\bm{e}(d,v,B)=\bm{\Psi}(d,v,B)\bm{x},$ (48)
where the dependence upon the initial range $d$ of the target, the range-rate $v$ of the target, and $B$ has been made explicit.
IV-B Approximated model
We derive here an approximated model: the goal is to obtain an input-output relationship with a structure similar to the one in (32), which will turn out useful in Sec. V-A.
Notice first that (39) and (40) differ by the term
$e^{-\mathrm{i}2\pi\frac{k^{\prime}}{N}}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{d}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}-M}{M}}\\ \times\frac{B}{N}\mathcal{D}_{M}\!\left(-\left(\frac{l+M}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right)\mathbbm{1}_{\{b=0\}},$ (49)
whose magnitude tends to zero for large $N/B$. Upon neglecting the above term in (40) and upon ignoring the fractional part of the target range, then (41) can be approximated as
$Y_{{\rm DD}}[k,l]\approx\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{NM}l}\Theta_{k^{\prime},l^{\prime}}[k,l]$
×$\quad\times\frac{1}{B}\sum_{b=0}^{B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k^{\prime}-k_{\rm int}-k_{\rm fra}}{B}\right)b}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k^{\prime}-k_{\rm int}-k_{\rm fra}}{N}\right)$
×$\quad\times\mathcal{D}_{M}\!\left(-\frac{l^{\prime}-l_{b,{\rm int}}}{M}\right)X_{\rm DD}\Big{[}[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\Big{]},$ (50)
where
$\Theta_{k^{\prime},l^{\prime}}[k,l]=\left\{\begin{matrix}1&l^{\prime}\leq l\leq M-1,\\ \mathrm{e}^{-\mathrm{i}2\pi\frac{[k-k^{\prime}]_{N}}{N}}&0\leq l\leq l^{\prime}-1.\\ \end{matrix}\right.$ (51)
Based on (50), we can now write
$\bm{y}\approx\alpha\big{(}\bm{X}\odot\bm{\Theta}\odot\bm{\Lambda}(v)\big{)}\bm{\phi}(d,v,B),$ (52)
where $\bm{X}$ and $\bm{\phi}(d,v,B)$ are the same quantities in (32), while $\bm{\Theta}$ and $\bm{\Lambda}(v)$ are $NM\times NM$ matrices with $(\bm{\Theta})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Theta_{k^{\prime},l^{\prime}}[k,l]$ and $\big{(}\bm{\Lambda}(v)\big{)}_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{NM}l}=\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda M}l}$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$.
To proceed, notice that $\{\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda M}l}\}_{l=-\infty}^{\infty}$ is a complex exponential whose numerical frequency depends upon the target range-rate. Since $\frac{v_{\max}T}{\lambda}<1$ and the samples $\{\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda M}l}\}_{l=0}^{M-1}$ present an overall phase variation less than $2\pi$, we propose to approximate $\bm{\Lambda}(v)$ by $\bm{\Lambda}(v_{\max}/2)$. Hence, we finally obtain
$\bm{y}\approx\alpha\bm{\Xi}\bm{\phi}(d,v,B),$ (53)
where $\bm{\Xi}=\bm{X}\odot\bm{\Theta}\odot\bm{\Lambda}(v_{\max}/2)$. By comparing (32) with (53), it is seen that the symbol matrix $\bm{X}$ in (32) is now replaced in (52) by the modified symbol matrix $\bm{\Xi}$ to account for the presence of rectangular filters. To compare (47) and (53), we compute the correspoding normalized root mean square error (NRMSE), namely, $(\mathrm{E}\left[\|\bm{\Xi}\bm{\phi}(d,v,B)-\bm{\Psi}(d,v,B)\bm{x}\|^{2}\right]/\mathrm{E}\left[\|\bm{\Psi}(d,v,B)\bm{x}\|^{2}\right])^{1/2}$, where the statistical expectation is with respect to $\bm{x}$, $d$, and $v$. For the system setup considered in Sec. VI, we found an NRMSE of about $0.07$ for any $B$, thus confirming the goodness of the approximated model in (53).
V Estimation of the target parameters
Upon accounting for the presence of the additive noise, after OTFS demodulation the received signal can be written as
$\bm{z}=\bm{y}+\bm{\omega}\in\mathbb{C}^{NM},$ (54)
where $\bm{y}$ is the target echo and $\bm{\omega}$ is a complex circularly-symmetric Gaussian vector with covariance matrix $\sigma_{\omega}^{2}\bm{I}_{NM}$. Based on $\bm{z}$ and on the echo model $\bm{y}=\alpha\bm{e}(d,v,B)$ provided in Secs. III-B and IV-A, the radar receiver is faced with the problem of estimating the unknown target parameters, namely, $d$, $v$, and $\alpha$. We resort here to an ML estimation procedure; for a given $B$, the corresponding estimates are [43]
$\big{[}\hat{d}(B);\,\hat{v}(B)\big{]}$ $=\operatorname*{\arg\max}_{\begin{subarray}{c}[d;\,v]\in\mathcal{S}\end{subarray}}\frac{\big{|}\bm{e}(d,v,B)^{\mathsf{H}}\bm{z}\big{|}^{2}}{\|\bm{e}(d,v,B)\|^{2}},$ (55a)
$\hat{\alpha}(B)$ $=\frac{\bm{e}\big{(}\hat{d}(B),\hat{v}(B),B\big{)}^{\mathsf{H}}\bm{z}}{\big{\|}\bm{e}\big{(}\hat{d}(B),\hat{v}(B),B\big{)}\big{\|}^{2}},$ (55b)
where $\mathcal{S}=[0,r_{\max}]\times[0,v_{\max}]$. In practice, the continuous set $\mathcal{S}$ is replaced by a uniform rectangular grid $\mathcal{G}$ containing $G_{\rm r}=\lfloor r_{\max}/\Sigma_{\rm r}\rfloor$ and $G_{\rm rr}=\lfloor v_{\max}/\Sigma_{\rm rr}\rfloor$ points along the range and range-rate dimensions, respectively, where $\Sigma_{\rm r}\leq R_{\rm r}$ and $\Sigma_{\rm rr}\leq R_{\rm rr}$ are the corresponding point spacing. Since the cardinality of $\mathcal{G}$ and the size of $\bm{z}$ scale with $MN$, the computation of (55a) may become demanding for large $M$ and $N$, which is often the case in OTFS-based systems. To overcome this problem, we propose a two-step estimation procedure which operates as follows.
  1. 1.
    As shown in Sec. V-A, we first exploit the signal sparsity to compute a coarse estimate of the target response and, as a by-product, coarse estimates of the initial range and range-rate, denoted as $\tilde{d}(B)$ and $\tilde{v}(B)$, respectively.
  2. 2.
    Then, we restrict the search set in (55a) to a neighborhood of the above coarse estimates; in particular, we consider a uniform rectangular grid $\tilde{\mathcal{G}}$ that is centered at $\big{[}\tilde{d}(B);\,\tilde{v}(B)\big{]}$ and has $\tilde{G}_{\rm r}$ and $\tilde{G}_{\rm rr}$ points along the range and the range-rate dimension spaced of $\Sigma_{\rm r}$ and $\Sigma_{\rm rr}$, respectively, with $\tilde{G}_{\rm r}\ll G_{\rm r}$ and $\tilde{G}_{\rm rr}\ll G_{\rm rr}$.
V-A Coarse estimation of the initial range and range-rate
Consider first ideal shaping filters. Then, from (32) the received signal can be expended as $\bm{z}=\bm{X}\bm{f}+\bm{\omega}$, where $\bm{f}=\alpha\bm{\phi}$ is the target response vector. [^7] For brevity, the dependence upon $d$, $v$, and $B$ is kept concealed here. To proceed, notice that $\bm{f}$ can be partitioned into $M$ sub-vectors (blocks) of length $N$, namely, $\bm{f}=(\bm{f}_{0};\cdots;\bm{f}_{M-1})$, where $\bm{f}_{m}=\big{(}\alpha\Phi[0,m];\ldots;\alpha\Phi[N-1,m]\big{)}$; according to Remark 2, $\bm{f}$ is a block sparse vector for large $N/B$ and $M$. Notice next that the size of $\bm{z}$ can be reduced by employing a $NK\times NM$ selection matrix $\bm{S}$ whose rows are formed by any $NK$ distinct rows of $\bm{I}_{NM}$; the parameter $K\in\{1,\ldots,M\}$ rules the implementation complexity of the coarse estimator, and in practice we can select $K\ll M$. The compressed measurement $\bm{z}_{s}=\bm{S}\bm{z}\in\mathbb{C}^{NK}$ can be expanded as
$\bm{z}_{s}=\bm{X}_{s}\bm{f}+\bm{\omega}_{s}=\sum_{m=0}^{M-1}\bm{X}_{s,m}\bm{f}_{m}+\bm{\omega}_{s},$ (56)
where $\bm{\omega}_{s}=\bm{S}\bm{\omega}$, $\bm{X}_{s}=(\bm{X}_{s,0}\,\cdots\,\bm{X}_{s,M-1})=\bm{S}\bm{X}$, and $\bm{X}_{s,0},\ldots,\bm{X}_{s,M-1}\in\mathbb{C}^{NK\times N}$.
Based on (56), an estimate of the target response vector, say $\tilde{\bm{f}}$, can be computed by resorting to the BOMP algorithm [35]. Let $\bm{z}^{(1)}_{s}=\bm{z}_{s}$ and $\mathcal{I}^{(1)}=\{0,\ldots,M-1\}$; then, at the $\ell$-th iteration, for $\ell\geq 1$, the index of the block in the dictionary matrix $\bm{X}_{s}$ best matched to $\bm{z}^{(\ell)}_{s}$ is first identified as follows
$i^{(\ell)}=\operatorname*{\arg\max}_{i\in\mathcal{I}^{(\ell)}}\Big{\|}\bm{X}_{s,i}^{\mathsf{H}}\bm{z}^{(\ell)}_{s}\Big{\|}^{2};$ (57)
then, an estimate of the sub-vectors $\bm{f}_{i^{(1)}},\ldots,\bm{f}_{i^{(\ell)}}$ corresponding to the currently selected blocks of $\bm{X}_{s}$ is computed by solving a least-square problem, namely,
$\big{(}\tilde{\bm{f}}_{i^{(1)}}\cdots\tilde{\bm{f}}_{i^{(\ell)}}\big{)}=\operatorname*{\arg\max}_{(\bm{f}_{i^{(1)}}\cdots\bm{f}_{i^{(\ell)}})\in\mathbb{C}^{N\times\ell}}\Big{\|}\bm{z}_{s}-\sum_{q=1}^{\ell}\bm{X}_{s,i^{(q)}}\bm{f}_{i^{(q)}}\Big{\|}^{2};$ (58)
finally, the residual vector and the search set are updated as $\bm{z}^{(\ell+1)}_{s}=\bm{z}^{(\ell)}_{s}-\sum_{q=1}^{\ell}\bm{X}_{s,i^{(q)}}\tilde{\bm{f}}_{i^{(q)}}$ and $\mathcal{I}^{(\ell+1)}=\mathcal{I}^{(\ell)}\setminus\{i^{(\ell)}\}$, respectively, and the procedure is stopped when $\ell=K$ or $\|\bm{z}^{(\ell+1)}_{s}\|^{2}<\epsilon NK\sigma_{\omega}^{2}$, with $\epsilon$ being a positive threshold. If the procedure ends in $\tilde{\ell}$ iterations, an estimate $\tilde{\bm{f}}$ of the target response vector is obtained from the indexes $i^{(1)},\ldots,i^{(\tilde{\ell})}$ and the sub-vectors $\tilde{\bm{f}}_{i^{(1)}},\ldots,\tilde{\bm{f}}_{i^{(\tilde{\ell})}}$. At this point, the entries of $\tilde{\bm{f}}$ are arranged into the target response sequence $\tilde{F}[k,l]=(\tilde{\bm{f}})_{lN+k+1}$, for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$, which provides an estimate of $\alpha\Phi[k,l]$; finally, coarse estimates of the initial range and range-rate of target are obtained from the location of the maximum element of $|\tilde{F}[k,l]|$, in keeping with the discussion provided in Remark 4.
When a rectangular shaping filter is employed, from (53), the received signal can be approximately expanded as $\bm{z}\approx\alpha\bm{\Xi}\bm{\phi}+\bm{\omega}$; accordingly, the above procedure can still be used after replacing $\bm{X}$ by the modified symbol matrix $\bm{\Xi}$.
V-B Extension to multiple targets
The above estimation procedure can be combined with the CLEAN algorithm in [36, 37, 38, 39, 15] to handle multiple targets at low complexity. The CLEAN algorithm iteratively reconstructs one target echo at the time, after removing the interference caused by the previously-estimated target echoes from the data. More specifically, assume that $P>1$ targets are present in the inspected region and denote by $\hat{\bm{y}}_{p}(B)=\hat{\alpha}_{p}(B)\bm{e}\big{(}\hat{d}_{p}(B),\hat{v}_{p}(B),B\big{)}$ the reconstructed target echo in the $p$-th iteration, where $\hat{d}_{p}(B)$, $\hat{v}_{p}(B)$, and $\hat{\alpha}_{p}(B)$ are the estimated initial range, range-rate, and amplitude, respectively; then, the proposed two-step estimation procedure is iterated on the cleaned data $\bm{z}-\sum_{i=1}^{p-1}\hat{\bm{y}}_{i}(B)$, for $p=1,\ldots,P$.
Table I: System parameters
$c/\lambda$ Carrier frequency 4 GHz
$\Delta$ Subcarrier spacing 15 kHz
$M$ Number of subcarriers 512
$N$ Number of symbol intervals 128
$R_{\rm r}$ Range resolution 19.5 m
$R_{\rm rr}$ Range-rate resolution 4.4 m/s
$[\bar{r}_{\rm min},\,\bar{r}_{\rm max}]$ Inspected range interval [20, 39] km
$[\bar{v}_{\rm min},\,\bar{v}_{\rm max}]$ Inspected range-rate interval [15, 16] km/s
VI Numerical analysis
To assess the system performance, we consider the parameters in Table I. The maximum variation of the target range during the OTFS frame is $\bar{v}_{\rm max}NT=136$ m, corresponding to about $7$ times the range resolution $R_{\rm r}$; the initial range $d$ and the range-rate $v$ are randomly generated, while a Swerling I fluctuation model is considered for the amplitude $\alpha$. The information symbols $\{X_{\rm DD}[k,l]\}$ are independently drawn from a unit-energy $4$-QAM constellation. The entries of the observed vector $\bm{z}$ are computed according to Eqs. (11), (14), and (18). Instead, the radar receiver is designed according to the echo model proposed in Secs. III and IV for ideal and rectangular shaping filters, respectively, that relies upon the $B$-fold stop-and-go approximation. In the implementation of the proposed two-step estimator, we set $\Sigma_{\rm r}=R_{\rm r}/100$, $\Sigma_{\rm rr}=R_{\rm rr}/100$, $\tilde{G}_{\rm r}=401$, $\tilde{G}_{\rm rr}=401$, $K=4P$, and $\epsilon=1$.
For future reference, we define the signal-to-noise ratio (SNR) of the target as $\mathrm{SNR}=\mathrm{E}[|\alpha|^{2}]NM/\sigma_{\omega}^{2}$. Also, let $d$, $v$, and $\alpha$ be the true values of the initial range, range-rate, and amplitude of the target, and let $\hat{d}$, $\hat{v}$, and $\hat{\alpha}$ be the corresponding estimates; then, we define the root mean square error (RMSE) in the estimation of the initial range, the RMSE in the estimation of the range-rate, and the NRMSE in the estimation of the amplitude as $\mathrm{RMSE}_{\rm r}=(\mathrm{E}[|d-\hat{d}|^{2}])^{1/2}$, $\mathrm{RMSE}_{\rm rr}=(\mathrm{E}[|v-\hat{v}|^{2}])^{1/2}$, and $\mathrm{NRMSE}_{\rm\alpha}=(\mathrm{E}\left[|\alpha-\hat{\alpha}|^{2}\right]/\mathrm{E}\left[|\alpha|^{2}\right])^{1/2}$, respectively. Next, we discuss two examples with one and multiple targets, respectively.
VI-A Single target
Refer to caption Refer to caption Refer to caption
Figure 3: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $B=1,2,4,8,16,32,64,128$ for $\mathrm{SNR}=10,20,30$ dB, when a high- or low-speed target is present and ideal shaping filters are employed. Both the ML and the proposed estimator are considered for comparison.
Refer to caption Refer to caption Refer to caption
Figure 4: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $\mathrm{SNR}$ for $B=4,8,16$, when a high-speed target is present and the proposed estimator is employed. Both ideal and rectangular shaping filters are considered.
Refer to caption Refer to caption Refer to caption
Figure 5: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $\bar{v}_{\rm min}$, when rectangular filters are used, $\mathrm{SNR}=15$ dB, and $\bar{v}_{\rm max}=\bar{v}_{\rm min}+1$ km/s. The proposed estimator is implemented for $B=1,2,4,8,16$. For comparison, the performance of the estimators presented in [20] and [28] is also evaluated.
[…]
Next, we consider the use of rectangular shaping filters. Fig. 4 shows the behavior of $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\rm\alpha}$ as functions of $\mathrm{SNR}$ when the proposed estimator is employed and $B=4,8,16$. For comparison, results obtained with ideal shaping filters are also shown. Notably, the more practical rectangular filters achieve estimation performance that closely approximates that of the ideal filters.
Finally, we compare the performance of the proposed estimator with the methods in [20] and [28], which do not explicitly account for range migration within the OTFS frame. To evaluate performance across varying mobility scenarios, we set $\bar{v}_{\rm max}=\bar{v}_{\rm min}+1$ km/s and vary $\bar{v}_{\rm min}$ from 0 to 15 km/s. Fig. 5 plots the metrics $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\alpha}$ versus $\bar{v}_{\rm min}$ for rectangular shaping filters and $\mathrm{SNR}=15$ dB. The performance of the estimators in [20] and [28] deteriorates rapidly when $\bar{v}_{\rm min}$ exceeds 1 km/s. In this regime, the maximum range variation during the OTFS frame, given by $\bar{v}_{\rm max}NT$, becomes comparable to or exceeds the range resolution $R_{\rm r}$, resulting in a mismatch between the assumed and actual echo models. In contrast, the proposed estimator maintains robust performance across a broader range of $\bar{v}_{\rm min}$ by increasing $B$, thereby enabling a finer stop-and-go approximation in (7). Performance degradation only occurs when the range variation over $N/B$ symbol intervals, i.e., $\bar{v}_{\rm max}NT/B$, becomes comparable to or exceeds $R_{\rm r}$. Notably, with $B=16$, the proposed estimator delivers consistent performance across all tested values of $\bar{v}_{\rm min}$, in keeping with the findings in Fig. 3. Furthermore, as $\bar{v}_{\rm min}\rightarrow 0$, its performance aligns with that of the existing methods, confirming its effectiveness and practical relevance even for low-speed targets.
VI-B Multiple targets
Refer to caption Refer to caption Refer to caption
Figure 6: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $\mathrm{SNR}_{2}$, when rectangular filters are used, $P=3$ high-speed targets are present, $B=16$, and $\mathrm{SNR}_{1}/\mathrm{SNR}_{2}=\mathrm{SNR}_{2}/\mathrm{SNR}_{3}=10$. The performance of the proposed CLEAN-based estimator and the corresponding single-target performance bound are reported.
Refer to caption Refer to caption Refer to caption
Figure 7: $|\bm{F}[k,l]|/\sigma_{\omega}$ vs $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ at the first (top), second (middle), and third (bottom) iteration of the CLEAN procedure, when rectangular filters are employed, $P=3$ high-speed targets are present, $B=16$, $\mathrm{SNR}_{1}=30$ dB, $\mathrm{SNR}_{2}=20$ dB, and $\mathrm{SNR}_{3}=10$ dB.
We use here rectangular shaping filters and consider $P=3$ targets. Let $\mathrm{SNR}_{p}$ be the SNR of Target $p$ for $p=1,\ldots,P$. For each target, Fig. 6 reports the metrics $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\rm\alpha}$ versus $\mathrm{SNR}_{2}$, when $B=16$, $\mathrm{SNR}_{1}/\mathrm{SNR}_{2}=\mathrm{SNR}_{2}/\mathrm{SNR}_{3}=10$, and the proposed CLEAN-based estimator is employed. For comparison, we also plot the achievable performance when only a single target is present and the ML estimator in (55) is employed: we refer to it as the single-target bound. The proposed procedure is able to effectively estimate the parameters of each target, providing a performance close to the single-target bound. For a single instance and $\mathrm{SNR}_{2}=20$ dB, Fig. 7 shows $|\tilde{F}[k,l]|/\sigma_{\omega}$ versus $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ in each iteration of the CLEAN procedure, where $\tilde{F}[k,l]$ is the target response sequence recovered by the BOMP algorithm. It is seen that the echo of Target 1 (the strongest one) is recovered in Iteration 1 and subtracted out from the observed data; then, the echoes of Targets 2 and 3 are recovered in Iterations 2 and 3, respectively.
VII Conclusions
In this work, we have investigated the use of OTFS communication signals for estimating the parameters of space targets experiencing range migration, which may occur in LEO satellite applications. Our major technical contribution is the derivation of a novel model for the target echo, assuming that the target range changes $B$ times within an OTFS frame; in particular, we have shown that range migration causes a predictable spread in the sparse target response in the delay-Doppler domain. Additionally, we have developed an approximate implementation of the maximum likelihood estimator for the target’s initial range, range-rate, and amplitude, which avoids exhaustive search in the delay-Doppler domain. If combined with the CLEAN algorithm, the proposed procedure can also handle multiple targets. A significant finding is that the estimation accuracy for high-speed targets is similar to that for low-speed targets when range migration is properly accounted for at the design stage.
Future studies should consider that the radar receiver may not have perfect knowledge of the emitted symbols. In such cases, the use of dedicated radar symbols may be considered, and the optimal resource allocation between sensing and communication tasks should be studied. Furthermore, using multiple coordinated LEO satellites for signal transmission and/or reception may allow for estimating the initial position and velocity vectors of the target rather than only its initial range and range-rate. Finally, the use of multiple OTFS frames of different sizes may provide additional degrees of freedom and help resolve possible range and range-rate ambiguities.
We provide here the proof to Proposition 1. Upon plugging (35) and (36) in (18), we have
$Y_{\rm DD}[k,l]$ $=\frac{1}{\sqrt{NM}}\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)}$
×$\quad\times\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n,m^{\prime}]X_{\rm TF}[n,m^{\prime}]$
$\quad+\frac{1}{\sqrt{NM}}\sum_{n=1}^{N-1}\sum_{m=0}^{M-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)}$
×$\quad\times\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n-1,m^{\prime}]X_{\rm TF}[n-1,m^{\prime}]$
$=Y_{\rm DD}^{\rm ici}[k,l]+Y_{\rm DD}^{\rm isi}[k,l].$ (59)
As to the term $Y_{\rm DD}^{\rm ici}[k,l]$ in (59), we have:
$Y_{\rm DD}^{\rm ici}[k,l]$ $=\frac{1}{NM}\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)}\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n,m^{\prime}]$
×$\quad\times\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{\mathrm{i}2\pi\left(\frac{nk^{\prime}}{N}-\frac{m^{\prime}l^{\prime}}{M}\right)}$
$=\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\Pi_{k,l}^{\rm ici}[k^{\prime},l^{\prime}],$ (60)
where
$\Pi_{k,l}^{\rm ici}[k^{\prime},l^{\prime}]$ $=\frac{1}{NM}\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n,m^{\prime}]$
×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}\right)}e^{j2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}.$ (61)
$\Pi_{k,l}^{\rm ici}[k^{\prime},l^{\prime}]=\frac{\alpha}{NM^{2}}\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}\sum_{m^{\prime}=0}^{M-1}\left[\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}nT}\mathrm{e}^{-\mathrm{i}2\pi m^{\prime}\Delta\frac{r\left(nT\right)}{c}}\sum_{p=0}^{\big{\lceil}M-\frac{r\left(nT\right)M}{cT}\big{\rceil}-1}\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left(\frac{pT}{M}+\frac{r\left(nT\right)}{c}\right)}\right]\mathrm{e}^{\mathrm{i}2\pi n\left(\frac{k^{\prime}-k}{N}\right)}\mathrm{e}^{\mathrm{i}2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}$
$=\frac{\alpha}{N}\sum_{n=0}^{N-1}\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(nT\right)}{c}}\sum_{p=0}^{\big{\lceil}M-\frac{r\left(nT\right)M}{cT}\big{\rceil}-1}\left[\frac{1}{M}\sum_{m=0}^{M-1}\mathrm{e}^{-\mathrm{i}2\pi\frac{m}{M}\left(p+\frac{r\left(nT\right)}{c}\frac{M}{T}-l\right)}\right]\left[\frac{1}{M}\sum_{m^{\prime}=0}^{M-1}\mathrm{e}^{\mathrm{i}2\pi\frac{m^{\prime}}{M}(p-l^{\prime})}\right]\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{p}{M}}$
$=\sum_{b=0}^{B-1}\frac{\alpha}{N}\sum_{n=bN/B}^{(b+1)N/B-1}\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(nT\right)}{c}}\sum_{p=0}^{\big{\lceil}M-\frac{r\left(nT\right)M}{cT}\big{\rceil}-1}\mathcal{D}_{M}\!\left(\frac{p+\frac{r\left(nT\right)}{c}\frac{M}{T}-l}{M}\right)\delta\big{[}[p-l^{\prime}]_{M}\big{]}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{p}{M}}$
$=\frac{\alpha}{B}\sum_{b=0}^{B-1}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\tfrac{k-k^{\prime}}{N}-\tfrac{vT}{\lambda}\right)\sum_{p=0}^{\big{\lceil}M-\frac{r\left(bT_{B}\right)M}{cT}\big{\rceil}-1}\mathcal{D}_{M}\!\left(\tfrac{p+\frac{r\left(bT_{B}\right)}{c}\frac{M}{T}-l}{M}\right)\delta\big{[}[p-l^{\prime}]_{M}\big{]}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{p}{M}}.$ (62)
Upon expanding $\Pi_{k,l}^{\rm ici}[k^{\prime},l^{\prime}]$ as shown in (62) at the top of this page, we obtain
$Y_{{\rm DD}}^{\rm ici}[k,l]$ $=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{\big{\lceil}M-\frac{r\left(bT_{B}\right)M}{cT}\big{\rceil}-1}X_{\rm DD}[k^{\prime},l^{\prime}]$
×$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}}{M}}$
×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)$
×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (63)
As to the term $Y_{\rm DD}^{\rm isi}[k,l]$ in (59), we have:
$Y_{\rm DD}^{\rm isi}[k,l]$ $=\frac{1}{NM}\sum_{n=1}^{N-1}\sum_{m=0}^{M-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)}\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n-1,m^{\prime}]$
×$\quad\times\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{\mathrm{i}2\pi\left(\frac{(n-1)k^{\prime}}{N}-\frac{m^{\prime}l^{\prime}}{M}\right)}$
$=\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\Pi_{k,l}^{\rm isi}[k^{\prime},l^{\prime}]\mathrm{e}^{-\mathrm{i}2\pi\frac{k^{\prime}}{N}},$ (64)
where
$\Pi_{k,l}^{\rm isi}[k^{\prime},l^{\prime}]$ $=\frac{1}{NM}\sum_{n=1}^{N-1}\sum_{m=0}^{M-1}\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n-1,m^{\prime}]$
×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}\right)}e^{j2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}.$ (65)
$\Pi_{k,l}^{\rm isi}[k^{\prime},l^{\prime}]=\frac{\alpha}{NM^{2}}\!\sum_{n=1}^{N-1}\sum_{m=0}^{M-1}\sum_{m^{\prime}=0}^{M-1}\!\left[\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}nT}\mathrm{e}^{-\mathrm{i}2\pi m^{\prime}\Delta\frac{r\left((n-1)T\right)}{c}}\!\!\!\!\sum_{p=\left\lceil M-\frac{r\left((n-1)T\right)M}{cT}\right\rceil}^{M-1}\!\!\!\!\!\!\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left(\frac{pT}{M}-T+\frac{r\left((n-1)T\right)}{c}\right)}\!\right]\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}\right)}\!\mathrm{e}^{\mathrm{i}2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}$
$=\frac{\alpha}{N}\sum_{n=1}^{N-1}\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left((n-1)T\right)}{c}}\!\!\sum_{p=\left\lceil M-\frac{r\left((n-1)T\right)}{c}\frac{M}{T}\right\rceil}^{M-1}\!\!\left[\frac{1}{M}\sum_{m=0}^{M-1}\mathrm{e}^{-\mathrm{i}2\pi\frac{m}{M}\left(p-M+\frac{r\left((n-1)T\right)}{c}\frac{M}{T}-l\right)}\right]\left[\frac{1}{M}\sum_{m^{\prime}=0}^{M-1}\mathrm{e}^{\mathrm{i}2\pi\frac{m^{\prime}}{M}(p-M-l^{\prime})}\right]\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{p-M}{M}}$
$=\sum_{b=0}^{B-1}\frac{\alpha}{N}\sum_{\begin{subarray}{c}n=bN/B\\ n\neq 0\end{subarray}}^{(b+1)N/B-1}\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left((n-1)T\right)}{c}}\sum_{p=\left\lceil M-\frac{r\left((n-1)T\right)}{c}\frac{M}{T}\right\rceil}^{M-1}\mathcal{D}_{M}\!\left(\frac{p-M+\frac{r\left((n-1)T\right)}{c}\frac{M}{T}-l}{M}\right)\delta\big{[}[p-l^{\prime}]_{M}\big{]}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{p-M}{M}}$
$=\frac{\alpha}{B}\sum_{b=0}^{B-1}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\left[\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]$
×$\quad\times\sum_{p=\left\lceil M-\frac{r\left(bT_{B}\right)}{c}\frac{M}{T}\right\rceil}^{M-1}\mathcal{D}_{M}\!\left(\frac{p-M+\frac{r\left(bT_{B}\right)}{c}\frac{M}{T}-l}{M}\right)\delta\big{[}[p-M-l^{\prime}]_{M}\big{]}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{p-M}{M}}.$ (66)
Upon expanding $\Pi_{k,l}^{\rm isi}[k^{\prime},l^{\prime}]$ as shown in (66) at the top of the next page, we obtain
$Y_{{\rm DD}}^{\rm ici}[k,l]$ $=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=\big{\lceil}M-\frac{r\left(bT_{B}\right)M}{cT}\big{\rceil}}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{-\mathrm{i}2\pi\frac{k^{\prime}}{N}}$
×$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}-M}{M}}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}$
×$\quad\times\left[\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]$
×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}+M}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (67)
Finally, (37) follows from (63) and (67).

### v2

@@
 Finally, numerical examples are provided to evaluate the estimation performance.
-Index Terms:
-Communication-centric ISAC, opportunistic sensing, OTFS modulation, LEO satellites, space targets, high-speed targets, range migration, maximum likelihood estimation.
 I Introduction
@@
 OTFS-based transceivers exploit channel diversity across both time and frequency, thus providing better communication performance in doubly-dispersive fading channels as compared to schemes using orthogonal frequency division multiplexing [8, 9, 10].
-They are also suitable for radar sensing, as the underlying signal processing resembles that of pulsed-Doppler radars [11]; in particular, the range and range-rate [^1] In this study, range-rate refers to the radial component of velocity, i.e., the rate at which the distance between the object and the observer is changing.
-of prospective scattering objects can be inferred from the estimated delay-Doppler channel response [12].
+They are also suitable for radar sensing, as the underlying signal processing resembles that of pulsed-Doppler radars [11]; in particular, the range and range-rate of prospective scattering objects can be inferred from the estimated delay-Doppler channel response [12].
 These unique features enable OTFS-based LEO satellites to inherently support ISAC in various forms [13, 14].
 Unlike terrestrial stations, LEO-based stations have the capability to also cover aerial and space regions, and this expanded coverage opens up new application scenarios, such as real-time environmental monitoring, enhanced global navigation systems, and space exploration.
-Communication-centric ISAC solutions appear particularly promising in this context, as existing data signals can be opportunistically reused to add the sensing functions at a low cost, without altering the underlying communication protocol or affecting its performance [12, 15, 16, 17, 18]; this approach is indeed considered in this study.
+Communication-centric ISAC solutions appear particularly promising in this context, as existing data signals can be opportunistically reused to add the sensing functions at a low cost, without altering the underlying communication protocol or affecting its performance [12, 15, 16, 17, 18, 19]; this approach is indeed considered in this study.
 Several solutions for signal estimation and detection have been deployed for OTFS-based applications, which exploit the channel sparsity in the delay-Doppler domain, and different complexity-performance tradeoffs have been achieved by modeling the range and range-rate of any path component as either integer or fractional multiples of the range and range-rate resolutions, respectively: [^1] For a given path component, the delay and Doppler parameters are equivalent to its range and range-rate, respectively, with a convenient scaling factor; these terms are often used interchangeably in the literature, and, for the remainder of this study, we will prefer the latter terminology.
@@
 However, this assumption may not hold true for satellite-based space sensing, as the significant relative motion between the transmitter, monitored target, and receiver can cause range migration during the OTFS frame.
-With a transmitter on a LEO satellite and a radar receiver on the same or a different satellite, or an aircraft at a lower altitude, non-cooperative space targets (e.g., space debris, asteroids, spacecraft, and unresponsive satellites) may have a range-rate of several km/s; for a system with $M=512$ subcarriers, $N=128$ symbol intervals, a subcarrier spacing of $\Delta=15$ kHz, a symbol interval of $T=1/\Delta$, and a range-rate of $\bar{v}=15$ km/s, the range variation in the OTFS frame is $\bar{v}NT=128$ m and greatly exceeds the range resolution $c/(2M\Delta)\approx 19.5$ m, where $c$ is the speed of light.
+With a transmitter on a LEO satellite and a radar receiver on an aircraft at a lower altitude, non-cooperative space targets (e.g., space debris, asteroids, spacecrafts, missiles, and unresponsive satellites) may have a bistatic range-rate of several km/s; for a system with $M=512$ subcarriers, $N=128$ symbol intervals, a subcarrier spacing of $\Delta=15$ kHz, a symbol interval of $T=1/\Delta$, and a bistatic range-rate of $\bar{v}=15$ km/s, the bistatic range variation in the OTFS frame is $\bar{v}NT=128$ m and greatly exceeds the bistatic range resolution $c/(M\Delta)\approx 39$ m, where $c$ is the speed of light.
 In this study, we investigate the use of OTFS communication signals emitted by LEO satellites for estimating the initial range, range-rate, and amplitude of space targets experiencing range migration in the OTFS frame, henceforth referred to as high-speed targets for brevity.
@@
 Additionally, we show that range migration results in a spread of this response across both dimensions.
-  • We propose an approximate implementation of the maximum likelihood (ML) estimator for target parameters, which avoids exhaustive delay-Doppler search.
+  • We propose an approximate implementation of the maximum likelihood (ML) estimator for target parameters, which avoids exhaustive two-dimensional search in the delay-Doppler plane.
 First, a block orthogonal matching pursuit (BOMP) algorithm [36] is used to obtain coarse information on the initial range and range-rate.
@@
 $\mathbbm{1}_{\mathcal{A}}$ is the indicator function of the condition $\mathcal{A}$, i.e., $\mathbbm{1}_{\mathcal{A}}=1$, if $\mathcal{A}$ holds true, and $\mathbbm{1}_{\mathcal{A}}=0$ otherwise.
-$[\,\cdot\,]_{M}$ denotes the mod $M$ operation.
+$\cdot\,]_{M}$ denotes the mod $M$ operation.
 $\delta(t)$ and $\delta[n]$ denote the Dirac delta function and the discrete unit sample function, respectively.
@@
 The radar receiver aims to estimate the initial range (delay), range-rate (Doppler shift), and amplitude of high-speed targets illuminated by the communication transmitter; it may be mounted on the same platform hosting the communication transmitter (monostatic configuration) or on a separate device (bistatic configuration), e.g., another LEO satellites, an aircraft, a high altitude platform, or a terrestrial station.
-The receiver is synchronized with the transmitter and knows the emitted symbols: for a monostatic configuration, this is readily verified [15], whereas, for a bistatic configuration, this requires the presence of some coordination [^4] For example, when the radar receiver is on a different LEO satellite, the coordination may be established via an inter-satellite link.
-or a separate reference channel able to demodulate the communication signal [40, 41].
-For a more illustration, we consider a bistatic scenario, as shown in Fig.
-1, but the following developments similarly apply to the monostatic case.
-We initially assume the presence of a single target to simplify exposition.
+The receiver is synchronized with the transmitter and knows the emitted symbols: for a monostatic configuration, this is readily verified [15, 16], whereas, for a bistatic configuration, this requires the presence of some coordination [^3] For example, when the radar receiver is on a different LEO satellite, the coordination may be established via an inter-satellite link.
+or a separate reference channel able to demodulate the communication signal [41, 42, 43].
+For illustration, we consider a bistatic radar configuration, as depicted in Fig.
+1.
+In this setup, the bistatic range of a target is defined as the total distance traveled by the probing signal (first from the communication transmitter to the target, and then from the target to the radar receiver).
+Correspondingly, the bistatic range-rate refers to the time derivative of the bistatic range, i.e., the rate at which the bistatic range changes due to motion of the transmitter, the target, or the receiver.
+It is important to note that the following analysis also applies to the monostatic case, where the transmitter and receiver are co-located; in such a configuration, the bistatic range and range-rate simplify to twice the monostatic range and range-rate, respectively.
+To streamline the exposition, we will omit the adjective "bistatic" from this point forward and use the terms range and range-rate to refer to the bistatic quantities unless otherwise specified.
+Finally, we begin by assuming the presence of a single target to simplify the initial discussion.
 Figure 1: System geometry.
 II-A Transmit signal
-[…]
+Consider an OTFS frame with bandwidth $M\Delta$ and duration $NT$, where $M$ and $N$ are the number of subcarriers and symbol intervals, respectively, $\Delta$ is the subcarrier spacing, and $T=1/\Delta$ is the symbol interval, while $1/(M\Delta)$ and $1/(NT)$ are the delay and Doppler resolutions, respectively.
+Up to a positive scaling factor accounting for the radiated energy, the baseband signal emitted by the communication transmitter is
 $x(t)=\sum_{n=0}^{N-1}\underbrace{\sum_{m=0}^{M-1}X_{\rm TF}[n,m]g_{\rm tx}\!\left(t-nT\right)\mathrm{e}^{\mathrm{i}2\pi m\Delta(t-nT)}}_{x_{n}(t)},$ (1)
@@
 By considering $g_{\rm tx}(t)$ time limited in $[0,T]$, then $x(t)$ has support $[0,NT]$.
-Also, by assuming the spectrum of $g_{\rm tx}(t)$ contained in $[-\Delta/2,\Delta/2]$, then the spectrum of $x(t)$ is contained in $[-\Delta/2,M\Delta-\Delta/2]$.
-[^6] Since a time limited signal cannot be strictly band limited, some spectral leakage outside $[-\Delta/2,M\Delta-\Delta/2]$ is present.
+Also, by assuming the spectrum of $g_{\rm tx}(t)$ contained in $\Delta/2,\Delta/2]$, then the spectrum of $x(t)$ is contained in $\Delta/2,M\Delta-\Delta/2]$.
+[^4] Since a time limited signal cannot be strictly band limited, some spectral leakage outside $\Delta/2,M\Delta-\Delta/2]$ is present.
 In practice, the transmit shaping filter may not be exactly zero outside $[0,T]$ to reduce such spectral leakage; in this latter case, $x(t)$ is only approximately contained in $[0,NT]$.
@@
 The position of the target at time $t\in[0,NT]$ is $\bm{p}(t)=\bm{p}-\bm{v}t\in\mathbb{R}^{3}$, where $\bm{p}$ and $\bm{v}$ are the initial position and the velocity, respectively; similarly, the positions of communication transmitter and the radar receiver at time $t\in[0,NT]$ are $\bm{p}_{\rm tx}(t)=\bm{p}_{\rm tx}-\bm{v}_{\rm tx}t\in\mathbb{R}^{3}$ and $\bm{p}_{\rm rx}(t)=\bm{p}_{\rm rx}-\bm{v}_{\rm rx}t\in\mathbb{R}^{3}$, respectively.
-Accordingly, for $t\in[0,NT]$, the (bistatic) target range is
+Accordingly, for $t\in[0,NT]$, the target range is
 $\bar{r}(t)=\left\|\bm{p}(t)-\bm{p}_{\rm tx}(t)\right\|+\left\|\bm{p}(t)-\bm{p}_{\rm rx}(t)\right\|.$ (3)
@@
 is the baseband representation of the target echo with respect to the reference frequency $\frac{c+\bar{v}_{\min}}{\lambda}$ and $r(t)=\bar{r}(t)-\bar{r}_{\min}$ is the excess target range (i.e., the target range in excess with respect to $\bar{r}_{\min}$).
-The support of $y(t)$ is contained in the time interval $[0,NT+r_{\max}/c]$, and its spectrum is negligible outside the frequency interval $[-\Delta/2,M\Delta+v_{\max}-\Delta/2]$.
+The support of $y(t)$ is contained in the time interval $0,NT+r_{\max}/c]$, and its spectrum is negligible outside the frequency interval $\Delta/2,M\Delta+v_{\max}/\lambda-\Delta/2]$.
 Define the excess initial range $d=\bar{d}-\bar{r}_{\min}\in[0,r_{\max}]$, the excess range-rate $v=\bar{v}-\bar{v}_{\min}\in[0,v_{\max}]$, and $\alpha=\bar{\alpha}\mathrm{e}^{-\mathrm{i}2\pi\frac{d}{\lambda}}$.
@@
 $y_{n}(t)$ $=\alpha x_{n}\left(t-\frac{r\left(nT\right)}{c}\right)\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}t}$
-[…]
+$=\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}h_{n}(\tau,\nu)x_{n}\left(t-\tau\right)\mathrm{e}^{\mathrm{i}2\pi\nu(t-\tau)}d\tau d\nu$ (12)
 and
@@
 The OTFS demodulator first transform $y(t)$ into the time-frequency domain via the Wigner Transform; in particular, the following measurements are obtained
-[…]
+$Y_{\rm TF}[n,m]=\int_{-\infty}^{\infty}y(t)g_{\rm rx}^{*}(t-nT)\mathrm{e}^{-\mathrm{i}2\pi m\Delta(t-nT)}dt,$ (14)
 for $n=0,\ldots,N-1$ and $m=0,\ldots,M-1$, where $g_{\rm rx}(t)$ is the receive shaping filter with unit energy, effective temporal duration $T$ and effective spectral width $\Delta$. Denote by
-[…]
+$\gamma(\tau,\nu)=\int_{-\infty}^{\infty}g_{\rm tx}(\beta)g^{*}_{\rm rx}(\beta-\tau)\mathrm{e}^{-\mathrm{i}2\pi\nu(\beta-\tau)}d\beta$ (15)
 the cross-ambiguity function of $g_{\rm tx}(t)$ and $g_{\rm rx}(t)$; then, the following proposition establishes a connection between $\{Y_{\rm TF}[n,m]\}$ and $\{X_{\rm TF}[n,m]\}$ that generalizes the result in [8, Theorem 1] to the case where the variation of the target range is ignored only within each symbol interval.
@@
 $Y_{\rm TF}[n,m]=\sum_{n^{\prime}=0}^{N-1}\sum_{m^{\prime}=0}^{M-1}\alpha X_{\rm TF}[n^{\prime},m^{\prime}]$
-×$\quad\times\int_{-\infty}^{\infty}g_{\rm tx}\!\left(t-n^{\prime}T-\frac{r\left(n^{\prime}T\right)}{c}\right)g^{*}_{\rm rx}\!\left(t-nT\right)$
-×$\quad\times\mathrm{e}^{\mathrm{i}2\pi m^{\prime}\Delta\left(t-n^{\prime}T-\frac{r\left(n^{\prime}T\right)}{c}\right)}\mathrm{e}^{-\mathrm{i}2\pi m\Delta\left(t-nT\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}t}dt$
+$\quad\times\int_{-\infty}^{\infty}g_{\rm tx}\!\left(t-n^{\prime}T-\frac{r\left(n^{\prime}T\right)}{c}\right)g^{*}_{\rm rx}\!\left(t-nT\right)$
+$\quad\times\mathrm{e}^{\mathrm{i}2\pi m^{\prime}\Delta\left(t-n^{\prime}T-\frac{r\left(n^{\prime}T\right)}{c}\right)}\mathrm{e}^{-\mathrm{i}2\pi m\Delta\left(t-nT\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}t}dt$
 $=\sum_{n^{\prime}=0}^{N-1}\sum_{m^{\prime}=0}^{M-1}X_{\rm TF}[n^{\prime},m^{\prime}]$
-×$\quad\times\alpha\mathrm{e}^{\mathrm{i}2\pi m\Delta\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c}\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\left(n^{\prime}T+\frac{r\left(n^{\prime}T\right)}{c}\right)}$
-×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c}\right)}$
-×$\quad\times\int_{-\infty}^{\infty}g_{\rm tx}\!\left(\beta\right)g^{*}_{\rm rx}\!\left(\beta-(n-n^{\prime})T+\frac{r\left(n^{\prime}T\right)}{c}\right)$
-×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left(\beta-(n-n^{\prime})T+\frac{r\left(n^{\prime}T\right)}{c}\right)}d\beta$
+$\quad\times\alpha\mathrm{e}^{\mathrm{i}2\pi m\Delta\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c}\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\left(n^{\prime}T+\frac{r\left(n^{\prime}T\right)}{c}\right)}$
+$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c}\right)}$
+$\quad\times\int_{-\infty}^{\infty}g_{\rm tx}\!\left(\beta\right)g^{*}_{\rm rx}\!\left(\beta-(n-n^{\prime})T+\frac{r\left(n^{\prime}T\right)}{c}\right)$
+$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left(\beta-(n-n^{\prime})T+\frac{r\left(n^{\prime}T\right)}{c}\right)}d\beta$
 $=\sum_{n^{\prime}=0}^{N-1}\sum_{m^{\prime}=0}^{M-1}X_{\rm TF}[n^{\prime},m^{\prime}]\alpha\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}nT}\mathrm{e}^{-\mathrm{i}2\pi m^{\prime}\Delta\frac{r\left(n^{\prime}T\right)}{c}}$
-×$\quad\times\gamma\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c},(m-m^{\prime})\Delta-\frac{v}{\lambda}\right),$
+$\quad\times\gamma\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c},(m-m^{\prime})\Delta-\frac{v}{\lambda}\right),$
 which completes the proof. ∎
@@
 $\Phi[k,l]$ $=\frac{1}{B}\sum_{b=0}^{B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k}{N}-\frac{vT}{\lambda}\right)$
-×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (23)
+$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (23)
 Proof.
@@
 Finally, notice that we have
-(($\star$)( ⋆ ) $=M\alpha\sum_{n=0}^{N-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}$
-×$\quad\times\left[\frac{1}{M}\sum_{m=0}^{M-1}\mathrm{e}^{\mathrm{i}2\pi\left(\frac{l-l^{\prime}}{M}-\frac{r\left(nT\right)\Delta}{c}\right)m}\right]$
+$(\star)$ $=M\alpha\sum_{n=0}^{N-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}$
+$\quad\times\left[\frac{1}{M}\sum_{m=0}^{M-1}\mathrm{e}^{\mathrm{i}2\pi\left(\frac{l-l^{\prime}}{M}-\frac{r\left(nT\right)\Delta}{c}\right)m}\right]$
 $=M\alpha\sum_{n=0}^{N-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}$
-×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(nT\right)\Delta}{c}\right)\right)$
+$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(nT\right)\Delta}{c}\right)\right)$
 $=NM\alpha\sum_{b=0}^{B-1}\frac{1}{B}\left[\frac{B}{N}\sum_{n=bN/B}^{(b+1)N/B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}\right]$
-×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right)$ (25)
+$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right)$ (25)
 $=NM\alpha\Phi[k-k^{\prime},l-l^{\prime}],$
@@
 III-A Some insights on Proposition 2
-Recall that $\mathcal{D}_{Q}(\nu)$ is a periodic function of $\nu$ with period one and is zero at multiple of $1/Q$, except for $\nu=0,\pm 1,\pm 2,\ldots$ where it is one; also, $|\mathcal{D}_{Q}(\nu)|$ has a main-lobe centered at 00 of width $2/Q$ (between zero crossings) and $Q-2$ sidelobes of width $1/Q$ (between zero crossings) in $[-0.5,0.5]$.
+Recall that $\mathcal{D}_{Q}(\nu)$ is a periodic function of $\nu$ with period one and is zero at multiple of $1/Q$, except for $\nu=0,\pm 1,\pm 2,\ldots$ where it is one; also, $|\mathcal{D}_{Q}(\nu)|$ has a main-lobe centered at 0 of width $2/Q$ (between zero crossings) and Q2Q-2 sidelobes of width $1/Q$ (between zero crossings) in [0.5,0.5][-0.5,0.5].
 The first sidelobe is approximately $13$ dB down from the main-lobe, and the subsequent sidelobes fall off at about $6$ dB per octave [44]; following [8], for large $Q$, we can assume $|\mathcal{D}_{Q}(\nu)|\approx 0$ for $\nu\notin[-D/Q,D/Q]$, with $0<2D+1\ll Q$, where the non-negative integer $D$ is a design parameter that depends upon the desired approximation error.
@@
 Remark 2.
-For large $N/B$ and $M$, the sums in (26) only contain few dominant terms. To illustrate this point, write first the range at time $bT_{B}$ and the range-rate as
+For large $N/B$ and $M$, the sums in (26) only contain few dominant terms.
+To illustrate this point, denote by $R_{\rm r}=c/(M\Delta)$ and $R_{\rm rr}=\lambda/(NT)$ the range and range-rate resolutions [41], respectively, and write the range at time $bT_{B}$ and the range-rate as
 $r\left(bT_{B}\right)$ $=(l_{b,{\rm int}}+l_{b,{\rm fra}})R_{\rm r},$ (27a)
@@
 $\Phi[k,l]$ $=\frac{1}{B}\sum_{b=0}^{B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k_{\rm int}-k_{\rm fra}}{B}\right)b}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k_{\rm int}-k_{\rm fra}}{N}\right)$
-×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l_{b,{\rm int}}-l_{b,{\rm fra}}}{M}\right)\right).$ (28)
+$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l_{b,{\rm int}}-l_{b,{\rm fra}}}{M}\right)\right).$ (28)
 Finally, from the properties of the Dirichlet function discussed in Sec.
-III-A, we have $|\Phi[k,l]|\approx 0$ if $[k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $[l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$, for $b=0,\ldots,B-1$: more on this in Remark 4.
+III-A, we have $|\Phi[k,l]|\approx 0$ if $k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$, for $b=0,\ldots,B-1$: more on this in Remark 4.
 ∎
@@
 $\Psi_{b,k^{\prime},l^{\prime}}[k,l]=\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}}{M}}$
-×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k}{N}-\frac{vT}{\lambda}\right)$
-×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right),$ (39)
-if $10\leq l^{\prime}\leq\big{\lceil}M-\frac{r\left(bT_{B}\right)M}{cT}\big{\rceil}-$ 1, and
+$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k}{N}-\frac{vT}{\lambda}\right)$
+$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right),$ (39)
+if $0\leq l^{\prime}\leq\big\lceil M-\frac{r\left(bT_{B}\right)M}{cT}\big\rceil-1$, and
 $\Psi_{b,k^{\prime},l^{\prime}}[k,l]=\mathrm{e}^{-\mathrm{i}2\pi\frac{k^{\prime}}{N}}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}-M}{M}}$
-×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\left[\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k}{N}-\frac{vT}{\lambda}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]$
-×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l+M}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right),$ (40)
+$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\left[\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k}{N}-\frac{vT}{\lambda}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]$
+$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l+M}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right),$ (40)
 if $\big\lceil M-\frac{r\left(bT_{B}\right)M}{cT}\big\rceil\leq l^{\prime}\leq M-1$.
@@
 For large $N/B$ and $M$, the summations in (37) only contain few dominant dominant terms; indeed, from the properties of the Dirichlet function discussed in Sec.
-III-A, we have $|\Psi_{k^{\prime},l^{\prime}}[k,l]|\approx 0$ if $[k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $[l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$, for $b=0,\ldots,B-1$.
+III-A, we have $|\Psi_{k^{\prime},l^{\prime}}[k,l]|\approx 0$ if $k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$, for $b=0,\ldots,B-1$.
 This result is similar to that obtained for ideal shaping filters in Remark 2.
@@
 $Y_{{\rm DD}}[k,l]=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{k^{\prime}=0}^{N-1}$
-×$\quad\times\Bigg{[}\sum_{l^{\prime}=0}^{M-1-l_{b,{\rm int}}-\lfloor l_{b,{\rm fra}}\rfloor}X_{\rm DD}[k^{\prime},l^{\prime}]\Psi_{b,k^{\prime},l^{\prime}}[k-k^{\prime},l-l^{\prime}]$
+$\quad\times\Bigg[\sum_{l^{\prime}=0}^{M-1-l_{b,{\rm int}}-\lfloor l_{b,{\rm fra}}\rfloor}X_{\rm DD}[k^{\prime},l^{\prime}]\Psi_{b,k^{\prime},l^{\prime}}[k-k^{\prime},l-l^{\prime}]$
 $\quad+\sum_{l^{\prime}=M-l_{b,{\rm int}}-\lfloor l_{b,{\rm fra}}\rfloor}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\Psi_{b,k^{\prime},l^{\prime}}[k-k^{\prime},l-l^{\prime}]\Bigg]$
 $=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{\bar{k}=k_{\rm int}-k}^{N+k_{\rm int}-k-1}$
-×$\quad\times\Bigg{[}\sum_{\bar{l}_{b}=l_{b,{\rm int}}-l}^{M-l-1-\lfloor l_{b,{\rm fra}}\rfloor}\Psi_{b,k^{\prime},l^{\prime}}[k_{\rm int}-\bar{k},l_{b,{\rm int}}-\bar{l}_{b}]$
-×$\quad\times X_{\rm DD}\Big{[}[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big{]}$
+$\quad\times\Bigg[\sum_{\bar{l}_{b}=l_{b,{\rm int}}-l}^{M-l-1-\lfloor l_{b,{\rm fra}}\rfloor}\Psi_{b,k^{\prime},l^{\prime}}[k_{\rm int}-\bar{k},l_{b,{\rm int}}-\bar{l}_{b}]$
+$\quad\times X_{\rm DD}\Big[[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big]$
 $\quad+\sum_{\bar{l}_{b}=-l-\lfloor l_{b,{\rm fra}}\rfloor}^{l_{b,{\rm int}}-l-1}\Psi_{b,k^{\prime},l^{\prime}}[k_{\rm int}-\bar{k},l_{b,{\rm int}}-\bar{l}_{b}-M]$
-×$\quad\times X_{\rm DD}\Big{[}[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big{]}\Bigg{]}$
+$\quad\times X_{\rm DD}\Big[[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big]\Bigg]$
 $=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{\bar{k}=-\lfloor N/2\rfloor}^{\lceil N/2\rceil-1}\sum_{\bar{l}_{b}=-l-\lfloor l_{b,{\rm fra}}\rfloor}^{M-l-1-\lfloor l_{b,{\rm fra}}\rfloor}$
-×$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l_{b,{\rm int}}+l_{b,{\rm fra}}}{M}}\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l-l_{b,{\rm int}}+\bar{l}_{b}}{M}}$
-×$\quad\times\mathrm{e}^{\mathrm{i}2\pi\left(\frac{\bar{k}+k_{\rm fra}}{B}\right)b}A_{b,\bar{k},\bar{l}_{b}}[k,l]\mathcal{D}_{M}\!\left(\frac{\bar{l}_{b}+l_{b,{\rm fra}}}{M}\right)$
-×$\quad\times X_{\rm DD}\Big{[}[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big{]},$ (41)
+$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l_{b,{\rm int}}+l_{b,{\rm fra}}}{M}}\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l-l_{b,{\rm int}}+\bar{l}_{b}}{M}}$
+$\quad\times\mathrm{e}^{\mathrm{i}2\pi\left(\frac{\bar{k}+k_{\rm fra}}{B}\right)b}A_{b,\bar{k},\bar{l}_{b}}[k,l]\mathcal{D}_{M}\!\left(\frac{\bar{l}_{b}+l_{b,{\rm fra}}}{M}\right)$
+$\quad\times X_{\rm DD}\Big[[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big],$ (41)
 where $A_{b,\bar{k},\bar{l}_{b}}[k,l]$ is equal to
@@
 $\left[\mathcal{D}_{\frac{N}{B}}\!\left(-\frac{\bar{k}+k_{\rm fra}}{N}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]\mathrm{e}^{-\mathrm{i}2\pi\frac{[k-k_{\rm int}+\bar{k}]_{N}}{N}},$ (43)
-if $-l-\lfloor l_{b,{\rm fra}}\rfloor\leq\bar{l}_{b}\leq l_{b,{\rm int}}-l-1$. If $l_{b,{\rm int}}=l_{{\rm int}}$ and $l_{b,{\rm fra}}=0$ for $b=0,\ldots,B-1$, then (41) becomes
+if $l-\lfloor l_{b,{\rm fra}}\rfloor\leq\bar{l}_{b}\leq l_{b,{\rm int}}-l-1$. If $l_{b,{\rm int}}=l_{{\rm int}}$ and $l_{b,{\rm fra}}=0$ for $b=0,\ldots,B-1$, then (41) becomes
 $Y_{{\rm DD}}[k,l]=\alpha\sum_{\bar{k}=-\lfloor N/2\rfloor}^{\lceil N/2\rceil-1}\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l_{{\rm int}}}{M}}\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l-l_{{\rm int}}}{M}}\\ \times A_{\bar{k}}[k,l]X_{\rm DD}\Big[[k-k_{\rm int}+\bar{k}]_{N},[l-l_{{\rm int}}]_{M}\Big],$ (44)
@@
 $\left[\mathcal{D}_{N}\!\left(-\frac{\bar{k}+k_{\rm fra}}{N}\right)-\frac{1}{N}\right]\mathrm{e}^{-\mathrm{i}2\pi\frac{[k-k_{\rm int}+\bar{k}]_{N}}{N}},$ (46)
-if $10\leq l\leq l_{{\rm int}}-$ 1, that is consistent with [8, Theorem 2]. ∎
+if $0\leq l\leq l_{{\rm int}}-1$, that is consistent with [8, Theorem 2]. ∎
 Remark 8.
@@
 $Y_{{\rm DD}}[k,l]\approx\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{NM}l}\Theta_{k^{\prime},l^{\prime}}[k,l]$
-×$\quad\times\frac{1}{B}\sum_{b=0}^{B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k^{\prime}-k_{\rm int}-k_{\rm fra}}{B}\right)b}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k^{\prime}-k_{\rm int}-k_{\rm fra}}{N}\right)$
-×$\quad\times\mathcal{D}_{M}\!\left(-\frac{l^{\prime}-l_{b,{\rm int}}}{M}\right)X_{\rm DD}\Big{[}[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\Big{]},$ (50)
+$\quad\times\frac{1}{B}\sum_{b=0}^{B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k^{\prime}-k_{\rm int}-k_{\rm fra}}{B}\right)b}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k^{\prime}-k_{\rm int}-k_{\rm fra}}{N}\right)$
+$\quad\times\mathcal{D}_{M}\!\left(-\frac{l^{\prime}-l_{b,{\rm int}}}{M}\right)X_{\rm DD}\Big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\Big],$ (50)
 where
-$\Theta_{k^{\prime},l^{\prime}}[k,l]=\left\{\begin{matrix}1&l^{\prime}\leq l\leq M-1,\\ \mathrm{e}^{-\mathrm{i}2\pi\frac{[k-k^{\prime}]_{N}}{N}}&0\leq l\leq l^{\prime}-1.\\ \end{matrix}\right.$ (51)
+[…]
 Based on (50), we can now write
@@
 $M$ Number of subcarriers 512
-$N$ Number of symbol intervals 128
-$R_{\rm r}$ Range resolution 19.5 m
-$R_{\rm rr}$ Range-rate resolution 4.4 m/s
-$[\bar{r}_{\rm min},\,\bar{r}_{\rm max}]$ Inspected range interval [20, 39] km
-$[\bar{v}_{\rm min},\,\bar{v}_{\rm max}]$ Inspected range-rate interval [15, 16] km/s
+NN Number of symbol intervals 128
+$R_{\rm r}$ Range resolution $\approx$ 39 m
+$R_{\rm rr}$ Range-rate resolution $\approx$ 8.8 m/s
+$\bar{r}_{\rm min},\,\bar{r}_{\rm max}]$ Inspected range interval [20, 39] km
+$\bar{v}_{\rm min},\,\bar{v}_{\rm max}]$ Inspected range-rate interval [15, 16] km/s
 VI Numerical analysis
 To assess the system performance, we consider the parameters in Table I.
-The maximum variation of the target range during the OTFS frame is $\bar{v}_{\rm max}NT=136$ m, corresponding to about $7$ times the range resolution $R_{\rm r}$; the initial range $d$ and the range-rate $v$ are randomly generated, while a Swerling I fluctuation model is considered for the amplitude $\alpha$.
+The maximum variation of the target range during the OTFS frame is $\bar{v}_{\rm max}NT\approx 136$ m, corresponding to about $3.5$ times the range resolution $R_{\rm r}$; the initial range $d$ and the range-rate $v$ are randomly generated, while a Swerling I fluctuation model is considered for the amplitude $\alpha$.
 The information symbols $\{X_{\rm DD}[k,l]\}$ are independently drawn from a unit-energy $4$-QAM constellation.
@@
 $Y_{\rm DD}[k,l]$ $=\frac{1}{\sqrt{NM}}\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)}$
-×$\quad\times\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n,m^{\prime}]X_{\rm TF}[n,m^{\prime}]$
+$\quad\times\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n,m^{\prime}]X_{\rm TF}[n,m^{\prime}]$
 $\quad+\frac{1}{\sqrt{NM}}\sum_{n=1}^{N-1}\sum_{m=0}^{M-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)}$
-×$\quad\times\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n-1,m^{\prime}]X_{\rm TF}[n-1,m^{\prime}]$
+$\quad\times\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n-1,m^{\prime}]X_{\rm TF}[n-1,m^{\prime}]$
 $=Y_{\rm DD}^{\rm ici}[k,l]+Y_{\rm DD}^{\rm isi}[k,l].$ (59)
@@
 $Y_{\rm DD}^{\rm ici}[k,l]$ $=\frac{1}{NM}\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)}\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n,m^{\prime}]$
-×$\quad\times\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{\mathrm{i}2\pi\left(\frac{nk^{\prime}}{N}-\frac{m^{\prime}l^{\prime}}{M}\right)}$
+$\quad\times\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{\mathrm{i}2\pi\left(\frac{nk^{\prime}}{N}-\frac{m^{\prime}l^{\prime}}{M}\right)}$
 $=\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\Pi_{k,l}^{\rm ici}[k^{\prime},l^{\prime}],$ (60)
@@
 $\Pi_{k,l}^{\rm ici}[k^{\prime},l^{\prime}]$ $=\frac{1}{NM}\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n,m^{\prime}]$
-×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}\right)}e^{j2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}.$ (61)
+$\quad\times\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}\right)}e^{j2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}.$ (61)
 $\Pi_{k,l}^{\rm ici}[k^{\prime},l^{\prime}]=\frac{\alpha}{NM^{2}}\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}\sum_{m^{\prime}=0}^{M-1}\left[\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}nT}\mathrm{e}^{-\mathrm{i}2\pi m^{\prime}\Delta\frac{r\left(nT\right)}{c}}\sum_{p=0}^{\big\lceil M-\frac{r\left(nT\right)M}{cT}\big\rceil-1}\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left(\frac{pT}{M}+\frac{r\left(nT\right)}{c}\right)}\right]\mathrm{e}^{\mathrm{i}2\pi n\left(\frac{k^{\prime}-k}{N}\right)}\mathrm{e}^{\mathrm{i}2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}$
@@
 $Y_{{\rm DD}}^{\rm ici}[k,l]$ $=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{\big\lceil M-\frac{r\left(bT_{B}\right)M}{cT}\big\rceil-1}X_{\rm DD}[k^{\prime},l^{\prime}]$
-×$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}}{M}}$
-×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)$
-×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (63)
+$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}}{M}}$
+$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)$
+$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (63)
 As to the term $Y_{\rm DD}^{\rm isi}[k,l]$ in (59), we have:
 $Y_{\rm DD}^{\rm isi}[k,l]$ $=\frac{1}{NM}\sum_{n=1}^{N-1}\sum_{m=0}^{M-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)}\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n-1,m^{\prime}]$
-×$\quad\times\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{\mathrm{i}2\pi\left(\frac{(n-1)k^{\prime}}{N}-\frac{m^{\prime}l^{\prime}}{M}\right)}$
+$\quad\times\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{\mathrm{i}2\pi\left(\frac{(n-1)k^{\prime}}{N}-\frac{m^{\prime}l^{\prime}}{M}\right)}$
 $=\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\Pi_{k,l}^{\rm isi}[k^{\prime},l^{\prime}]\mathrm{e}^{-\mathrm{i}2\pi\frac{k^{\prime}}{N}},$ (64)
@@
 $\Pi_{k,l}^{\rm isi}[k^{\prime},l^{\prime}]$ $=\frac{1}{NM}\sum_{n=1}^{N-1}\sum_{m=0}^{M-1}\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n-1,m^{\prime}]$
-×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}\right)}e^{j2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}.$ (65)
+$\quad\times\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}\right)}e^{j2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}.$ (65)
 $\Pi_{k,l}^{\rm isi}[k^{\prime},l^{\prime}]=\frac{\alpha}{NM^{2}}\!\sum_{n=1}^{N-1}\sum_{m=0}^{M-1}\sum_{m^{\prime}=0}^{M-1}\!\left[\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}nT}\mathrm{e}^{-\mathrm{i}2\pi m^{\prime}\Delta\frac{r\left((n-1)T\right)}{c}}\!\!\!\!\sum_{p=\left\lceil M-\frac{r\left((n-1)T\right)M}{cT}\right\rceil}^{M-1}\!\!\!\!\!\!\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left(\frac{pT}{M}-T+\frac{r\left((n-1)T\right)}{c}\right)}\!\right]\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}\right)}\!\mathrm{e}^{\mathrm{i}2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}$
@@
 $=\frac{\alpha}{B}\sum_{b=0}^{B-1}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\left[\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]$
-×$\quad\times\sum_{p=\left\lceil M-\frac{r\left(bT_{B}\right)}{c}\frac{M}{T}\right\rceil}^{M-1}\mathcal{D}_{M}\!\left(\frac{p-M+\frac{r\left(bT_{B}\right)}{c}\frac{M}{T}-l}{M}\right)\delta\big{[}[p-M-l^{\prime}]_{M}\big{]}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{p-M}{M}}.$ (66)
+$\qquad\times\sum_{p=\left\lceil M-\frac{r\left(bT_{B}\right)}{c}\frac{M}{T}\right\rceil}^{M-1}\mathcal{D}_{M}\!\left(\frac{p-M+\frac{r\left(bT_{B}\right)}{c}\frac{M}{T}-l}{M}\right)\delta\big[[p-M-l^{\prime}]_{M}\big]\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{p-M}{M}}.$ (66)
 Upon expanding $\Pi_{k,l}^{\rm isi}[k^{\prime},l^{\prime}]$ as shown in (66) at the top of the next page, we obtain
 $Y_{{\rm DD}}^{\rm ici}[k,l]$ $=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=\big\lceil M-\frac{r\left(bT_{B}\right)M}{cT}\big\rceil}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{-\mathrm{i}2\pi\frac{k^{\prime}}{N}}$
-×$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}-M}{M}}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}$
-×$\quad\times\left[\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]$
-×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}+M}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (67)
+$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}-M}{M}}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}$
+$\quad\times\left[\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]$
+$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}+M}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (67)
 Finally, (37) follows from (63) and (67).

### v3

@@
 Abstract
-This study investigates a communication-centric integrated sensing and communication (ISAC) system that utilizes orthogonal time frequency space (OTFS) modulated signals emitted by low Earth orbit (LEO) satellites to estimate the parameters of space targets experiencing range migration, henceforth referred to as high-speed targets.
-Leveraging the specific signal processing performed by OTFS transceivers, we derive a novel input-output model for the echo generated by a high-speed target in scenarios where ideal and rectangular shaping filters are employed.
-Our findings reveal that the target response exhibits a sparse structure in the delay-Doppler domain, dependent solely upon the initial range and range-rate; notably, range migration causes a spread in the target response, marking a significant departure from previous studies.
-Utilizing this signal structure, we propose an approximate implementation of the maximum likelihood estimator for the target’s initial range, range-rate, and amplitude.
-The estimation process involves obtaining coarse information on the target response using a block orthogonal matching pursuit algorithm, followed by a refinement step using a bank of matched filters focused on a smaller range and range-rate region.
+This study investigates a communication-centric integrated sensing and communication system that utilizes orthogonal time-frequency space (OTFS) modulated signals emitted by low Earth orbit satellites to estimate the parameters of space targets experiencing range migration, hereinafter referred to as high-speed targets.
+Leveraging the signal samples produced by off-the-shelf OTFS demodulators, we derive a novel input-output model for the echo generated by a high-speed target when ideal and rectangular shaping filters are employed.
+Our findings reveal that the target response exhibits a sparse structure in the delay-Doppler domain, whose support is determined by the target initial-range and range-rate.
+Range migration induces a structured spread of this response, which is explicitly characterized in the paper and differs from that in previous models.
+We propose an approximate implementation of the maximum likelihood estimator for the target initial-range, range-rate, and amplitude.
+The estimation process first obtains coarse information on the target response using a block orthogonal matching pursuit algorithm, followed by a refinement step based on a bank of matched filters focused on a smaller initial-range/range-rate region.
+The proposed single-target procedure is extended to multiple targets via iterative estimation, reconstruction, and cancellation of dominant echoes.
 Finally, numerical examples are provided to evaluate the estimation performance.
 I Introduction
-Future wireless networks and Internet-of-Things (IoT) will utilize terrestrial and non-terrestrial stations to ensure seamless and ubiquitous connectivity [1].
-Among spaceborne transceivers, low Earth orbit (LEO) satellites are attracting much attention for their lower latency and higher data rates [2]; also, they may efficiently provide integrated sensing and communications (ISAC) functionalities over large regions with the same signal waveform and hardware architecture [3, 4, 5].
-Orthogonal time frequency space (OTFS) modulation is a promising transmission scheme for future LEO constellations [6, 7].
-OTFS-based transceivers exploit channel diversity across both time and frequency, thus providing better communication performance in doubly-dispersive fading channels as compared to schemes using orthogonal frequency division multiplexing [8, 9, 10].
-They are also suitable for radar sensing, as the underlying signal processing resembles that of pulsed-Doppler radars [11]; in particular, the range and range-rate of prospective scattering objects can be inferred from the estimated delay-Doppler channel response [12].
-These unique features enable OTFS-based LEO satellites to inherently support ISAC in various forms [13, 14].
-Unlike terrestrial stations, LEO-based stations have the capability to also cover aerial and space regions, and this expanded coverage opens up new application scenarios, such as real-time environmental monitoring, enhanced global navigation systems, and space exploration.
-Communication-centric ISAC solutions appear particularly promising in this context, as existing data signals can be opportunistically reused to add the sensing functions at a low cost, without altering the underlying communication protocol or affecting its performance [12, 15, 16, 17, 18, 19]; this approach is indeed considered in this study.
-Several solutions for signal estimation and detection have been deployed for OTFS-based applications, which exploit the channel sparsity in the delay-Doppler domain, and different complexity-performance tradeoffs have been achieved by modeling the range and range-rate of any path component as either integer or fractional multiples of the range and range-rate resolutions, respectively: [^1] For a given path component, the delay and Doppler parameters are equivalent to its range and range-rate, respectively, with a convenient scaling factor; these terms are often used interchangeably in the literature, and, for the remainder of this study, we will prefer the latter terminology.
-see [20] and references therein.
-For example, the work in [8] has derived a widely-used input-output signal model under the assumption of integer range and fractional range-rate for both ideal and rectangular shaping filters and has presented a message passing algorithm for data symbol detection.
-[^2] Hereafter, integer (fractional) range or range-rate means that the range or range-rate is an integer (fractional) multiple of the respective resolution.
-Leveraging the model in [8], a pilot-aided channel estimation has been developed in [21]; here, pilot, guard, and data symbols are arranged so as to avoid interference between pilot and data symbols at the receiver, thus facilitating channel estimation and data detection.
-Another channel estimation method based on a sparse Bayesian learning algorithm has been provided in [22] and, then, extended in [23] to handle both fractional range and range-rate.
-More recently, a channel estimation method relying on the insertion of random pilots has been presented in [24].
-On a different side, leveraging the model in [8], the work in [12] has proposed to exploit existing OTFS communication signals for radar sensing to implement a simplified form of communication-centric ISAC.
-This idea has been further elaborated in [25], where the Cramér-Rao lower bounds in the estimation of the range and range-rate are derived, in [26], where different pulse shapes are investigated, in [27], where a modified orthogonal matching pursuit algorithm with fractional refinement is proposed, and in [28], where the estimation of the target parameter is formulated as an atomic norm minimization.
-The 3D-structured sparsity of MIMO-OTFS channel is further explored in [29], where the downlink channel estimation problem is formulated as a sparse signal recovery problem.
-In addition, OTFS signals have been used for joint localization and communications in [30] and for target tracking in [31].
-Previous studies assume that the length (range) of a channel path between the transmitter and receiver remains constant within the OTFS frame, which is typically valid for terrestrial applications.
-However, this assumption may not hold true for satellite-based space sensing, as the significant relative motion between the transmitter, monitored target, and receiver can cause range migration during the OTFS frame.
-With a transmitter on a LEO satellite and a radar receiver on an aircraft at a lower altitude, non-cooperative space targets (e.g., space debris, asteroids, spacecrafts, missiles, and unresponsive satellites) may have a bistatic range-rate of several km/s; for a system with $M=512$ subcarriers, $N=128$ symbol intervals, a subcarrier spacing of $\Delta=15$ kHz, a symbol interval of $T=1/\Delta$, and a bistatic range-rate of $\bar{v}=15$ km/s, the bistatic range variation in the OTFS frame is $\bar{v}NT=128$ m and greatly exceeds the bistatic range resolution $c/(M\Delta)\approx 39$ m, where $c$ is the speed of light.
-In this study, we investigate the use of OTFS communication signals emitted by LEO satellites for estimating the initial range, range-rate, and amplitude of space targets experiencing range migration in the OTFS frame, henceforth referred to as high-speed targets for brevity.
-We follow a communication-centric approach where the existing waveform and protocol are not modified to maintain the communication performance unaltered; instead, the focus is on the design of the opportunistic radar receiver and on the analysis of its achievable performance.
-Sensing high-speed targets is a challenging problem, and previous studies in the radar literature have primarily utilized unmodulated or modulated coherent pulse trains (typically, a train of linear frequency modulated pulses) and developed motion compensation strategies operating in the time or frequency domain to integrate consecutive pulses [32, 33, 34, 35].
-However, the impact of range migration on OTFS signals is unexplored, and existing motion compensation methodologies cannot directly be applied in the delay-Doppler domain when off-the-shelf OTFS modulator and demodulator are employed.
-This study aims to provide initial insights on parameter estimation of high-speed targets using OTFS communication signals, without changing the structure of existing transceivers.
+Future wireless networks and Internet-of-Things (IoT) systems will utilize non-terrestrial stations to ensure seamless and ubiquitous connectivity [1].
+Among space transceivers, low Earth orbit (LEO) satellites are attracting significant attention due to their low latency, high data rates, and capability to support integrated sensing and communication (ISAC) [2, 3, 4], enabling applications such as environmental monitoring, global navigation, and space exploration.
+Orthogonal time frequency space (OTFS) modulation is a promising transmission scheme for future LEO constellations [5].
+OTFS-based transceivers exploit channel diversity across both time and frequency, thus providing improved robustness in doubly-dispersive fading channels compared to orthogonal frequency division multiplexing [6, 7].
+They are also suitable for radar sensing, as the underlying signal processing resembles that of pulsed-Doppler radars [8]; in particular, the range and range-rate parameters of the detected objects are inferred from the estimated delay-Doppler channel response.
+[^1] In this work, the range is defined as the propagation distance of the signal path, while the range-rate is defined as its time derivative.
+Hence, the range and range-rate provide a convenient physical parameterization of the delay and Doppler shifts along the propagation path, respectively.
+These features enable OTFS-based LEO satellites to support ISAC in various forms [9, 10].
+Communication-centric ISAC appears particularly promising in this context, as existing signals can be reused for sensing at low cost without altering the underlying communication protocol [11, 12, 13, 14, 15]; this approach is considered in this study.
+A large body of work has addressed signal detection and channel estimation for OTFS systems by exploiting the inherent sparsity of the delay-Doppler representation when range migration over the processed frame can be ignored, an assumption that is usually valid for terrestrial scenarios; see [16] and references therein.
+For example, [6] derived a widely-used input-output signal model under the assumption of integer range and fractional range-rate for both ideal and rectangular shaping filters, along with a message passing algorithm for data detection.
+[^2] Integer (fractional) range or range-rate means that the range or range-rate is an integer (fractional) multiple of the respective resolution.
+Based on this model, pilot-aided channel estimation schemes have been proposed in [17], where guard symbols are introduced to avoid interference between pilot and data symbols at the receiver.
+Sparse Bayesian learning approaches have been developed in [18] and extended in [19] to handle both fractional range and range-rate.
+On the sensing side, [11] proposed to exploit OTFS communication signals for radar sensing within a communication-centric ISAC framework.
+This idea has been further developed in several directions, including the derivation of Cramér–Rao lower bounds for range and range-rate estimation [20], the analysis of pulse shaping effects [21], the design of sparse recovery algorithms with fractional refinement [22], and atomic norm-based formulations for target parameter estimation [23].
+The 3D structured sparsity of MIMO-OTFS channels has been investigated in [24], where the downlink channel estimation problem is formulated as a sparse signal recovery task.
+In addition, OTFS signals have been employed for joint localization and communication in [25] and for target tracking in [26].
+Assuming the range of each propagation path to be constant may no longer be valid in satellite-based sensing.
+For instance, in a scenario with the communication transmitter on a LEO satellite and the radar receiver on an airborne platform, non-cooperative space targets (e.g., debris, asteroids, spacecraft, or missiles) may exhibit range rates of several km/s.
+When considering a system with $M=512$ subcarriers, $N=128$ symbol intervals, a subcarrier spacing of $\Delta=15$ kHz, a symbol interval of $T=1/\Delta$, and a range rate of $\bar{v}=15$ km/s, the range variation over the OTFS frame is $\bar{v}NT=128$ m, which greatly exceeds the range resolution $c/(M\Delta)\approx 39$ m, where $c$ is the speed of light.
+This work investigates the use of OTFS communication signals for estimating the parameters of high-speed targets undergoing range migration within the OTFS frame.
+We adopt a communication-centric approach in which the waveform is not modified to preserve communication performance, and we focus on characterizing the target echo model, designing the radar receiver, and analyzing estimation performance.
+Sensing high-speed targets is challenging, and prior radar studies have primarily relied on unmodulated or modulated coherent pulse trains (typically, a train of linear frequency-modulated pulses) and developed motion-compensation strategies in the time or frequency domain to coherently integrate consecutive pulses [27, 28, 29, 30].
+However, the impact of range migration on OTFS signals remains unexplored, and existing motion-compensation techniques cannot be directly applied in the delay-Doppler domain under standard modulation and demodulation processing.
+This study aims to provide initial insights into parameter estimation for high-speed targets using OTFS communication signals, without altering the transceiver structure.
 The major contributions are summarized below.
-  • We derive a novel model for the echo produced by a high-speed target considering both ideal and rectangular shaping filters; in particular, the target range is assumed constant over a block of a few symbols and changes from block-to-block consistently with the range-rate.
-The proposed model subsumes the one in [8] when the target range migration is negligible and the target range is an integer multiple of the range resolution.
-  • We demonstrate that the response of high-speed targets maintains a sparse structure in the delay-Doppler domain, determined entirely by the initial range and range-rate.
-Additionally, we show that range migration results in a spread of this response across both dimensions.
-  • We propose an approximate implementation of the maximum likelihood (ML) estimator for target parameters, which avoids exhaustive two-dimensional search in the delay-Doppler plane.
-First, a block orthogonal matching pursuit (BOMP) algorithm [36] is used to obtain coarse information on the initial range and range-rate.
-Next, a refined estimate of the target parameters is derived using a bank of matched filters focused on a narrower delay-Doppler region.
-The proposed procedure can be combined with the CLEAN algorithm in [37, 38, 39, 40, 16] to handle multiple targets.
-  • Numerical simulations are conducted to validate the proposed echo model and assess the performance of the target parameter estimator under both ideal and rectangular shaping filters.
-Results are also compared against those obtained using the methods in [21] and [29] that assume no range migration occurs within the OTFS frame.
+  • We derive a model for the echo produced by a high-speed target in the delay-Doppler domain, considering both ideal and rectangular shaping filters; the target range is assumed constant over a block of a few symbols, but changes from block to block consistently with the range-rate.
+Ideal filters yield a circular convolution model, whereas rectangular filters break this convolution structure because of inter-carrier interference (ICI) and inter-symbol interference (ISI).
+This model subsumes the one in [6] when the target-range migration is negligible.
+  • We demonstrate that the response of high-speed targets maintains a sparse structure in the delay-Doppler domain, determined by its initial-range (i.e., the target range at the beginning of the OTFS frame) and range-rate.
+Specifically, range migration results in an expanded support of this response not captured by existing models.
+  • We propose an approximate implementation of the maximum likelihood (ML) estimator for target parameters, which avoids exhaustive search over the entire parameter space.
+First, a block orthogonal matching pursuit (BOMP) algorithm [31] is used to obtain coarse information on the initial-range and range-rate.
+Next, a refined estimate of the target parameters is derived using a bank of matched filters focused on a narrower region.
+The proposed two-step estimator can be combined with the CLEAN algorithm in [32, 33, 34, 35, 12] to handle multiple targets.
+  • Numerical simulations are conducted to validate the proposed echo model and assess the performance of the proposed two-step estimator under different system configurations and numbers of targets, also in comparison with the estimators in [17] and [24].
 Organization
@@
 Sec.
-II presents the system description and the design assumptions.
-Secs.
-III and IV provide the proposed signal model for ideal and rectangular shaping filters, respectively.
+II presents the system description.
 Sec.
-V illustrates the proposed procedure for target parameter estimation.
+III presents the delay–Doppler input–output model under ideal and rectangular shaping filters.
 Sec.
-VI discusses some numerical examples.
+IV contains the proposed procedure for target parameter estimation.
 Sec.
-VII contains the conclusions.
-Finally, some mathematical derivations are deferred to the Appendix.
+V discusses some numerical examples.
+Sec.
+VI contains the conclusions.
+Finally, the appendices contain some mathematical derivations.
 Notation
@@
 II System description
-Consider a sensing system consisting of a communication transmitter and a radar receiver.
-The communication transmitter is mounted on a LEO satellite, uses an OTFS modulation, and operates with a carrier frequency $c/\lambda$, where $\lambda$ is the carrier wavelength; the emitted waveform is specified by the underlying communication protocol and cannot be modified to adjust the radar needs.
-The radar receiver aims to estimate the initial range (delay), range-rate (Doppler shift), and amplitude of high-speed targets illuminated by the communication transmitter; it may be mounted on the same platform hosting the communication transmitter (monostatic configuration) or on a separate device (bistatic configuration), e.g., another LEO satellites, an aircraft, a high altitude platform, or a terrestrial station.
-The receiver is synchronized with the transmitter and knows the emitted symbols: for a monostatic configuration, this is readily verified [15, 16], whereas, for a bistatic configuration, this requires the presence of some coordination [^3] For example, when the radar receiver is on a different LEO satellite, the coordination may be established via an inter-satellite link.
-or a separate reference channel able to demodulate the communication signal [41, 42, 43].
-For illustration, we consider a bistatic radar configuration, as depicted in Fig.
+Consider a system consisting of a communication transmitter and a radar receiver.
+The communication transmitter is mounted on a LEO satellite, uses an OTFS modulation, and operates at a carrier frequency $c/\lambda$, where $\lambda$ is the carrier wavelength; the emitted waveform is specified by the underlying communication protocol and is not modified here.
+The radar receiver aims to estimate the initial-range, range-rate, and amplitude of high-speed targets illuminated by the communication transmitter; it may be mounted on the same platform hosting the communication transmitter (monostatic configuration) or on a separate device (bistatic configuration), e.g., another LEO satellite, an aircraft, a high altitude platform, or a terrestrial station.
+The receiver is synchronized with the transmitter and knows the emitted symbols: for a monostatic configuration, this is readily verified [12], whereas, for a bistatic configuration, this requires some coordination [^3] When the radar receiver is on a different LEO satellite, the coordination may be established via an inter-satellite link.
+or a separate reference channel [36, 37, 38].
+For illustration, in this study we consider the bistatic configuration depicted in Fig.
 1.
-In this setup, the bistatic range of a target is defined as the total distance traveled by the probing signal (first from the communication transmitter to the target, and then from the target to the radar receiver).
-Correspondingly, the bistatic range-rate refers to the time derivative of the bistatic range, i.e., the rate at which the bistatic range changes due to motion of the transmitter, the target, or the receiver.
-It is important to note that the following analysis also applies to the monostatic case, where the transmitter and receiver are co-located; in such a configuration, the bistatic range and range-rate simplify to twice the monostatic range and range-rate, respectively.
-To streamline the exposition, we will omit the adjective "bistatic" from this point forward and use the terms range and range-rate to refer to the bistatic quantities unless otherwise specified.
-Finally, we begin by assuming the presence of a single target to simplify the initial discussion.
+We assume point-like targets, each generating a single line-of-sight propagation path.
+The point-like model is appropriate when the target extent is small relative to the range resolution and allows us to isolate and analyze the effects of range migration.
+The single-path model is consistent with the considered space-sensing scenario, where secondary scattering components are typically negligible or absent.
+Extensions to distributed targets or to more general propagation conditions involving multiple echoes from the same target are left for future work.
+For clarity of exposition, we initially focus on a single target, while multiple targets are addressed in Sec.
+IV-B.
 Figure 1: System geometry.
 II-A Transmit signal
-Consider an OTFS frame with bandwidth $M\Delta$ and duration $NT$, where $M$ and $N$ are the number of subcarriers and symbol intervals, respectively, $\Delta$ is the subcarrier spacing, and $T=1/\Delta$ is the symbol interval, while $1/(M\Delta)$ and $1/(NT)$ are the delay and Doppler resolutions, respectively.
-Up to a positive scaling factor accounting for the radiated energy, the baseband signal emitted by the communication transmitter is
+Consider an OTFS frame with bandwidth $M\Delta$ and duration $NT$, where $M$ and $N$ are the number of subcarriers and symbol intervals, respectively, $\Delta$ is the subcarrier spacing, and $T=1/\Delta$ is the symbol interval.
+Accordingly, the corresponding range and range-rate resolutions are [36]
+$R_{\rm r}$ $=c/(M\Delta),$ (1a)
+$R_{\rm rr}$ $=\lambda/(NT),$ (1b)
+[…]
 $x(t)=\sum_{n=0}^{N-1}\underbrace{\sum_{m=0}^{M-1}X_{\rm TF}[n,m]g_{\rm tx}\!\left(t-nT\right)\mathrm{e}^{\mathrm{i}2\pi m\Delta(t-nT)}}_{x_{n}(t)},$ (2)
@@
 By considering $g_{\rm tx}(t)$ time limited in $[0,T]$, then $x(t)$ has support $[0,NT]$.
-Also, by assuming the spectrum of $g_{\rm tx}(t)$ contained in $\Delta/2,\Delta/2]$, then the spectrum of $x(t)$ is contained in $\Delta/2,M\Delta-\Delta/2]$.
-[^4] Since a time limited signal cannot be strictly band limited, some spectral leakage outside $\Delta/2,M\Delta-\Delta/2]$ is present.
+Also, by assuming the spectrum of $g_{\rm tx}(t)$ is contained in $0,\Delta]$, then the spectrum of $x(t)$ is contained in [^5] Since a time-limited signal cannot be strictly band-limited, some spectral leakage outside $0,M\Delta]$ is present.
 In practice, the transmit shaping filter may not be exactly zero outside $[0,T]$ to reduce such spectral leakage; in this latter case, $x(t)$ is only approximately contained in $[0,NT]$.
-II-B Design assumptions
+$0,M\Delta]$.
+II-B Target kinematics and design assumptions
 During an OTFS frame, the target, the communication transmitter, and the radar receiver can be assumed to move with a constant velocity.
@@
 $\bar{r}(t)=\bar{d}-\bar{v}t,$ (5)
-where $\bar{d}$ and $\bar{v}$ are the initial range and the range-rate, respectively, defined as
+where $\bar{d}$ and $\bar{v}$ are the initial-range and the range-rate, respectively, defined as
 $\bar{d}$ $=\left\|\bm{p}-\bm{p}_{\rm tx}\right\|+\left\|\bm{p}-\bm{p}_{\rm rx}\right\|,$ (6a)
 $\bar{v}$ $=\frac{\left(\bm{p}-\bm{p}_{\rm tx}\right)^{\mathsf{T}}\!\left(\bm{v}_{\rm tx}-\bm{v}\right)}{\left\|\bm{p}-\bm{p}_{\rm tx}\right\|}+\frac{\left(\bm{p}-\bm{p}_{\rm rx}\right)^{\mathsf{T}}\!\left(\bm{v}_{\rm rx}-\bm{v}\right)}{\left\|\bm{p}-\bm{p}_{\rm rx}\right\|}.$ (6b)
-Denote by $\bar{r}_{\min}$ and $\bar{r}_{\max}$ the minimum and maximum ranges, respectively, and by $\bar{v}_{\min}$ and $\bar{v}_{\max}$ the minimum and maximum range-rates, respectively.
+Hereafter, we make the following design assumptions.
+II-B1 Unambiguous initial-range and range-rate intervals
+Denote by $\bar{r}_{\min}$ and $\bar{r}_{\max}$ the minimum and maximum initial-ranges, respectively, and by $\bar{v}_{\min}$ and $\bar{v}_{\max}$ the minimum and maximum range-rates, respectively.
 These values are tied to the prior uncertainty on the target location and mobility, as resulting from a previous detection phase (if the radar is operating in search mode) or as predicted by the tracking stage (if the radar is operating in tracking mode), and are assumed known.
-Also, let $r_{\max}=\bar{r}_{\max}-\bar{r}_{\min}$ and $v_{\max}=\bar{v}_{\max}-\bar{v}_{\min}$ be the length of the inspected range interval and range-rate interval, respectively.
+Also, let $r_{\max}=\bar{r}_{\max}-\bar{r}_{\min}$ and $v_{\max}=\bar{v}_{\max}-\bar{v}_{\min}$ be the length of the inspected initial-range and range-rate intervals, respectively.
 Then, we assume
@@
 $\nu_{\max}$ $=\frac{v_{\max}}{\lambda}<\Delta,$ (7b)
-so as to avoid ambiguity in the estimation of the initial range and the range-rate of the target, respectively.
-Finally, let $N$ be a positive integer multiple of $B$.
-Then, we assume that $\bar{r}(t)$ is constant over at least $N/B$ consecutive symbol intervals: we refer to this as the $B$-fold stop-and-go approximation on the range variation over the OTFS frame.
-Under this design assumption, we can express:
+so as to avoid ambiguity in the estimation of the initial-range and the range-rate of the target, respectively.
+[^6] For the parameters adopted in Sec.
+V ($\Delta=15$ kHz, $M=512$, $N=128$, carrier frequency $4$ GHz), we have $c/\Delta\approx 19.9$ km and $\lambda\Delta\approx 1.12$ km/s, respectively.
+By varying $\bar{r}_{\min}$ and $\bar{v}_{\min}$, any initial-range interval of length $r_{\max}<19.9$ km and any range-rate interval of length $v_{\max}<1.12$ km/s can be jointly inspected without ambiguity.
+II-B2 Blockwise constant-range approximation
+Let $B$ be a divisor of $N$.
+Then, the OTFS frame can be partitioned into $B$ equal blocks of $N/B$ symbol intervals.
+For receiver design, the target range is assumed constant within each block.
+Accordingly, we write
 $\bar{r}(t)=\bar{r}(bT_{B}),\quad\forall t\in[bT_{B},(b+1)T_{B}),$ (8)
-with $T_{B}=NT/B$ and $b=0,\ldots,B-1$.
-For a fixed system bandwidth, increasing $B$ improves the model’s accuracy, particularly when the target exhibits higher range-rates or the OTFS frame duration is long (see Sec.
-VI for further discussion).
-Notably, prior studies assume $B=1$ at the design stage, effectively neglecting any range migration across the entire OTFS frame.
-In contrast, setting $B=N$ limits this approximation to only a single symbol interval, thereby capturing range migration with a much finer granularity.
+where $T_{B}=NT/B$ and $b=0,\ldots,B-1$.
+Increasing $B$ improves the model accuracy, particularly when the target exhibits higher range-rates or when the OTFS frame duration is long (see Sec.
+V for further discussion).
+A practical design guideline is to select $B$ such that the range variation over each block, i.e., $\bar{v}_{\max}NT/B$, is significantly smaller than the range resolution $R_{\rm r}$ (see also Sec.
+V).
+Prior studies assume $B=1$, which is appropriate only for applications in which target-range migration within the OTFS frame can be neglected.
 II-C Received signal
@@
 is the baseband representation of the target echo with respect to the reference frequency $\frac{c+\bar{v}_{\min}}{\lambda}$ and $r(t)=\bar{r}(t)-\bar{r}_{\min}$ is the excess target range (i.e., the target range in excess with respect to $\bar{r}_{\min}$).
-The support of $y(t)$ is contained in the time interval $0,NT+r_{\max}/c]$, and its spectrum is negligible outside the frequency interval $\Delta/2,M\Delta+v_{\max}/\lambda-\Delta/2]$.
-Define the excess initial range $d=\bar{d}-\bar{r}_{\min}\in[0,r_{\max}]$, the excess range-rate $v=\bar{v}-\bar{v}_{\min}\in[0,v_{\max}]$, and $\alpha=\bar{\alpha}\mathrm{e}^{-\mathrm{i}2\pi\frac{d}{\lambda}}$.
-If the variation of the target range in a symbol interval can be ignored (which is the case under the $B$-fold stop-and-go approximation), then (10) can be expanded as
+Note that the support of $y(t)$ is contained in the time interval $0,NT+r_{\max}/c]$, and its spectrum is negligible outside the frequency interval $0,M\Delta+v_{\max}/\lambda]$.
+Define the excess initial-range $d=\bar{d}-\bar{r}_{\min}\in[0,r_{\max}]$, the excess range-rate $v=\bar{v}-\bar{v}_{\min}\in[0,v_{\max}]$, and $\alpha=\bar{\alpha}\mathrm{e}^{-\mathrm{i}2\pi\frac{d}{\lambda}}$.
+Ignoring the variation of the target range in a symbol interval, then (11) can be expanded as
 $y(t)=\sum_{n=0}^{N-1}y_{n}(t),$ (12)
@@
 $h_{n}(\tau,\nu)=\alpha\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(nT\right)}{c}}\delta\left(\tau-\frac{r\left(nT\right)}{c}\right)\delta\left(\nu-\frac{v}{\lambda}\right).$ (14)
-Notice here that $y_{n}(t)$ is the output of a doubly-selective channel with delay-Doppler spread function $h_{n}(\tau,\nu)$ when the input is $x_{n}(t)$; the spread function depends upon $n$, as the target range can change during the OTFS frame.
+Notice that $y_{n}(t)$ is the output of a doubly-selective channel with delay-Doppler spread function $h_{n}(\tau,\nu)$ when the input is $x_{n}(t)$; the spread function depends upon $n$, as the target range can change during the OTFS frame.
 II-D OTFS demodulation
-The radar receiver is implemented on a device than relies on its OTFS demodulator to obtain discrete samples of the received signal $y(t)$ in the delay-Doppler domain; such samples are then processed to estimate the target parameters as shown in Sec.
-V.
-The OTFS demodulator first transform $y(t)$ into the time-frequency domain via the Wigner Transform; in particular, the following measurements are obtained
+The radar receiver relies on an OTFS demodulator to obtain discrete samples of the received signal in the delay-Doppler domain.
+Such samples are then processed to estimate the target parameters as shown in Sec.
+IV.
+First, the OTFS demodulator transforms $y(t)$ into the time-frequency domain via the Wigner Transform; in particular, the following measurements are obtained
 $Y_{\rm TF}[n,m]=\int_{-\infty}^{\infty}y(t)g_{\rm rx}^{*}(t-nT)\mathrm{e}^{-\mathrm{i}2\pi m\Delta(t-nT)}dt,$ (15)
@@
 $\gamma(\tau,\nu)=\int_{-\infty}^{\infty}g_{\rm tx}(\beta)g^{*}_{\rm rx}(\beta-\tau)\mathrm{e}^{-\mathrm{i}2\pi\nu(\beta-\tau)}d\beta$ (16)
-the cross-ambiguity function of $g_{\rm tx}(t)$ and $g_{\rm rx}(t)$; then, the following proposition establishes a connection between $\{Y_{\rm TF}[n,m]\}$ and $\{X_{\rm TF}[n,m]\}$ that generalizes the result in [8, Theorem 1] to the case where the variation of the target range is ignored only within each symbol interval.
+the cross-ambiguity function of $g_{\rm tx}(t)$ and $g_{\rm rx}(t)$; then, the following proposition establishes a connection between $\{Y_{\rm TF}[n,m]\}$ and $\{X_{\rm TF}[n,m]\}$ that generalizes the previous result in [6, Theorem 1].
+The proof is reported in Appendix A.
 Proposition 1.
@@
 $H_{n,m}[n^{\prime},m^{\prime}]=\alpha\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}nT}\mathrm{e}^{-\mathrm{i}2\pi m^{\prime}\Delta\frac{r\left(n^{\prime}T\right)}{c}}\\ \times\gamma\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c},(m-m^{\prime})\Delta-\frac{v}{\lambda}\right).$ (18)
-Proof.
-Upon plugging (11) into (14), we have
-$Y_{\rm TF}[n,m]=\sum_{n^{\prime}=0}^{N-1}\sum_{m^{\prime}=0}^{M-1}\alpha X_{\rm TF}[n^{\prime},m^{\prime}]$
-$\quad\times\int_{-\infty}^{\infty}g_{\rm tx}\!\left(t-n^{\prime}T-\frac{r\left(n^{\prime}T\right)}{c}\right)g^{*}_{\rm rx}\!\left(t-nT\right)$
-$\quad\times\mathrm{e}^{\mathrm{i}2\pi m^{\prime}\Delta\left(t-n^{\prime}T-\frac{r\left(n^{\prime}T\right)}{c}\right)}\mathrm{e}^{-\mathrm{i}2\pi m\Delta\left(t-nT\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}t}dt$
-$=\sum_{n^{\prime}=0}^{N-1}\sum_{m^{\prime}=0}^{M-1}X_{\rm TF}[n^{\prime},m^{\prime}]$
-$\quad\times\alpha\mathrm{e}^{\mathrm{i}2\pi m\Delta\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c}\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\left(n^{\prime}T+\frac{r\left(n^{\prime}T\right)}{c}\right)}$
-$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c}\right)}$
-$\quad\times\int_{-\infty}^{\infty}g_{\rm tx}\!\left(\beta\right)g^{*}_{\rm rx}\!\left(\beta-(n-n^{\prime})T+\frac{r\left(n^{\prime}T\right)}{c}\right)$
-$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left(\beta-(n-n^{\prime})T+\frac{r\left(n^{\prime}T\right)}{c}\right)}d\beta$
-$=\sum_{n^{\prime}=0}^{N-1}\sum_{m^{\prime}=0}^{M-1}X_{\rm TF}[n^{\prime},m^{\prime}]\alpha\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}nT}\mathrm{e}^{-\mathrm{i}2\pi m^{\prime}\Delta\frac{r\left(n^{\prime}T\right)}{c}}$
-$\quad\times\gamma\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c},(m-m^{\prime})\Delta-\frac{v}{\lambda}\right),$
-which completes the proof. ∎
-Finally, the OTFS demodulator transforms $Y_{\rm TF}[n,m]$ into the delay-Doppler domain via the Symplectic Finite Fourier Transform; in particular, we have
+Next, the OTFS demodulator transforms $Y_{\rm TF}[n,m]$ into the delay-Doppler domain via the Symplectic Finite Fourier Transform; in particular, we have
 $Y_{\rm DD}[k,l]=\frac{1}{\sqrt{NM}}\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}Y_{\rm TF}[n,m]\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)},$ (19)
 for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$.
-In the Secs.
-III and IV we will establish an explicit relationship between $\{Y_{\rm DD}[n,m]\}$ and $\{X_{\rm DD}[n,m]\}$ for the cases where ideal and rectangular shaping filters are employed, respectively.
-III Ideal shaping filters
-Consider ideal shaping filters satisfying the bi-orthogonality condition [8, Sec.
+III Delay–Doppler Input–Output Model
+We establish here a relationship between $\{Y_{\rm DD}[k,l]\}$ and $\{X_{\rm DD}[k,l]\}$ for two representative shaping filters, namely ideal and rectangular filters.
+Ideal shaping filters lead to a simplified input–output relationship, and the resulting analysis provides useful insights into the impact of range migration and serves as a benchmark.
+In contrast, the more practical rectangular shaping filters yield a more involved relationship.
+III-A Ideal shaping filters
+Ideal shaping filters satisfies the bi-orthogonality condition [6, Sec.
 III].
-Under this assumption, the ambiguity function satisfies $\gamma(\tau,\nu)=1$ for $(\tau,\nu)\in[-\tau_{\max},\tau_{\max}]\times[-\nu_{\max},+\nu_{\max}]$ and $\gamma(\tau,\nu)=0$ for $(\tau,\nu)\in[nT-\tau_{\max},nT+\tau_{\max}]\times[m\Delta-\nu_{\max},m\Delta+\nu_{\max}]$ for all pairs of integers $(m,n)\neq(0,0)$.
-Although practical waveforms can only approximate this condition to a certain extent, the use of ideal shaping filters offers valuable design insights, and the resulting system performance serves as a benchmark for comparison with that achievable using rectangular shaping filters.
-In this case, (16) becomes
+In this case, the ambiguity function satisfies $\gamma(\tau,\nu)=1$ for $(\tau,\nu)\in[-\tau_{\max},\tau_{\max}]\times[-\nu_{\max},\nu_{\max}]$, and $\gamma(\tau,\nu)=0$ for $(\tau,\nu)\in[nT-\tau_{\max},nT+\tau_{\max}]\times[m\Delta-\nu_{\max},m\Delta+\nu_{\max}]$ for all integer pairs $(m,n)\neq(0,0)$.
+Consequently, (17) becomes
 $Y_{\rm TF}[n,m]=H_{n,m}[n,m]X_{\rm TF}[n,m],$ (20)
@@
 $H_{n,m}[n,m]=\alpha\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}nT}\mathrm{e}^{-\mathrm{i}2\pi\frac{r\left(nT\right)}{c}m\Delta}.$ (21)
-Also, denote by
+To proceed, denote by
 $\mathcal{D}_{Q}(\nu)=\frac{1}{Q}\sum_{q=0}^{Q-1}\mathrm{e}^{-\mathrm{i}2\pi\nu q}=\frac{\mathrm{e}^{-\mathrm{i}2\pi\nu Q}-1}{Q\left(\mathrm{e}^{-\mathrm{i}2\pi\nu}-1\right)}$ (22)
-the periodic sinc function (also called Dirichlet function) of degree $Q$.
-Then, the following proposition establishes a connection between $\{Y_{\rm DD}[k,l]\}$ and $\{X_{\rm DD}[k,l]\}$ that generalizes the result in [8, Proposition 2] to the case where the target range can change every $N/B$ symbol intervals.
+the periodic $\mathrm{sinc}$ function (also called Dirichlet function) [^7] Recall that $\mathcal{D}_{Q}(\nu)$ is a periodic function of $\nu$ with period one and is zero at multiple of $1/Q$, except for $\nu=0,\pm 1,\pm 2,\ldots$ where it is one; also, $|\mathcal{D}_{Q}(\nu)|$ has a main-lobe centered at 0 of width $2/Q$ (between zero crossings) and Q2Q-2 sidelobes of width $1/Q$ (between zero crossings) in [0.5,0.5][-0.5,0.5].
+The first sidelobe is approximately $13$ dB down from the main-lobe, and the subsequent sidelobes fall off at about $6$ dB per octave [39].
+of degree $Q$.
+Then, the following result extends [6, Proposition 2] to the case of target range variations every $N/B$ symbol intervals.
+The proof is reported in Appendix B.
 Proposition 2.
-Assume that ideal shaping filters are employed. Then, under the $B$-fold stop-and-go approximation on the range variation over the OFTS frame, we have
+Assume that ideal shaping filters are employed. Under the blockwise constant-range approximation, we have
 $Y_{\rm DD}[k,l]=\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\!X_{\rm DD}[k^{\prime},l^{\prime}]\Phi[k-k^{\prime},l-l^{\prime}],$ (23)
@@
 $\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (24)
-Proof.
-After exploiting (2) and (19), Eq. (18) becomes
-$Y_{\rm DD}[k,l]=\frac{1}{NM}\!\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\!X_{\rm DD}[k^{\prime},l^{\prime}]\\ \times\underbrace{\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}H_{n,m}[n,m]\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}\right)n}\mathrm{e}^{\mathrm{i}2\pi\left(\frac{l-l^{\prime}}{M}\right)m}}_{(\star)}.$ (24)
-Finally, notice that we have
-$(\star)$ $=M\alpha\sum_{n=0}^{N-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}$
-$\quad\times\left[\frac{1}{M}\sum_{m=0}^{M-1}\mathrm{e}^{\mathrm{i}2\pi\left(\frac{l-l^{\prime}}{M}-\frac{r\left(nT\right)\Delta}{c}\right)m}\right]$
-$=M\alpha\sum_{n=0}^{N-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}$
-$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(nT\right)\Delta}{c}\right)\right)$
-$=NM\alpha\sum_{b=0}^{B-1}\frac{1}{B}\left[\frac{B}{N}\sum_{n=bN/B}^{(b+1)N/B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}\right]$
-$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right)$ (25)
-$=NM\alpha\Phi[k-k^{\prime},l-l^{\prime}],$
-where (25) follows from (7). ∎
-III-A Some insights on Proposition 2
-Recall that $\mathcal{D}_{Q}(\nu)$ is a periodic function of $\nu$ with period one and is zero at multiple of $1/Q$, except for $\nu=0,\pm 1,\pm 2,\ldots$ where it is one; also, $|\mathcal{D}_{Q}(\nu)|$ has a main-lobe centered at 0 of width $2/Q$ (between zero crossings) and Q2Q-2 sidelobes of width $1/Q$ (between zero crossings) in [0.5,0.5][-0.5,0.5].
-The first sidelobe is approximately $13$ dB down from the main-lobe, and the subsequent sidelobes fall off at about $6$ dB per octave [44]; following [8], for large $Q$, we can assume $|\mathcal{D}_{Q}(\nu)|\approx 0$ for $\nu\notin[-D/Q,D/Q]$, with $0<2D+1\ll Q$, where the non-negative integer $D$ is a design parameter that depends upon the desired approximation error.
-Accordingly, the following remarks are in order.
+The following remarks and example provide further insight into Proposition 2.
 Remark 1.
-The sequence $\Phi[k,l]$ in (23) is a periodic function of $k$ and $l$ with period $N$ and $M$, respectively. This implies that (22) is a two-dimensional circular convolution, whereby
+$\Phi[k,l]$ is periodic in $k$ and $l$ with period $N$ and $M$, respectively. This implies that (23) is a two-dimensional circular convolution, which can be rewritten as
 $Y_{\rm DD}[k,l]$ $=\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\Phi\Big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\Big]$ (25a)
-$=\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\!X_{\rm DD}\Big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\Big]\Phi[k^{\prime},l^{\prime}],$ (26b)
-for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$. ∎
+$=\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}\Big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\Big]\Phi[k^{\prime},l^{\prime}].$ (25b)
+Here, $\Phi[k,l]$ is the convolution kernel relating the transmitted and received delay-Doppler symbols.
+Accordingly, $\alpha\Phi[k,l]$ represents the target response in the delay-Doppler domain, and $\Phi[k,l]$ can be interpreted as the associated point-spread function, specified by the target initial range and range-rate.
 Remark 2.
-For large $N/B$ and $M$, the sums in (26) only contain few dominant terms.
-To illustrate this point, denote by $R_{\rm r}=c/(M\Delta)$ and $R_{\rm rr}=\lambda/(NT)$ the range and range-rate resolutions [41], respectively, and write the range at time $bT_{B}$ and the range-rate as
+Proposition 2 subsumes [6, Proposition 2] as a special case. To illustrate this point, first use (1) to write the range at time $bT_{B}$ and the range-rate as
 $r\left(bT_{B}\right)$ $=(l_{b,{\rm int}}+l_{b,{\rm fra}})R_{\rm r},$ (26a)
 $v$ $=(k_{\rm int}+k_{\rm fra})R_{\rm rr},$ (26b)
-for $b=0,\ldots,B-1$, where $l_{b,{\rm int}}\in\{0,1,\ldots,M-1\}$, $k_{\rm int}\in\{0,1,\ldots,N-1\}$, and $l_{b,{\rm fra}},\,k_{\rm fra}\in(-0.5,0.5]$. Then, from (23) and (27) we have
+where $l_{b,{\rm int}}\in\{0,1,\ldots,M-1\}$, $k_{\rm int}\in\{0,1,\ldots,N-1\}$, and $l_{b,{\rm fra}},\,k_{\rm fra}\in(-0.5,0.5]$.
+Next, assume no range migration, whereby $l_{b,{\rm int}}=l_{{\rm int}}$ and $l_{b,{\rm fra}}=l_{{\rm fra}}$ for $b=0,\ldots,B-1$.
+Upon plugging (26) in (25b), we obtain
+$Y_{\rm DD}[k,l]=\alpha\sum_{\bar{k}=-\lfloor N/2\rfloor}^{\lceil N/2\rceil-1}\sum_{\bar{l}=-\lfloor M/2\rfloor}^{\lceil M/2\rceil-1}\mathcal{D}_{N}\!\left(-\frac{\bar{k}+k_{\rm fra}}{N}\right)\\ \times\mathcal{D}_{M}\!\left(\frac{\bar{l}+l_{\rm fra}}{M}\right)X_{\rm DD}\Big[[k-k_{\rm int}+\bar{k}]_{N},[l-l_{\rm int}+\bar{l}]_{M}\Big],$ (27)
+consistently with [6, Proposition 2]. In addition, if $k_{\rm fra}=l_{\rm fra}=0$, then (27) becomes
+$Y_{\rm DD}[k,l]$ $=\alpha X_{\rm DD}\Big[[k-k_{\rm int}]_{N},[l-l_{{\rm int}}]_{M}\Big].$ (28)
+Remark 3.
+For large $N/B$ and $M$, the sum in (25) contains only a few dominant terms. Indeed, by substituting (26) into (24), we obtain
 $\Phi[k,l]$ $=\frac{1}{B}\sum_{b=0}^{B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k_{\rm int}-k_{\rm fra}}{B}\right)b}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k_{\rm int}-k_{\rm fra}}{N}\right)$
 $\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l_{b,{\rm int}}-l_{b,{\rm fra}}}{M}\right)\right).$ (29)
-Finally, from the properties of the Dirichlet function discussed in Sec.
-III-A, we have $|\Phi[k,l]|\approx 0$ if $k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$, for $b=0,\ldots,B-1$: more on this in Remark 4.
-∎
-Remark 3.
-Assume that $l_{b,{\rm int}}=l_{{\rm int}}$ and $l_{b,{\rm fra}}=l_{{\rm fra}}$ for $b=0,\ldots,B-1$, which means that there is no range migration and that the target range is an integer multiple of the range resolution.
-Then, upon plugging (28) in (26b) and after straightforward manipulations, we obtain
-$Y_{\rm DD}[k,l]=\alpha\sum_{\bar{k}=-\lfloor N/2\rfloor}^{\lceil N/2\rceil-1}\sum_{\bar{l}=-\lfloor M/2\rfloor}^{\lceil M/2\rceil-1}\mathcal{D}_{N}\!\left(-\frac{\bar{k}+k_{\rm fra}}{N}\right)\\ \times\mathcal{D}_{M}\!\left(\frac{\bar{l}+l_{\rm fra}}{M}\right)X_{\rm DD}\Big[[k-k_{\rm int}+\bar{k}]_{N},[l-l_{\rm int}+\bar{l}]_{M}\Big],$ (29)
-consistently with [8, Proposition 2]. In addition, if $k_{\rm fra}=l_{\rm fra}=0$, then (29) becomes
-$Y_{\rm DD}[k,l]$ $=\alpha X_{\rm DD}\Big[[k-k_{\rm int}]_{N},[l-l_{{\rm int}}]_{M}\Big].$ (30)
-Remark 4.
+For large $Q$, $|\mathcal{D}_{Q}(\nu)|\approx 0$ for $\nu\notin[-D/Q,D/Q]$, with $0<2D+1\ll Q$, where the non-negative integer $D$ controls the approximation error [6].
+Accordingly, $|\Phi[k,l]|\approx 0$ if $k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$.
 Refer to caption    Refer to caption    Refer to caption
-Figure 2: $|\Phi[k,l]|$ vs $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ for Cases 1, 2, and 3, when $B=4$, $N=M=64$, and ideal shaping filters are employed.
-The two-dimensional sequence $\alpha\Phi[k,l]$ represent the target response in the delay-Doppler domain.
-For a given $B$, $\Phi[k,l]$ is specified by the initial range and the range-rate of the target and satisfy $|\Phi[k,l]|\leq 1$; to get more insights on its structure, we discuss a toy example with $B=4$ and $N=M=64$.
+Figure 2: $|\Phi[k,l]|$ vs $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$, when $B=4$, $N=M=64$, $l_{0,{\rm int}}=42$, $k_{\rm int}=32$, and ideal shaping filters.
+Case 1 (left): $l_{1,{\rm int}}=l_{2,{\rm int}}=l_{3,{\rm int}}=42$ and $l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=k_{\rm fra}=0$.
+Case 2 (center): $l_{1,{\rm int}}=l_{2,{\rm int}}=l_{3,{\rm int}}=42$ and $l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=k_{\rm fra}=0.5$.
+Case 3 (right): $l_{1,{\rm int}}=43$, $l_{2,{\rm int}}=44$, $l_{3,{\rm int}}=45$, and $l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=k_{\rm fra}=0.5$.
+Example 1.
+To provide physical insight into the effect of range migration on the target point-spread function in (24), we consider a toy example with $B=4$, $N=M=64$, $l_{0,{\rm int}}=42$, and $k_{\rm int}=32$.
 Fig.
-2 shows $|\Phi[k,l]|$ versus $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ in different cases:
-  1. 1.
-    $k_{\rm fra}=l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=0$, $k_{\rm int}=32$, and $l_{0,{\rm int}}=\cdots=l_{3,{\rm int}}=42$;
-  2. 2.
-    $k_{\rm fra}=l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=0.5$, $k_{\rm int}=32$, and $l_{0,{\rm int}}=\cdots=l_{3,{\rm int}}=42$;
-  3. 3.
-    $k_{\rm fra}=l_{0,{\rm fra}}=\cdots=l_{B-1,{\rm fra}}=0.5$, $k_{\rm int}=50$, $l_{0,{\rm int}}=42$, $l_{1,{\rm int}}=43$, $l_{2,{\rm int}}=44$, and $l_{3,{\rm int}}=45$.
-In Case 1, the range of the target is not changing in the OTFS frame; also, the range and range-rate of the target are multiple of $R_{\rm r}$ and $R_{\rm rr}$, respectively.
-In this case, $|\Phi[k,l]|$ only contains one non-zero entry, whose location depends upon the range and range-rate of the target, in keeping with the analytical model in (30).
-By comparing Cases 1 and 2, it is seen that a fractional range and a fractional range-rate cause a spread of $|\Phi[k,l]|$ along both domains, in keeping with (29).
-By comparing Cases 2 and 3, it seen that a variation of the target range every $N/B$ symbol intervals causes an additional spread of $|\Phi[k,l]|$ in both domains: this is the novel result established by Proposition 2.
-Finally, notice that most elements of $|\Phi[k,l]|$ are much smaller than one, in keeping with Remark 2; hence, the response of a high-speed target still has a sparse structure in the delay-Doppler domain, determined entirely by the initial range and range-rate.
-∎
-III-B Vectorized model
-The samples $\{Y_{\rm DD}[k,l]\}$ can be organized into the vector $\bm{y}\in\mathbb{C}^{NM}$, where $(\bm{y})_{lN+k+1}=Y_{\rm DD}[k,l]$, for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$.
-In the following, we refer to $\bm{y}$ as the target echo.
-From (26a) we have
+2 reports $|\Phi[k,l]|$ versus $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ for different target kinematics.
+In Case 1, we assume no range migration (i.e., $l_{1,{\rm int}}=l_{2,{\rm int}}=l_{3,{\rm int}}=42$), integer range (i.e., $l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=0$), and integer range-rate (i.e., $k_{\rm fra}=0$).
+It is seen that $|\Phi[k,l]|$ contains a single non-zero entry at $(k,l)=(32,42)$, as predicted by (28).
+In Case 2, we assume no range migration, fractional range with $l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=0.5$, and fractional range-rate with $k_{\rm fra}=0.5$.
+By comparing Cases 1 and 2, it is observed that fractional range and range-rate induce a spreading of $|\Phi[k,l]|$ along both the delay and Doppler dimensions, as predicted by (27).
+In Case 3, we assume that the target range varies four times over the OTFS frame, with $l_{1,{\rm int}}=43$, $l_{2,{\rm int}}=44$, $l_{3,{\rm int}}=45$ and $k_{\rm fra}=l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=0.5$.
+By comparing Cases 2 and 3, range migration produces an additional spreading of $|\Phi[k,l]|$, which constitutes the main novelty captured by Proposition 2.
+In all cases, most entries of $|\Phi[k,l]|$ remain significantly smaller than one, in agreement with Remark 3.
+Hence, the response of a high-speed target preserves a sparse structure, albeit with an expanded support.
+III-A1 Vectorized model
+The samples $\{Y_{\rm DD}[k,l]\}$ can be arranged into the vector $\bm{y}\in\mathbb{C}^{NM}$, where $(\bm{y})_{lN+k+1}=Y_{\rm DD}[k,l]$, for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$.
+From (25a), we obtain
 $\bm{y}=\alpha\bm{\Phi}\bm{x},$ (30)
-where $\bm{\Phi}$ is an $NM\times NM$ matrix with $(\bm{\Phi})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Phi\big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\big]$ and $\bm{x}$ is an $NM$-dimensional vector with $(\bm{x})_{lN+k+1}=X_{\rm DD}[k,l]$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$.
-Alternatively, from (26b) we have
+where $\bm{\Phi}\in\mathbb{C}^{NM\times NM}$ has entries $(\bm{\Phi})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Phi\big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\big]$ and $\bm{x}\in\mathbb{C}^{NM}$ is defined by $(\bm{x})_{lN+k+1}=X_{\rm DD}[k,l]$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$.
+Alternatively, from (25b), we have
 $\bm{y}=\alpha\bm{X}\bm{\phi},$ (31)
-where $\bm{X}$ is an $NM\times NM$ matrix with $(\bm{X})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=X_{\rm DD}\big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\big]$ and $\bm{\phi}$ is an $NM$-dimensional vector with $(\bm{\phi})_{lN+k+1}=\Phi[k,l]$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$.
-In the following, we refer to
-$\bm{e}(d,v,B)=\bm{\Phi}(d,v,B)\bm{x}=\bm{X}\bm{\phi}(d,v,B)$ (33)
-as the normalized target echo, where we have made explicit the dependence upon the initial range $d$ of the target, the range-rate $v$ of the target, and the parameter $B$.
-IV Rectangular shaping filters
-Assume here $g_{\rm tx}(t)=g_{\rm rx}(t)=\sqrt{1/T}\mathrm{rect}(t/T)$. Then, $\gamma(\tau,\nu)=0$ for $|\tau|\geq T$; also, for large $M$, recall that [8]
+where $\bm{X}\in\mathbb{C}^{NM\times NM}$ has entries $(\bm{X})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=X_{\rm DD}\big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\big]$ and $\bm{\phi}\in\mathbb{C}^{NM}$ is defined by $(\bm{\phi})_{lN+k+1}=\Phi[k,l]$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$.
+Since the matrix $\bm{\Phi}$ and the vector $\bm{\phi}$ depend on the target initial-range and range-rate, we will denote them by $\bm{\Phi}(d,v)$ and $\bm{\phi}(d,v)$ when we want to make explicit this dependence.
+III-B Rectangular shaping filters
+Assume now $g_{\rm tx}(t)=g_{\rm rx}(t)=\sqrt{1/T}\,\mathrm{rect}(t/T)$. This implies that $\gamma(\tau,\nu)=0$ for $|\tau|\geq T$. Moreover, for large $M$, we have [6]
 $\gamma(\tau,\nu)=\frac{T}{M}\sum_{p=0}^{M-1}g_{\rm tx}\left(\frac{pT}{M}\right)g^{*}_{\rm rx}\left(\frac{pT}{M}-\tau\right)\mathrm{e}^{-\mathrm{i}2\pi\nu\left(\frac{pT}{M}-\tau\right)}\\ =\begin{cases}\displaystyle\frac{1}{M}\sum_{p=0}^{\lceil M+\tau M/T\rceil-1}\!\!\mathrm{e}^{-\mathrm{i}2\pi\nu T\left(\frac{p}{M}-\frac{\tau}{T}\right)},&\tau\in(-T,0),\\ \displaystyle\frac{1}{M}\sum_{p=\lceil\tau M/T\rceil}^{M-1}\!\!\mathrm{e}^{-\mathrm{i}2\pi\nu T\left(\frac{p}{M}-\frac{\tau}{T}\right)},&\tau\in[0,T).\end{cases}$ (32)
-To proceed, notice that (16) becomes here
+Consequently, (17) simplifies as follows. For $n=1,\ldots,N-1$ and $m=0,\ldots,M-1$, we have
 $Y_{\rm TF}[n,m]$ $=H_{n,m}[n,m]X_{\rm TF}[n,m]$
@@
 $\quad+\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n-1,m^{\prime}]X_{\rm TF}[n-1,m^{\prime}],$ (33)
-for $n=1,\ldots,N-1$ and $m=0,\ldots,M-1$, where the second and third terms on the right hand side are the inter carrier interference (ICI) and inter symbol interference (ISI) [8], respectively.
-For $n=0$ no ISI is present, as a guard interval is inserted between consecutive OTFS frames, and (16) becomes
+where the second and third terms on the right-hand side correspond to ICI and ISI, respectively [6].
+For $n=0$, instead, no ISI is present due to the guard interval between consecutive OTFS frames, and (17) reduces to
 $Y_{\rm TF}[0,m]$ $=H_{0,m}[0,m]X_{\rm TF}[0,m]$
@@
 for $m=0,\ldots,M-1$.
-Then, the following proposition establishes a connection between $\{Y_{\rm DD}[k,l]\}$ and $\{X_{\rm DD}[k,l]\}$ that generalizes the result in [8, Theorem 2] to the case where the target range is not a multiple of $R_{\rm r}$ and can change every $N/B$ symbol intervals; the proof is in the Appendix.
+The following result extends [6, Theorem 2] to the case where the target range is not an integer multiple of $R_{\rm r}$ and may vary every $N/B$ symbol intervals. The proof is provided in Appendix C.
 Proposition 3.
-Assume rectangular shaping filters. Then, under the $B$-fold stop-and-go approximation on the range variation over the OFTS frame, we have
+Assume rectangular-shaped filters. Under the blockwise constant-range approximation, we have
 $Y_{\rm DD}[k,l]=\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\!X_{\rm DD}[k^{\prime},l^{\prime}]\Psi_{k^{\prime},l^{\prime}}[k-k^{\prime},l-l^{\prime}],$ (35)
@@
 if $\big\lceil M-\frac{r\left(bT_{B}\right)M}{cT}\big\rceil\leq l^{\prime}\leq M-1$.
-The following remarks on Proposition 3 are now in order.
+The following remarks and example provide further insight into Proposition 3.
 Remark 4.
-The summations in (37) cannot be regarded as a two-dimensial convolution, since $\Psi_{k^{\prime},l^{\prime}}[k,l]$ depends upon $k^{\prime}$ and $l^{\prime}$: this is in sharp contrast with the result obtained for ideal shaping filters in Remark 1.
-Notice in passing that, for fixed $k^{\prime}$ and $l^{\prime}$, $\Psi_{k^{\prime},l^{\prime}}[k,l]$ is a periodic function of $k$ and $l$ with period $N$ and $M$, respectively, while, for fixed $k$ and $l$, it is a periodic function of $k^{\prime}$ with period $N$.
-∎
+The right-hand side of (35) is not a two-dimensional convolution, since $\Psi_{k^{\prime},l^{\prime}}[k,l]$ depends on $k^{\prime}$ and $l^{\prime}$ due to ISI and ICI.
+This is in sharp contrast with the ideal shaping filter case.
+For fixed $k^{\prime}$ and $l^{\prime}$, $\Psi_{k^{\prime},l^{\prime}}[k,l]$ acts as a location-dependent kernel relating the input symbol $X_{\rm DD}[k^{\prime},l^{\prime}]$ to the output symbols $Y_{\rm DD}[k,l]$, and is periodic in $k$ and $l$ with period $N$ and $M$, respectively.
+Accordingly, $\alpha\Psi_{k^{\prime},l^{\prime}}[k,l]$ represents the target response in the delay-Doppler domain to an input symbol located at $(k^{\prime},l^{\prime})$, and $\Psi_{k^{\prime},l^{\prime}}[k,l]$ can be interpreted as a location-dependent point-spread function, specified by the target initial range and range-rate.
 Remark 5.
-For large $N/B$ and $M$, the summations in (37) only contain few dominant dominant terms; indeed, from the properties of the Dirichlet function discussed in Sec.
-III-A, we have $|\Psi_{k^{\prime},l^{\prime}}[k,l]|\approx 0$ if $k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$, for $b=0,\ldots,B-1$.
-This result is similar to that obtained for ideal shaping filters in Remark 2.
-∎
-Remark 7.
-Proposition 3 subsumes [8, Theorem 2] as a special case. To verify this, use (27) to recast (37) as
+By using (26), the result in (35) can be recast as
 $Y_{{\rm DD}}[k,l]=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{k^{\prime}=0}^{N-1}$
-$\quad\times\Bigg[\sum_{l^{\prime}=0}^{M-1-l_{b,{\rm int}}-\lfloor l_{b,{\rm fra}}\rfloor}X_{\rm DD}[k^{\prime},l^{\prime}]\Psi_{b,k^{\prime},l^{\prime}}[k-k^{\prime},l-l^{\prime}]$
+$\quad\times\Bigg[\sum_{l^{\prime}=0}^{M-l_{b,{\rm int}}-\lfloor l_{b,{\rm fra}}\rfloor-1}X_{\rm DD}[k^{\prime},l^{\prime}]\Psi_{b,k^{\prime},l^{\prime}}[k-k^{\prime},l-l^{\prime}]$
 $\quad+\sum_{l^{\prime}=M-l_{b,{\rm int}}-\lfloor l_{b,{\rm fra}}\rfloor}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\Psi_{b,k^{\prime},l^{\prime}}[k-k^{\prime},l-l^{\prime}]\Bigg]$
@@
 $\left[\mathcal{D}_{N}\!\left(-\frac{\bar{k}+k_{\rm fra}}{N}\right)-\frac{1}{N}\right]\mathrm{e}^{-\mathrm{i}2\pi\frac{[k-k_{\rm int}+\bar{k}]_{N}}{N}},$ (44)
-if $0\leq l\leq l_{{\rm int}}-1$, that is consistent with [8, Theorem 2]. ∎
+if $0\leq l\leq l_{{\rm int}}-1$, that is consistent with [6, Theorem 2].
 Remark 6.
-Following the discussion in Remark 4, the two-dimensional sequence $\alpha\Psi_{k^{\prime},l^{\prime}}[k,l]$ can be regarded as the target response in the delay-Doppler domain for an input symbol located on the delay-Doppler grid point indexed by $k^{\prime}$ and $l^{\prime}$.
-We underline that the target response depends now upon the location of the input symbol in the delay-Doppler plane, which is in sharp contrast to the result previously obtained for ideal shaping filters.
-For a given $B$, the sequence $\Psi_{k^{\prime},l^{\prime}}[k,l]$ in (38) is entirely specified by the initial range and the range-rate of the target and satisfy $|\Psi_{k^{\prime},l^{\prime}}[k,l]|\leq 1$: this latter result is instead similar to that for ideal shaping filters.
-∎
-IV-A Vectorized model
-As in Sec III-B, the samples $\{Y_{\rm DD}[k,l]\}$ can be organized into the vector $\bm{y}\in\mathbb{C}^{NM}$ representing the target echo, where $(\bm{y})_{lN+k+1}=Y_{\rm DD}[k,l]$, for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$.
-From (37), we have
+For large $N/B$ and $M$, the sum in (35) only contains few dominant terms; indeed, we have $|\Psi_{k^{\prime},l^{\prime}}[k,l]|\approx 0$ if $k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$, for $b=0,\ldots,B-1$.
+This result is similar to that obtained for ideal shaping filters in Remark 3.
+Refer to caption    Refer to caption    Refer to caption
+Refer to caption    Refer to caption    Refer to caption
+Figure 3: $|\Psi_{k^{\prime},l^{\prime}}[k,l]|$ vs $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$, when $B=4$, $N=M=64$, and rectangular shaping filters are employed.
+Top and bottom rows are for $(k^{\prime},l^{\prime})=(0,0)$ and $(32,32)$, respectively.
+Cases 1 (left), 2 (center), and 3 (right) are defined as in Fig.
+2.
+Example 2.
+Fig.
+3 reports $|\Psi_{k^{\prime},l^{\prime}}[k,l]|$ versus $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ under the same parameter settings considered in Example 1, for the input locations $(l^{\prime},k^{\prime})=(0,0)$ and $(32,32)$.
+The resulting patterns show that the target responses exhibit similar overall support in the delay-Doppler domain, even though their shapes depend on the input location.
+For $(l^{\prime},k^{\prime})=(32,32)$, the Doppler profile is also affected by the term $\mathbbm{1}_{\{b=0\}}B/N$ in (38), which modifies the associated Dirichlet kernel due to the ISI induced by the rectangular shaping filters; this effect is not present for $(l^{\prime},k^{\prime})=(0,0)$.
+III-B1 Vectorized model
+The samples $\{Y_{\rm DD}[k,l]\}$ can again be arranged into the vector $\bm{y}\in\mathbb{C}^{NM}$. From (35), we obtain
 $\bm{y}=\alpha\bm{\Psi}\bm{x},$ (45)
-where $\bm{\Psi}$ is an $(NM\times NM)$-dimensional matrix with $(\bm{\Psi})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Psi_{k^{\prime},l^{\prime}}\big[k-k^{\prime},l-l^{\prime}\big]$.
-Interestingly, the models in (31) and (47) present the same structure; however, the involved matrices $\bm{\Phi}$ and $\bm{\Psi}$ are different.
-For future reference, we also define the normalized target echo as
-$\bm{e}(d,v,B)=\bm{\Psi}(d,v,B)\bm{x},$ (48)
-where the dependence upon the initial range $d$ of the target, the range-rate $v$ of the target, and $B$ has been made explicit.
-IV-B Approximated model
-We derive here an approximated model: the goal is to obtain an input-output relationship with a structure similar to the one in (32), which will turn out useful in Sec. V-A.
+where $\bm{\Psi}\in\mathbb{C}^{NM\times NM}$ has entries $(\bm{\Psi})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Psi_{k^{\prime},l^{\prime}}\big[k-k^{\prime},l-l^{\prime}\big]$.
+Note that the models in (30) and (45) share the same structure; however, the matrices $\bm{\Phi}$ and $\bm{\Psi}$ differ.
+We will denote $\bm{\Psi}$ by $\bm{\Psi}(d,v)$ when we want to make explicit its dependence on the target initial-range and range-rate.
+The input-output relationship in (35), or equivalently (45), captures the combined effects of range migration, ICI, and ISI.
+We next introduce a more tractable approximated model that preserves the dominant delay-Doppler spreading effects induced by range migration while simplifying the dependence of the target response on the transmitted-symbol location.
+III-B2 Approximated model
 Notice first that (37) and (38) differ by the term
 $e^{-\mathrm{i}2\pi\frac{k^{\prime}}{N}}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{d}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}-M}{M}}\\ \times\frac{B}{N}\mathcal{D}_{M}\!\left(-\left(\frac{l+M}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right)\mathbbm{1}_{\{b=0\}},$ (46)
-whose magnitude tends to zero for large $N/B$. Upon neglecting the above term in (40) and upon ignoring the fractional part of the target range, then (41) can be approximated as
+whose magnitude tends to zero as $N/B$ increases. Upon neglecting the above term in (38) and ignoring the fractional part of the target range, then (39) is approximated as
 $Y_{{\rm DD}}[k,l]\approx\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{NM}l}\Theta_{k^{\prime},l^{\prime}}[k,l]$
@@
 Based on (47), we can now write
-$\bm{y}\approx\alpha\big(\bm{X}\odot\bm{\Theta}\odot\bm{\Lambda}(v)\big)\bm{\phi}(d,v,B),$ (52)
-where $\bm{X}$ and $\bm{\phi}(d,v,B)$ are the same quantities in (32), while $\bm{\Theta}$ and $\bm{\Lambda}(v)$ are $NM\times NM$ matrices with $(\bm{\Theta})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Theta_{k^{\prime},l^{\prime}}[k,l]$ and $\big(\bm{\Lambda}(v)\big)_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{NM}l}=\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda M}l}$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$.
+$\bm{y}\approx\alpha\big(\bm{X}\odot\bm{\Theta}\odot\bm{\Lambda}(v)\big)\bm{\phi}(d,v),$ (49)
+where $\bm{X}$ and $\bm{\phi}(d,v)$ are the same quantities in (31), while $\bm{\Theta}$ and $\bm{\Lambda}(v)$ are $NM\times NM$ matrices with $(\bm{\Theta})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Theta_{k^{\prime},l^{\prime}}[k,l]$ and $\big(\bm{\Lambda}(v)\big)_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{NM}l}=\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda M}l}$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$.
 To proceed, notice that $\{\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda M}l}\}_{l=-\infty}^{\infty}$ is a complex exponential whose numerical frequency depends upon the target range-rate.
@@
 Hence, we finally obtain
-$\bm{y}\approx\alpha\bm{\Xi}\bm{\phi}(d,v,B),$ (53)
+$\bm{y}\approx\alpha\bm{\Xi}\bm{\phi}(d,v),$ (50)
 where $\bm{\Xi}=\bm{X}\odot\bm{\Theta}\odot\bm{\Lambda}(v_{\max}/2)$.
-By comparing (32) with (53), it is seen that the symbol matrix $\bm{X}$ in (32) is now replaced in (52) by the modified symbol matrix $\bm{\Xi}$ to account for the presence of rectangular filters.
-To compare (47) and (53), we compute the correspoding normalized root mean square error (NRMSE), namely, $(\mathrm{E}\left[\|\bm{\Xi}\bm{\phi}(d,v,B)-\bm{\Psi}(d,v,B)\bm{x}\|^{2}\right]/\mathrm{E}\left[\|\bm{\Psi}(d,v,B)\bm{x}\|^{2}\right])^{1/2}$, where the statistical expectation is with respect to $\bm{x}$, $d$, and $v$.
+By comparing (31) with (50), it is seen that the symbol matrix $\bm{X}$ in (31) is now replaced in (49) by the modified symbol matrix $\bm{\Xi}$ to account for the presence of rectangular shaping filters.
+To compare (45) and (50), we compute the corresponding normalized root mean square error (NRMSE), namely, $(\mathrm{E}\left[\|\bm{\Xi}\bm{\phi}(d,v)-\bm{\Psi}(d,v)\bm{x}\|^{2}\right]/\mathrm{E}\left[\|\bm{\Psi}(d,v)\bm{x}\|^{2}\right])^{1/2}$, where the statistical expectation is with respect to $\bm{x}$, $d$, and $v$.
 For the system setup considered in Sec.
-VI, we found an NRMSE of about $0.07$ for any $B$, thus confirming the goodness of the approximated model in (53).
-V Estimation of the target parameters
-Upon accounting for the presence of the additive noise, after OTFS demodulation the received signal can be written as
+V, we found an NRMSE of about $0.07$ for any choice of the receiver design parameter $B$, thus confirming the goodness of the approximated model in (50).
+IV Estimation of the target parameters
+Upon accounting for the additive noise, after OTFS demodulation the received signal can be written as
 $\bm{z}=\bm{y}+\bm{\omega}\in\mathbb{C}^{NM},$ (51)
-where $\bm{y}$ is the target echo and $\bm{\omega}$ is a complex circularly-symmetric Gaussian vector with covariance matrix $\sigma_{\omega}^{2}\bm{I}_{NM}$.
-Based on $\bm{z}$ and on the echo model $\bm{y}=\alpha\bm{e}(d,v,B)$ provided in Secs.
-III-B and IV-A, the radar receiver is faced with the problem of estimating the unknown target parameters, namely, $d$, $v$, and $\alpha$.
-We resort here to an ML estimation procedure; for a given $B$, the corresponding estimates are [45]
-$\big[\hat{d}(B);\,\hat{v}(B)\big]$ $=\operatorname*{\arg\max}_{\begin{subarray}{c}[d;\,v]\in\mathcal{S}\end{subarray}}\frac{\big|\bm{e}(d,v,B)^{\mathsf{H}}\bm{z}\big|^{2}}{\|\bm{e}(d,v,B)\|^{2}},$ (55a)
-$\hat{\alpha}(B)$ $=\frac{\bm{e}\big(\hat{d}(B),\hat{v}(B),B\big)^{\mathsf{H}}\bm{z}}{\big\|\bm{e}\big(\hat{d}(B),\hat{v}(B),B\big)\big\|^{2}},$ (55b)
-where $\mathcal{S}=[0,r_{\max}]\times[0,v_{\max}]$.
-In practice, the continuous set $\mathcal{S}$ is replaced by a uniform rectangular grid $\mathcal{G}$ containing $G_{\rm r}=\lfloor r_{\max}/\Sigma_{\rm r}\rfloor$ and $G_{\rm rr}=\lfloor v_{\max}/\Sigma_{\rm rr}\rfloor$ points along the range and range-rate dimensions, respectively, where $\Sigma_{\rm r}\leq R_{\rm r}$ and $\Sigma_{\rm rr}\leq R_{\rm rr}$ are the corresponding point spacing.
-Since the cardinality of $\mathcal{G}$ and the size of $\bm{z}$ scale with $MN$, the computation of (55a) may become demanding for large $M$ and $N$, which is often the case in OTFS-based systems.
-To overcome this problem, we propose a two-step estimation procedure which operates as follows.
-  1. 1.
-    As shown in Sec.
-V-A, we first exploit the signal sparsity to compute a coarse estimate of the target response and, as a by-product, coarse estimates of the initial range and range-rate, denoted as $\tilde{d}(B)$ and $\tilde{v}(B)$, respectively.
-  2. 2.
-    Then, we restrict the search set in (55a) to a neighborhood of the above coarse estimates; in particular, we consider a uniform rectangular grid $\tilde{\mathcal{G}}$ that is centered at $\big[\tilde{d}(B);\,\tilde{v}(B)\big]$ and has $\tilde{G}_{\rm r}$ and $\tilde{G}_{\rm rr}$ points along the range and the range-rate dimension spaced of $\Sigma_{\rm r}$ and $\Sigma_{\rm rr}$, respectively, with $\tilde{G}_{\rm r}\ll G_{\rm r}$ and $\tilde{G}_{\rm rr}\ll G_{\rm rr}$.
-V-A Coarse estimation of the initial range and range-rate
+where $\bm{\omega}$ is a circularly symmetric complex Gaussian vector with covariance matrix $\sigma_{\omega}^{2}\bm{I}_{NM}$.
+Based on $\bm{z}$ and the target echo model derived in Sec.
+III, the radar receiver is faced with the problem of estimating the unknown target parameters, namely $d$, $v$, and $\alpha$.
+The corresponding ML estimates are [40]
+[…]
+$\hat{\alpha}$ $=\frac{\bm{e}\big(\hat{d},\hat{v}\big)^{\mathsf{H}}\bm{z}}{\big\|\bm{e}\big(\hat{d},\hat{v}\big)\big\|^{2}},$ (52b)
+where $\mathcal{S}=[0,r_{\max}]\times[0,v_{\max}]$, and
+$\bm{e}(d,v)=\begin{cases}\bm{\Phi}(d,v)\bm{x},&\text{for ideal shaping filters},\\ \bm{\Psi}(d,v)\bm{x},&\text{for rectangular shaping filters}.\end{cases}$ (53)
+In practice, the continuous set $\mathcal{S}$ is replaced by a uniform rectangular grid $\mathcal{G}$ containing $G_{\rm r}=\lfloor r_{\max}/\Sigma_{\rm r}\rfloor$ and $G_{\rm rr}=\lfloor v_{\max}/\Sigma_{\rm rr}\rfloor$ points along the initial-range and range-rate domains, respectively, where $\Sigma_{\rm r}\leq R_{\rm r}$ and $\Sigma_{\rm rr}\leq R_{\rm rr}$ are the corresponding point spacing.
+A direct implementation of (52) entails a complexity of
+$\mathcal{O}\big(G_{\rm r}G_{\rm rr}(MN)^{2}\big).$ (54)
+In particular, the evaluation of (52) may become computationally demanding when a fine search grid is required.
+To alleviate this burden, we propose in Sec.
+IV-A a reduced-complexity two-step estimation procedure that exploits the sparse structure of the target echo.
+This procedure is further extended in Sec.
+IV-B to also handle multiple targets.
+IV-A Two-step estimation
+[…]
+IV-A1 Coarse estimation of the initial-range and range-rate
 Consider first ideal shaping filters.
-Then, from (32) the received signal can be expended as $\bm{z}=\bm{X}\bm{f}+\bm{\omega}$, where $\bm{f}=\alpha\bm{\phi}$ is the target response vector.
-[^5] For brevity, the dependence upon $d$, $v$, and $B$ is kept concealed here.
-To proceed, notice that $\bm{f}$ can be partitioned into $M$ sub-vectors (blocks) of length $N$, namely, $\bm{f}=(\bm{f}_{0};\cdots;\bm{f}_{M-1})$, where $\bm{f}_{m}=\big(\alpha\Phi[0,m];\ldots;\alpha\Phi[N-1,m]\big)$; according to Remark 2, $\bm{f}$ is a block sparse vector for large $N/B$ and $M$.
-Notice next that the size of $\bm{z}$ can be reduced by employing a $NK\times NM$ selection matrix $\bm{S}$ whose rows are formed by any $NK$ distinct rows of $\bm{I}_{NM}$; the parameter $K\in\{1,\ldots,M\}$ rules the implementation complexity of the coarse estimator, and in practice we can select $K\ll M$.
-The compressed measurement $\bm{z}_{s}=\bm{S}\bm{z}\in\mathbb{C}^{NK}$ can be expanded as
+From (31), the received signal can be written as $\bm{z}=\bm{X}\bm{f}+\bm{\omega}$, where $\bm{f}=\alpha\bm{\phi}$ is the target response vector.
+For brevity, the dependence on $d$, $v$, and $B$ is omitted in this section.
+To proceed, $\bm{f}$ is partitioned into $M$ subvectors (blocks) of length $N$, namely, $\bm{f}=(\bm{f}_{0};\cdots;\bm{f}_{M-1})$, where $\bm{f}_{m}=\big(\alpha\Phi[0,m];\ldots;\alpha\Phi[N-1,m]\big)$.
+According to Remark 3, $\bm{f}$ is block-sparse for large $N/B$ and $M$.
+Next, the dimension of $\bm{z}$ is reduced by means of an $NK\times NM$ selection matrix $\bm{S}$ whose rows are distinct rows of $\bm{I}_{NM}$.
+The parameter $K\in\{1,\ldots,M\}$ controls the implementation complexity.
+The compressed measurement $\bm{z}_{s}=\bm{S}\bm{z}\in\mathbb{C}^{NK}$ can be expressed as
 $\bm{z}_{s}=\bm{X}_{s}\bm{f}+\bm{\omega}_{s}=\sum_{m=0}^{M-1}\bm{X}_{s,m}\bm{f}_{m}+\bm{\omega}_{s},$ (55)
@@
 then, an estimate of the sub-vectors $\bm{f}_{i^{(1)}},\ldots,\bm{f}_{i^{(\ell)}}$ corresponding to the currently selected blocks of $\bm{X}_{s}$ is computed by solving a least-square problem, namely,
-$\big(\tilde{\bm{f}}_{i^{(1)}}\cdots\tilde{\bm{f}}_{i^{(\ell)}}\big)=\operatorname*{\arg\max}_{(\bm{f}_{i^{(1)}}\cdots\bm{f}_{i^{(\ell)}})\in\mathbb{C}^{N\times\ell}}\Big\|\bm{z}_{s}-\sum_{q=1}^{\ell}\bm{X}_{s,i^{(q)}}\bm{f}_{i^{(q)}}\Big\|^{2};$ (58)
+$\big(\tilde{\bm{f}}_{i^{(1)}}\cdots\tilde{\bm{f}}_{i^{(\ell)}}\big)=\operatorname*{\arg\min}_{(\bm{f}_{i^{(1)}}\cdots\bm{f}_{i^{(\ell)}})\in\mathbb{C}^{N\times\ell}}\Big\|\bm{z}_{s}-\sum_{q=1}^{\ell}\bm{X}_{s,i^{(q)}}\bm{f}_{i^{(q)}}\Big\|^{2};$ (57)
 finally, the residual vector and the search set are updated as $\bm{z}^{(\ell+1)}_{s}=\bm{z}^{(\ell)}_{s}-\sum_{q=1}^{\ell}\bm{X}_{s,i^{(q)}}\tilde{\bm{f}}_{i^{(q)}}$ and $\mathcal{I}^{(\ell+1)}=\mathcal{I}^{(\ell)}\setminus\{i^{(\ell)}\}$, respectively, and the procedure is stopped when $\ell=K$ or $\|\bm{z}^{(\ell+1)}_{s}\|^{2}<\epsilon NK\sigma_{\omega}^{2}$, with $\epsilon$ being a positive threshold.
 If the procedure ends in $\tilde{\ell}$ iterations, an estimate $\tilde{\bm{f}}$ of the target response vector is obtained from the indexes $i^{(1)},\ldots,i^{(\tilde{\ell})}$ and the sub-vectors $\tilde{\bm{f}}_{i^{(1)}},\ldots,\tilde{\bm{f}}_{i^{(\tilde{\ell})}}$.
-At this point, the entries of $\tilde{\bm{f}}$ are arranged into the target response sequence $\tilde{F}[k,l]=(\tilde{\bm{f}})_{lN+k+1}$, for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$, which provides an estimate of $\alpha\Phi[k,l]$; finally, coarse estimates of the initial range and range-rate of target are obtained from the location of the maximum element of $|\tilde{F}[k,l]|$, in keeping with the discussion provided in Remark 4.
+At this point, the entries of $\tilde{\bm{f}}$ are arranged into the sequence $\tilde{F}[k,l]=(\tilde{\bm{f}})_{lN+k+1}$, for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$, which provides an estimate of $\alpha\Phi[k,l]$; finally, coarse estimates of the target initial-range and range-rate are obtained from the location of the maximum element of $|\tilde{F}[k,l]|$.
 When a rectangular shaping filter is employed, from (50), the received signal can be approximately expanded as $\bm{z}\approx\alpha\bm{\Xi}\bm{\phi}+\bm{\omega}$; accordingly, the above procedure can still be used after replacing $\bm{X}$ by the modified symbol matrix $\bm{\Xi}$.
-V-B Extension to multiple targets
-The above estimation procedure can be combined with the CLEAN algorithm in [37, 38, 39, 40, 16] to handle multiple targets at low complexity.
-The CLEAN algorithm iteratively reconstructs one target echo at the time, after removing the interference caused by the previously-estimated target echoes from the data.
-More specifically, assume that $P>1$ targets are present in the inspected region and denote by $\hat{\bm{y}}_{p}(B)=\hat{\alpha}_{p}(B)\bm{e}\big(\hat{d}_{p}(B),\hat{v}_{p}(B),B\big)$ the reconstructed target echo in the $p$-th iteration, where $\hat{d}_{p}(B)$, $\hat{v}_{p}(B)$, and $\hat{\alpha}_{p}(B)$ are the estimated initial range, range-rate, and amplitude, respectively; then, the proposed two-step estimation procedure is iterated on the cleaned data $\bm{z}-\sum_{i=1}^{p-1}\hat{\bm{y}}_{i}(B)$, for $p=1,\ldots,P$.
+IV-A2 Implementation complexity
+We now discuss the implementation complexity of the proposed two-step estimator.
+In the first step, the BOMP algorithm is applied to the compressed model in (55).
+At the $\ell$-th iteration, the block-selection requires the computation of $\bm{X}_{s,i}^{\rm H}\bm{z}_{s}^{(\ell)}$ for all remaining blocks; since $\bm{X}_{s,i}\in\mathbb{C}^{NK\times N}$, this operation has complexity $\mathcal{O}((M-\ell+1)KN^{2})$.
+The following least-squares update involves the matrix obtained by concatenating the $\ell$ selected blocks; a standard dense least-squares implementation therefore requires $\mathcal{O}(K\ell^{2}N^{3}+\ell^{3}N^{3})$ operations.
+The residual updates have lower complexity and do not affect the dominant scaling.
+Hence, summing over at most $K$ iterations yields a complexity $\mathcal{O}(MK^{2}N^{2}+K^{4}N^{3})$ for the first step.
+In the second step, the metric in (52a) is evaluated over the local grid $\tilde{\mathcal{G}}$, with complexity $\mathcal{O}\big(\tilde{G}_{\rm r}\tilde{G}_{\rm rr}(MN)^{2}\big)$.
+Therefore, the overall complexity of the proposed two-step estimator is
+$\mathcal{O}\!\left(MK^{2}N^{2}+K^{4}N^{3}+\tilde{G}_{\rm r}\tilde{G}_{\rm rr}(MN)^{2}\right).$ (58)
+Since in practice $K\ll M$ and $\tilde{G}_{\rm r}\tilde{G}_{\rm rr}\ll G_{\rm r}G_{\rm rr}$, this complexity is substantially lower than that in (54).
+IV-B Extension to multiple targets
+Multiple targets can be handled through the iterative CLEAN algorithm [32, 33, 34, 35], which repeatedly estimates and removes the dominant target echo from the received signal.
+Let $\bm{r}^{(1)}=\bm{z}$ denote the initial residual.
+At the $p$-th iteration, the proposed two-step estimator is applied to $\bm{r}^{(p)}$ to extract the dominant remaining target and estimate its initial-range $\hat{d}^{(p)}$, range-rate $\hat{v}^{(p)}$, and amplitude $\hat{\alpha}^{(p)}$.
+The corresponding echo is then reconstructed as $\hat{\bm{y}}^{(p)}=\hat{\alpha}^{(p)}\bm{e}\big(\hat{d}^{(p)},\hat{v}^{(p)}\big)$, and the residual is updated as $\bm{r}^{(p+1)}=\bm{r}^{(p)}-\hat{\bm{y}}^{(p)}$ to mitigate the interference caused by previously extracted targets in the next iteration $p+1$.
+If the number of targets is known, say $P$, the procedure is repeated exactly $P$ times.
+Otherwise, the iterations are stopped when $\|\bm{r}^{(p+1)}\|^{2}$ falls below a given threshold or when a maximum number of iterations is reached.
+Since each iteration requires one execution of the proposed two-step estimator, with complexity given in (58), together with a residual update of complexity $\mathcal{O}(MN)$, the overall complexity scales linearly with the number of extracted targets.
+If $P$ targets are processed, the resulting complexity is
+$\mathcal{O}\big(P\big(MK^{2}N^{2}+K^{4}N^{3}+\tilde{G}_{\rm r}\tilde{G}_{\rm rr}(MN)^{2}\big)\big).$ (59)
+V Numerical analysis
+To assess the system performance, we consider the parameters in Table I.
+For these values, the maximum variation of the target range during one OTFS frame is $\bar{v}_{\rm max}NT\approx 136$ m, which corresponds to about $3.5$ times the range resolution $R_{\rm r}$.
+The initial-range $\bar{d}$ and range-rate $\bar{v}$ are randomly generated within the intervals $\bar{r}_{\rm min},\,\bar{r}_{\rm max}]$ and $\bar{v}_{\rm min},\,\bar{v}_{\rm max}]$, respectively, while a Swerling I fluctuation model is adopted for the amplitude $\alpha$.
+The information symbols $\{X_{\rm DD}[k,l]\}$ are independently drawn from a unit-energy $4$-QAM constellation.
+The entries of the observed vector $\bm{z}$ are generated according to (12), (15), and (19).
+The radar receiver, instead, is designed according to the model developed in Secs.
+III-A and III-B for ideal and rectangular shaping filters, respectively.
+Unless otherwise stated, we set $B=16$.
+In the implementation of the proposed two-step estimator, we use $\Sigma_{\rm r}=R_{\rm r}/100$, $\Sigma_{\rm rr}=R_{\rm rr}/100$, $\tilde{G}_{\rm r}=401$, $\tilde{G}_{\rm rr}=401$, $K=4P$, and $\epsilon=1$.
+For future reference, we define the target SNR as
+$\mathrm{SNR}=\mathrm{E}[|\alpha|^{2}]NM/\sigma_{\omega}^{2}.$ (60)
+Also, let $d$, $v$, and $\alpha$ denote the true values of the initial-range, range-rate, and amplitude of the target, and let $\hat{d}$, $\hat{v}$, and $\hat{\alpha}$ be the corresponding estimates.
+We define the root mean square error (RMSE) of the initial-range and range-rate estimates, and the NRMSE of the amplitude estimate, as $\mathrm{RMSE}_{\rm r}=(\mathrm{E}[|d-\hat{d}|^{2}])^{1/2}$, $\mathrm{RMSE}_{\rm rr}=(\mathrm{E}[|v-\hat{v}|^{2}])^{1/2}$, and $\mathrm{NRMSE}_{\rm\alpha}=(\mathrm{E}[|\alpha-\hat{\alpha}|^{2}]/\mathrm{E}[|\alpha|^{2}])^{1/2}$, respectively.
+In the following, we first examine the single-target case and then the multi-target scenario.
 Table I: System parameters
@@
 $R_{\rm rr}$ Range-rate resolution $\approx$ 8.8 m/s
-$\bar{r}_{\rm min},\,\bar{r}_{\rm max}]$ Inspected range interval [20, 39] km
+$\bar{r}_{\rm min},\,\bar{r}_{\rm max}]$ Inspected initial-range interval [20, 39] km
 $\bar{v}_{\rm min},\,\bar{v}_{\rm max}]$ Inspected range-rate interval [15, 16] km/s
-VI Numerical analysis
-To assess the system performance, we consider the parameters in Table I.
-The maximum variation of the target range during the OTFS frame is $\bar{v}_{\rm max}NT\approx 136$ m, corresponding to about $3.5$ times the range resolution $R_{\rm r}$; the initial range $d$ and the range-rate $v$ are randomly generated, while a Swerling I fluctuation model is considered for the amplitude $\alpha$.
-The information symbols $\{X_{\rm DD}[k,l]\}$ are independently drawn from a unit-energy $4$-QAM constellation.
-The entries of the observed vector $\bm{z}$ are computed according to Eqs.
-(11), (14), and (18).
-Instead, the radar receiver is designed according to the echo model proposed in Secs.
-III and IV for ideal and rectangular shaping filters, respectively, that relies upon the $B$-fold stop-and-go approximation.
-In the implementation of the proposed two-step estimator, we set $\Sigma_{\rm r}=R_{\rm r}/100$, $\Sigma_{\rm rr}=R_{\rm rr}/100$, $\tilde{G}_{\rm r}=401$, $\tilde{G}_{\rm rr}=401$, $K=4P$, and $\epsilon=1$.
-For future reference, we define the signal-to-noise ratio (SNR) of the target as $\mathrm{SNR}=\mathrm{E}[|\alpha|^{2}]NM/\sigma_{\omega}^{2}$.
-Also, let $d$, $v$, and $\alpha$ be the true values of the initial range, range-rate, and amplitude of the target, and let $\hat{d}$, $\hat{v}$, and $\hat{\alpha}$ be the corresponding estimates; then, we define the root mean square error (RMSE) in the estimation of the initial range, the RMSE in the estimation of the range-rate, and the NRMSE in the estimation of the amplitude as $\mathrm{RMSE}_{\rm r}=(\mathrm{E}[|d-\hat{d}|^{2}])^{1/2}$, $\mathrm{RMSE}_{\rm rr}=(\mathrm{E}[|v-\hat{v}|^{2}])^{1/2}$, and $\mathrm{NRMSE}_{\rm\alpha}=(\mathrm{E}\left[|\alpha-\hat{\alpha}|^{2}\right]/\mathrm{E}\left[|\alpha|^{2}\right])^{1/2}$, respectively.
-Next, we discuss two examples with one and multiple targets, respectively.
-VI-A Single target
+V-A Single target
 Refer to caption Refer to caption Refer to caption
 Figure 4: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $B=1,2,4,8,16,32,64,128$ for $\mathrm{SNR}=10,20,30$ dB, when a high- or low-speed target is present and ideal shaping filters are employed.
-Both the ML and the proposed estimator are considered for comparison.
+Both the ML and the proposed two-step estimator are considered for comparison.
 Refer to caption Refer to caption Refer to caption
-Figure 4: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $\mathrm{SNR}$ for $B=4,8,16$, when a high-speed target is present and the proposed estimator is employed.
+Figure 5: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $\mathrm{SNR}$ for $B=4,8,16$, when a high-speed target is present, and the proposed two-step estimator is employed.
 Both ideal and rectangular shaping filters are considered.
 Refer to caption Refer to caption Refer to caption
+Figure 6: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $B=1,2,4,8,16,32,64,128$ for different values NN and $\Delta$, when a high-speed target is present, rectangular shaping filters are employed, $\mathrm{SNR}=30$ dB, and the proposed two-step estimator is employed.
+Refer to caption Refer to caption Refer to caption
 Figure 7: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $\bar{v}_{\rm min}$, when rectangular filters are used, $\mathrm{SNR}=15$ dB, and $\bar{v}_{\rm max}=\bar{v}_{\rm min}+1$ km/s.
-The proposed estimator is implemented for $B=1,2,4,8,16$.
-For comparison, the performance of the estimators presented in [21] and [29] is also evaluated.
+The proposed two-step estimator is implemented for $B=1,2,4,8,16$.
+For comparison, the threshold-based estimator in [17] and the SOMP-based estimator in [24] are also included.
+We first examine the system performance with ideal shaping filters. We then consider rectangular shaping filters.
+V-A1 Ideal shaping filters
+Fig.
+4 presents the metrics $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\rm\alpha}$ as functions of $B$ for $\mathrm{SNR}=10,20,30$ dB.
+Observe that the parameters of a high-speed target cannot be accurately recovered when range variation during the OTFS frame is neglected at the system design (i.e., $B=1$).
+As $B$ increases from 1 to 16, estimation accuracy improves significantly, while further increases in $B$ yield diminishing returns.
+This behavior is governed by the ratio between the maximum range variation during one block of $N/B$ symbols and the range resolution, namely
+$\frac{\bar{v}_{\max}NT/B}{R_{\rm r}}=\frac{\bar{v}_{\max}MN}{cB},$ (61)
+which should be sufficiently smaller than one for accurate modeling.
+For the considered parameters, this ratio equals approximately $3.5/B$.
+Hence, for $B=1$, the blockwise constant-range approximation is inaccurate, whereas for $B=16$ the ratio reduces to about $0.22$, corresponding to a maximum range variation over $N/B$ symbols of approximately $8.5$ m.
+Such variation can be neglected with little performance loss.
 […]
-Next, we consider the use of rectangular shaping filters.
+Finally, note that the proposed two-step estimator achieves performance close to that of the ML estimator across all scenarios considered in Fig.
+4, at a much lower implementation complexity, underscoring its effectiveness and practical utility.
+V-A2 Rectangular shaping filters
 Fig.
-4 shows the behavior of $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\rm\alpha}$ as functions of $\mathrm{SNR}$ when the proposed estimator is employed and $B=4,8,16$.
+5 shows the behavior of $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\rm\alpha}$ as functions of $\mathrm{SNR}$ when the proposed two-step estimator is employed and $B=4,8,16$.
 For comparison, results obtained with ideal shaping filters are also shown.
-Notably, the more practical rectangular filters achieve estimation performance that closely approximates that of the ideal filters.
-Finally, we compare the performance of the proposed estimator with the methods in [21] and [29], which do not explicitly account for range migration within the OTFS frame.
-To evaluate performance across varying mobility scenarios, we set $\bar{v}_{\rm max}=\bar{v}_{\rm min}+1$ km/s and vary $\bar{v}_{\rm min}$ from 0 to 15 km/s.
+Notably, the more practical rectangular filters achieve performance close to that of the ideal filters.
 Fig.
-5 plots the metrics $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\alpha}$ versus $\bar{v}_{\rm min}$ for rectangular shaping filters and $\mathrm{SNR}=15$ dB.
+6 investigates the influence of the OTFS parameters $\Delta$ and $N$ on the estimation performance when the proposed two-step estimator is employed and $\mathrm{SNR}=30$ dB.
+The results show that the initial-range estimation accuracy improves as the subcarrier spacing $\Delta$ increases, since the occupied bandwidth $M\Delta$ becomes larger, whereas it is only marginally affected by variations of $N$ when SNR is kept constant.
+In contrast, the range-rate estimation accuracy benefits from increasing $N$ and reducing $\Delta$, since both choices enlarge the coherent observation time $NT$ (recall that $T=1/\Delta$).
+As for the impact of the parameter $B$, the ratio introduced in (61) remains the key quantity.
+Note that this ratio is independent of $\Delta$ and scales linearly with $N$.
+Therefore, the pairs $(N,\Delta)=(128,15\penalty 10000\ \text{kHz})$ and $(128,30\penalty 10000\ \text{kHz})$, represented by solid lines, exhibit a similar validity region for the blockwise constant-range approximation, whereas the remaining pairs with $N=256$, represented by dashed lines, require approximately twice the value of $B$ to achieve a comparable level of approximation accuracy.
+Finally, we evaluate the estimation performance of the proposed two-step estimator across varying mobility scenarios; to this end, we set $\bar{v}_{\rm max}=\bar{v}_{\rm min}+1$ km/s and vary $\bar{v}_{\rm min}$ from 0 to 15 km/s.
+For comparison, we also include the threshold-based estimator in [17] and the SOMP-based estimator in [24], both of which are designed under the assumption of a constant integer range within the OTFS frame.
+To ensure a fair comparison, the input delay-Doppler sequence $X_{{\rm DD}}[k,l]$ contains a single non-zero entry located at $(k,l)=(64,256)$, whereas all remaining entries are set to zero to create a guard region.
+[^8] In this setup, the complexity of the proposed two-step estimator simplifies to $\mathcal{O}\big(MK^{2}N^{2}+K^{4}N^{3}+\tilde{G}_{\text{r}}\tilde{G}_{\text{rr}}MN\big)$ because the vector $\bm{x}$ contains a single non-zero entry at a known location and the matrix-vector multiplication in (53) becomes trivial during the second step.
+In contrast, the benchmarks from [17] and [24] both have complexity $\mathcal{O}(MN)$.
+Fig.
+7 plots the metrics $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\alpha}$ versus $\bar{v}_{\rm min}$ for $\mathrm{SNR}=15$ dB.
 The performance of the estimators in [17] and [24] deteriorates rapidly when $\bar{v}_{\rm min}$ exceeds 1 km/s.
 In this regime, the maximum range variation during the OTFS frame, given by $\bar{v}_{\rm max}NT$, becomes comparable to or exceeds the range resolution $R_{\rm r}$, resulting in a mismatch between the assumed and actual echo models.
-In contrast, the proposed estimator maintains robust performance across a broader range of $\bar{v}_{\rm min}$ by increasing $B$, thereby enabling a finer stop-and-go approximation in (7).
-Performance degradation only occurs when the range variation over $N/B$ symbol intervals, i.e., $\bar{v}_{\rm max}NT/B$, becomes comparable to or exceeds $R_{\rm r}$.
-Notably, with $B=16$, the proposed estimator delivers consistent performance across all tested values of $\bar{v}_{\rm min}$, in keeping with the findings in Fig.
-3.
-Furthermore, as $\bar{v}_{\rm min}\rightarrow 0$, its performance aligns with that of the existing methods, confirming its effectiveness and practical relevance even for low-speed targets.
-VI-B Multiple targets
+In contrast, the proposed two-step estimator remains robust over a significantly wider mobility range by using larger values of $B$, which enable a finer blockwise constant-range approximation in (8).
+In particular, $B=16$ provides stable performance over all considered values of $\bar{v}_{\rm min}$, consistently with the trends already observed in Figs.
+4 and 6.
+Even when $\bar{v}_{\rm min}\rightarrow 0$, the proposed estimator still outperforms the existing methods due to the second refinement step, which improves the estimation accuracy at the cost of higher computational complexity.
+V-B Multiple targets
 Refer to caption Refer to caption Refer to caption
@@
 Refer to caption Refer to caption Refer to caption
-Figure 7: $|\bm{F}[k,l]|/\sigma_{\omega}$ vs $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ at the first (top), second (middle), and third (bottom) iteration of the CLEAN procedure, when rectangular filters are employed, $P=3$ high-speed targets are present, $B=16$, $\mathrm{SNR}_{1}=30$ dB, $\mathrm{SNR}_{2}=20$ dB, and $\mathrm{SNR}_{3}=10$ dB.
-We use here rectangular shaping filters and consider $P=3$ targets.
+Figure 9: $|\tilde{F}[k,l]|/\sigma_{\omega}$ vs $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ at the first (left), second (center), and third (right) iteration of the CLEAN procedure, when rectangular filters are employed, $P=3$ high-speed targets are present, $B=16$, $\mathrm{SNR}_{1}=30$ dB, $\mathrm{SNR}_{2}=20$ dB, and $\mathrm{SNR}_{3}=10$ dB.
+Refer to caption Refer to caption Refer to caption
+Figure 10: Average $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus range-rate separation, when rectangular filters are used, $P=2$ high-speed targets with the same initial-range are present, $B=16$, and $\mathrm{SNR}_{1}=\mathrm{SNR}_{2}=30$ dB.
+The performance of the proposed CLEAN-based estimator and the corresponding single-target performance bound are reported.
+We use rectangular-shaped filters here and consider $P=3$ targets.
 Let $\mathrm{SNR}_{p}$ be the SNR of Target $p$ for $p=1,\ldots,P$.
@@
 For comparison, we also plot the achievable performance when only a single target is present and the ML estimator in (52) is employed: we refer to it as the single-target bound.
-The proposed procedure is able to effectively estimate the parameters of each target, providing a performance close to the single-target bound.
+The proposed procedure effectively estimates the parameters of each target, achieving performance close to the single-target bound.
 For a single instance and $\mathrm{SNR}_{2}=20$ dB, Fig.
-7 shows $|\tilde{F}[k,l]|/\sigma_{\omega}$ versus $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ in each iteration of the CLEAN procedure, where $\tilde{F}[k,l]$ is the target response sequence recovered by the BOMP algorithm.
-It is seen that the echo of Target 1 (the strongest one) is recovered in Iteration 1 and subtracted out from the observed data; then, the echoes of Targets 2 and 3 are recovered in Iterations 2 and 3, respectively.
-VII Conclusions
-In this work, we have investigated the use of OTFS communication signals for estimating the parameters of space targets experiencing range migration, which may occur in LEO satellite applications.
-Our major technical contribution is the derivation of a novel model for the target echo, assuming that the target range changes $B$ times within an OTFS frame; in particular, we have shown that range migration causes a predictable spread in the sparse target response in the delay-Doppler domain.
-Additionally, we have developed an approximate implementation of the maximum likelihood estimator for the target’s initial range, range-rate, and amplitude, which avoids exhaustive search in the delay-Doppler domain.
-If combined with the CLEAN algorithm, the proposed procedure can also handle multiple targets.
-A significant finding is that the estimation accuracy for high-speed targets is similar to that for low-speed targets when range migration is properly accounted for at the design stage.
-Future studies should consider that the radar receiver may not have perfect knowledge of the emitted symbols.
-In such cases, the use of dedicated radar symbols may be considered, and the optimal resource allocation between sensing and communication tasks should be studied.
-Furthermore, using multiple coordinated LEO satellites for signal transmission and/or reception may allow for estimating the initial position and velocity vectors of the target rather than only its initial range and range-rate.
-Finally, the use of multiple OTFS frames of different sizes may provide additional degrees of freedom and help resolve possible range and range-rate ambiguities.
-We provide here the proof to Proposition 1. Upon plugging (35) and (36) in (18), we have
+9 shows $|\tilde{F}[k,l]|/\sigma_{\omega}$ versus $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ in each iteration of the CLEAN procedure, where $\tilde{F}[k,l]$ is the sequence recovered by the BOMP algorithm.
+It is seen that the target echoes are iteratively recovered and subtracted out from the observed data.
+Finally, we assess the ability of the proposed CLEAN-based estimator to resolve two targets having the same initial-range and SNR but different range-rates.
+Fig.
+10 reports $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\alpha}$ (averaged over the two targets) versus the range-rate separation.
+When the range-rate separation is below the nominal resolution $R_{\rm rr}$, the two targets are difficult to distinguish, and the estimation errors increase significantly.
+As the separation increases beyond $R_{\rm rr}$, the two targets become reliably separable, and the performance approaches the single-target bound.
+This confirms that $R_{\rm rr}$ provides a meaningful reference for range-rate discrimination.
+VI Conclusions
+In this work, we have investigated the use of OTFS communication signals for estimating the parameters of space targets experiencing range migration, as may arise in LEO satellite applications.
+Our main technical contribution is the derivation of a novel echo model that allows the target range to vary $B$ times within one OTFS frame.
+In particular, we have shown that range migration induces a spreading of the target response while preserving its sparse structure in the delay-Doppler domain.
+We have also developed a two-step estimator for the target initial-range, range-rate, and amplitude, achieving performance close to that of the maximum likelihood estimator at substantially lower complexity.
+When combined with the CLEAN algorithm, the proposed procedure can handle multiple targets.
+A significant finding is that the achievable estimation accuracy for high-speed targets is comparable to that of low-speed targets when range migration is properly accounted for in the receiver design.
+Future studies should consider that the radar receiver may not have perfect knowledge of the emitted communication symbols.
+In such cases, dedicated radar symbols may be considered, and the resource allocation between sensing and communication tasks should be investigated.
+Furthermore, using multiple coordinated LEO satellites for signal transmission and reception may allow for estimating the initial position and velocity vectors of the target rather than only its initial-range and range-rate.
+Finally, the use of multiple OTFS frames of different sizes may provide additional degrees of freedom and help resolve possible initial-range and range-rate ambiguities.
+Appendix A Proof of Proposition 1
+Upon plugging (12) into (15), we have
+$Y_{\rm TF}[n,m]=\sum_{n^{\prime}=0}^{N-1}\sum_{m^{\prime}=0}^{M-1}\alpha X_{\rm TF}[n^{\prime},m^{\prime}]$
+$\quad\times\int_{-\infty}^{\infty}g_{\rm tx}\!\left(t-n^{\prime}T-\frac{r\left(n^{\prime}T\right)}{c}\right)g^{*}_{\rm rx}\!\left(t-nT\right)$
+$\quad\times\mathrm{e}^{\mathrm{i}2\pi m^{\prime}\Delta\left(t-n^{\prime}T-\frac{r\left(n^{\prime}T\right)}{c}\right)}\mathrm{e}^{-\mathrm{i}2\pi m\Delta\left(t-nT\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}t}dt$
+$=\sum_{n^{\prime}=0}^{N-1}\sum_{m^{\prime}=0}^{M-1}X_{\rm TF}[n^{\prime},m^{\prime}]$
+$\quad\times\alpha\mathrm{e}^{\mathrm{i}2\pi m\Delta\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c}\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\left(n^{\prime}T+\frac{r\left(n^{\prime}T\right)}{c}\right)}$
+$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c}\right)}$
+$\quad\times\int_{-\infty}^{\infty}g_{\rm tx}\!\left(\beta\right)g^{*}_{\rm rx}\!\left(\beta-(n-n^{\prime})T+\frac{r\left(n^{\prime}T\right)}{c}\right)$
+$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left(\beta-(n-n^{\prime})T+\frac{r\left(n^{\prime}T\right)}{c}\right)}d\beta$
+$=\sum_{n^{\prime}=0}^{N-1}\sum_{m^{\prime}=0}^{M-1}X_{\rm TF}[n^{\prime},m^{\prime}]\alpha\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}nT}\mathrm{e}^{-\mathrm{i}2\pi m^{\prime}\Delta\frac{r\left(n^{\prime}T\right)}{c}}$
+$\quad\times\gamma\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c},(m-m^{\prime})\Delta-\frac{v}{\lambda}\right),$
+which completes the proof.
+Appendix B Proof of Proposition 2
+After exploiting (3) and (20), Eq. (19) becomes
+$Y_{\rm DD}[k,l]=\frac{1}{NM}\!\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\!X_{\rm DD}[k^{\prime},l^{\prime}]\\ \times\underbrace{\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}H_{n,m}[n,m]\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}\right)n}\mathrm{e}^{\mathrm{i}2\pi\left(\frac{l-l^{\prime}}{M}\right)m}}_{(\star)}.$ (62)
+Finally, notice that we have
+$(\star)$ $=M\alpha\sum_{n=0}^{N-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}$
+$\quad\times\left[\frac{1}{M}\sum_{m=0}^{M-1}\mathrm{e}^{\mathrm{i}2\pi\left(\frac{l-l^{\prime}}{M}-\frac{r\left(nT\right)\Delta}{c}\right)m}\right]$
+$=M\alpha\sum_{n=0}^{N-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}$
+$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(nT\right)\Delta}{c}\right)\right)$
+$=NM\alpha\sum_{b=0}^{B-1}\frac{1}{B}\left[\frac{B}{N}\sum_{n=bN/B}^{(b+1)N/B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}\right]$
+$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right)$ (63)
+$=NM\alpha\Phi[k-k^{\prime},l-l^{\prime}],$
+where (63) follows from (8).
+Appendix C Proof of Proposition 3
+Upon plugging (33) and (34) in (19), we have
 $Y_{\rm DD}[k,l]$ $=\frac{1}{\sqrt{NM}}\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)}$
@@
 $=\frac{\alpha}{B}\sum_{b=0}^{B-1}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\tfrac{k-k^{\prime}}{N}-\tfrac{vT}{\lambda}\right)\sum_{p=0}^{\big\lceil M-\frac{r\left(bT_{B}\right)M}{cT}\big\rceil-1}\mathcal{D}_{M}\!\left(\tfrac{p+\frac{r\left(bT_{B}\right)}{c}\frac{M}{T}-l}{M}\right)\delta\big[[p-l^{\prime}]_{M}\big]\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{p}{M}}.$ (67)
-Upon expanding $\Pi_{k,l}^{\rm ici}[k^{\prime},l^{\prime}]$ as shown in (62) at the top of this page, we obtain
+Upon expanding $\Pi_{k,l}^{\rm ici}[k^{\prime},l^{\prime}]$ as shown in (67) at the top of the next page, we obtain
 $Y_{{\rm DD}}^{\rm ici}[k,l]$ $=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{\big\lceil M-\frac{r\left(bT_{B}\right)M}{cT}\big\rceil-1}X_{\rm DD}[k^{\prime},l^{\prime}]$
@@
 Upon expanding $\Pi_{k,l}^{\rm isi}[k^{\prime},l^{\prime}]$ as shown in (71) at the top of the next page, we obtain
-$Y_{{\rm DD}}^{\rm ici}[k,l]$ $=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=\big\lceil M-\frac{r\left(bT_{B}\right)M}{cT}\big\rceil}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{-\mathrm{i}2\pi\frac{k^{\prime}}{N}}$
+$Y_{{\rm DD}}^{\rm isi}[k,l]$ $=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=\big\lceil M-\frac{r\left(bT_{B}\right)M}{cT}\big\rceil}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{-\mathrm{i}2\pi\frac{k^{\prime}}{N}}$
 $\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}-M}{M}}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}$